### v2

# LightRetriever: A LLM-based Hybrid Retrieval Architecture with 1000× Faster Query Inference

Abstract
Large Language Models (LLMs)-based hybrid retrieval uses LLMs to encode queries and documents into low-dimensional dense or high-dimensional sparse vectors. It retrieves documents relevant to search queries based on vector similarities. Documents are pre-encoded offline, while queries arrive in real-time, necessitating an efficient online query encoder. Although LLMs significantly enhance retrieval capabilities, serving deeply parameterized LLMs slows down query inference throughput and increases demands for online deployment resources. In this paper, we propose LightRetriever, a novel LLM-based hybrid retriever with extremely lightweight query encoders. Our method retains a full-sized LLM for document encoding, but reduces the workload of query encoding to no more than an embedding lookup. Compared to serving a full-sized LLM on an H800 GPU, our approach achieves over a 1000× speedup for query inference with GPU acceleration, and even a 20× speedup without GPU. Experiments on large-scale retrieval benchmarks demonstrate that our method generalizes well across diverse retrieval tasks, retaining an average of 95% full-sized performance.
1 Introduction
Given search queries and candidate documents, text retrieval employs dual-encoder architectures to encode representation vectors and search documents relevant to queries based on query-document similarities [62]. The encoded vectors could be condensed into low-dimensional vectors for dense retrieval or sparsified into high-dimensional vectors for sparse retrieval. Dense retrieval [29] compresses textual information into dense vectors with parameterized language models (LMs) and retrieves based on inner product similarity. Meanwhile, sparse retrieval [60] characterizes sparse term frequencies on a much broader vocabulary space, and then searches based on lexical overlap and term-based relevance metrics, such as direct impact search (Inner product) [3, 20] and BM25 [60]. Additionally, hybrid retrieval enhances retrieval abilities by interpolating their similarity scores [72] or ranks [15]. Text retrieval is broadly applied in various scenarios, such as web search [43], semantic textual similarity [8], fact checking [66], and retrieval augmented generation [39].
Figure 1: Architecture comparison of LLM-based hybrid retrieval with 1) Full-sized Query Inference (Previous works); and 2) Extreme Lightweight Query Inference (LightRetriever).
Deploying retrieval systems requires different processing workflows for queries and documents. As is shown in Figure 1, as search candidates, documents are encoded with a document encoder $Enc_{d}$ and then indexed into search databases [27, 60]. Because documents are obtained before the search operation, they are pre-computable and distributable, which can be indexed offline and distributed anywhere [79]. On the contrary, queries are input and encoded in real-time, which are non-pre-computable and non-distributable. The query encoder $Enc_{q}$ needs to be served online. Special GPU accelerators are mandatory if the query encoder is a deeply parameterized LM. Most existing works of text retrieval [29, 11, 71] utilize symmetric dual-encoder architectures, which share the same LM for both the query and document encoding or have an equal number of parameters on both sides. Recent state-of-the-art (SOTA) retrievers [71, 5] employ deeply parameterized LLMs as dual-encoders $Enc_{q}$ and $Enc_{d}$ for enhanced retrieval capabilities and domain adaptation abilities on various languages and tasks. However, serving giant LLMs slows down query inference throughput and depletes online deployment resources [37]. Nowadays, LLM-based text retrieval incurs non-negligible online workloads, necessitating special optimizations for query inference efficiency.
Modern Transformers-based LMs handle text modeling with the attention mechanism [68], including the contextual traits and deep interactions of co-occurrences on discrete tokens within certain texts. As is the case for existing LLM-based text retrievers [71, 5], queries and documents are modeled equally with the same LLM, enduring the same computational burdens. In contrast, traditional statistical retrieval algorithms, such as the popular TF-IDF [63] and BM25 [60], only characterize limited token co-occurrence by analyzing lexical overlaps with customized retrieval metrics. They do not handle the contextual traits and deep co-occurrences, and their encoders do not need to be deeply parametrized during inference. Experiments from [65] show that BM25 still performs well in many retrieval tasks, even beating some previous BERT-based retrievers trained on MSMARCO [52].
Built on the above observations, to minimize query service cost, we propose LightRetriever, a novel LLM-based hybrid retrieval architecture targeted at extreme lightweight query inference abilities. As is shown in Figure 1, we retain a full-sized LLM on the document side ($Enc_{d}$) for strong retrieval capacities, but completely remove the deep modeling of queries. 1) For dense retrieval, our method end-to-end trains cacheable token vectors from the LLM-based query encoder. The dense query representations are simply aggregated by averaging corresponding token vectors shown in the texts. All query token vectors are pre-computed and cached into a simple Embedding layer. Query inference costs are drastically reduced to only one Embedding lookup. 2) For sparse retrieval, similar to statistical retrieval algorithms, the mapping of a token and its corresponding frequency [^1] Here we directly use the number of a certain token that appeared in the tokenized query as its frequency. is directly used as the sparse query vector. Thus, no more encoding stage for sparse retrieval is needed. The dense similarity scores are computed with cosine similarity, while the sparse scores directly use inner product similarity. These scores are linearly interpolated [72] for final hybrid retrieval.
LightRetriever is super fast and lightweight with respect to query inference. Compared to serving a full-sized LLM (e.g., Mistral-7B [26]) on an H800 GPU, our method achieves over a 1000× speedup for query inference with GPU acceleration, and even a 20× speedup without GPU. Experiments on large-scale retrieval benchmarks demonstrate that our method generalizes well across diverse LLMs (including Mistral-7B, Llama-1B, 3B, 8B, and Qwen-1.5B, 3B, 7B) and retrieval benchmarks (including 15 English tasks from BeIR [65] and 8 Chinese tasks from CMTEB Retrieval [76]), retaining an average of 95% full-sized performance.
2 Algorithm
Figure 2: Full-sized symmetric dense retrieval for baseline training.
2.1 LM-based text retrieval
Definition
Given a search query $q$ and candidate documents $d=\{d^{+};d^{-}\}\in\mathcal{N}(q)$ (including irrelevant documents $d^{-}$ and relevant documents $d^{+}$) retrieved from indexes $\mathcal{N}$, LM-based text retrieval aims to learn optimized encoders $Enc_{q}$ and $Enc_{d}$ with parameters $\theta$, where the similarity of the query vector $v_{q}=Enc_{q}(q)$ and irrelevant document vectors $v_{d^{-}}=Enc_{d}(d^{-})$ are encouraged to be smaller than those of relevant documents $v_{d^{+}}=Enc_{d}(d^{+})$. Assuming the dot product is used as the similarity metric, the above optimization goals can be described as the following objective,
$\min_{\theta}\max_{d^{-}\in\mathcal{N}(q)}\mathcal{E}(q,d^{+},d^{-};\theta)\text{, where }\mathcal{E}(q,d^{+},d^{-};\theta)=v_{q}\cdot v_{d^{-}}-v_{q}\cdot v_{d^{+}}$ (1)
which lowers the upper boundary of ranking similarity errors $\mathcal{E}$ [12]. The above objective can be effectively optimized by the pairwise/listwise contrastive loss [10], margin loss [61], etc. Following [29], listwise contrastive loss ($\ell^{CL}$) is used in our work,
$\ell^{CL}=-\log\frac{e^{v_{q}\cdot v_{d^{+}}/\tau}}{e^{v_{q}\cdot v_{d^{+}}/\tau}+\sum_{d^{-}\in\mathcal{N}(q)}e^{v_{q}\cdot v_{d^{-}}/\tau}}.$ (2)
, where $\tau$ is the temperature. The retrieval capacity is strongly correlated with parametrized encoders. As discussed before, serving deep parametrized LMs, especially giant LLMs, consumes large amounts of deployment resources. Thus, our work aims to address this issue by designing a novel asymmetric retrieval architecture with extreme imbalance.
Instructed retrieval
Similar to instructed fine-tuning (SFT) [55], modern LM-based retrievers [64, 76, 70], especially LLM retrievers [71, 5, 38], customize task specific instructions for better domain adaption abilities. Specifically, as shown in Figure 2, the task instruction is prepended to the queries before retrieval. Although [64, 76] use instructions on both query and passage sides, most LLM-based retrievers [71, 5, 38] do not use passage instructions to reuse the pre-built indexes across different tasks. In our work, dense retrieval follows the query instructions of Mistral-E5 [71] and does not involve passage instructions. The sparse retrieval is incompatible with instructions because it does not involve learnable query encoders. Detailed instructions are shown in the Appendix A.4.
2.2 Dense retrieval of LightRetriever
For LLM-based retrieval, LightRetriever pushes the query inference efficiency to a new extreme by completely removing the deep modeling on the query side. For dense retrieval, our work end-to-end trains cacheable token vectors for online serving.
Figure 3: Dense retrieval of LightRetriever.
Training
Given a task instruction $Inst$ and query tokens $T_{q}=\{t_{0},t_{1},...,t_{n-1}\}$, each query token $t_{i}$ ($i\in\{0,...,n-1\}$) are prepended with the instruction. Then, the corresponding query token vectors $v_{t_{i}}^{\text{den}}$ are obtained by encoding [^2] Specifically, EOS pooling is used in our work by following Mistral-E5 [71]. with the query encoder [^3] A customized causal mask is designed to avoid repeated prompt computations, detailed in Appendix A.5.,
$v_{t_{i}}^{\text{den}}=Enc_{q}(Inst;t_{i})$ (3)
Then, the dense query vector $v_{q}^{\text{den}}$ of the whole query text is aggregated by averaging token vectors,
$v_{q}^{\text{den}}=\frac{1}{n}\sum_{i=0}^{n-1}v_{t_{i}}^{\text{den}}$ (4)
Contrastive loss is the main objective for dense training, as defined in Equation 2.
Following [9, 32], an auxiliary KL loss is also used to align the asymmetric similarity scores $S^{student}=v_{q}^{den}\cdot v_{d}^{den}$ to full-sized symmetric similarity scores $S^{teacher}=Enc_{q}(q)\cdot Enc_{d}(d)$,
$\ell^{Align}=\texttt{KL-Div}(S^{student},S^{teacher})$ (5)
Full-sized symmetric baselines could be trained with asymmetric encoders together. Thus, the computation of auxiliary loss is economical. Experiments show that this helps further narrow the performance gaps, but it is not mandatory.
Caching
Query tokens do not deeply interact with each other. Thus, the query token vectors are cacheable. Given a certain vocabulary with size $V$, all token vectors $v_{t}^{den},t\in\{0,...,V-1\}$ are precomputed as in Equation 3 and stored in a lookup Embedding $E$.
Online serving
LLM-based encoders are no longer needed during online serving. Instead, only one Embedding layer, with a single weight matrix shaped $[V,H]$, is hosted in RAM. $H$ is the dimension of LLM’s hidden states, which is also the dimension of dense vectors. The dense query vector is obtained with a simple lookup-then-average operation,
$v_{q}^{\text{den}}=\frac{1}{n}\sum_{i=0}^{n-1}v_{t_{i}}^{den}=\frac{1}{n}\sum_{i=0}^{n-1}E[t_{i}]$ (6)
The online serving of our dense retriever is extremely lightweight, because such Embedding lookup is super efficient with or without GPU hardware.
2.3 Sparse retrieval of LightRetriever
The sparse retrieval of LightRetriever steps towards lightweight even further, by completely removing the need for the sparse query encoder.
Figure 4: Sparse retrieval of LightRetriever.
Training
The sparse query vector $v_{q}^{spr}$ directly maps a token and its corresponding quantity,
$v_{q}^{\text{spr}}[t]=\left\{\begin{array}[]{ll}0&\text{if }t\notin T_{q}\\ \text{Number of }t&\text{if }t\in T_{q}\end{array}\right.$ (7)
where $t\in\{0,...,V-1\}$ represents valid token ids within LLM’s vocabulary. We obtain the sparse document vector $v_{d}^{\text{spr}}$ by first feeding the document $d$ into the pretrained language model (LM), resulting in the final-layer hidden states $h_{\text{last}}=LM(d)$. These hidden states are then projected into the vocabulary space to get logits $w$ via the model’s own language modeling head, parameterized by a projection matrix $P$. Following SPLADE [20], to induce sparsity and suppress dominant term frequencies, we apply a ReLU activation followed by a log-saturation function, and then perform a max-pooling operation of the logits $w$ over the sequence length dimension,
$v_{d}^{\text{spr}}=\max(w)\text{, where }w=\text{logits}=\ln(\max(h_{\text{last}}\cdot P,0)+1)$ (8)
Here, $v_{d}^{\text{spr}}\in\mathbb{R}^{V}$ denotes the final sparse representation vector of the document. Given sparse vectors $v_{q}^{\text{spr}}$ and $v_{d}^{\text{spr}}$, the same contrastive loss in Equation 2 and auxiliary loss in Equation 5 are applied for representation learning.
$v_{d}^{\text{spr}}$ is not naturally sparse enough, because LM tends to generate dense distributions. Thus, the sparsify regulator is needed during the training phase. Following SPLADE [20], given a batch of documents $d=\{d_{0},...,d_{i}\},i<\text{batch size }N$, the FLOPs regulator [56] is used for sparsification. This regulator first averages the logits across different documents within a certain batch, then computes the squared sum to get the final regulator.
$\ell_{\text{FLOPS}}=\sum_{t=0}^{V-1}\left(\frac{1}{N}\sum_{i=0}^{N-1}w_{t}^{(d_{i})}\right)^{2}$ (9)
Online serving
The sparse query vector $v_{q}^{\text{spr}}$ during online serving follows the same rule in Equation 7. No more query encoder for sparse retrieval is needed.
3 Experiments
3.1 Settings
Datasets
A total of 23 training sets (including 20 English and 3 Chinese) are used in our experiments by following [47, 71]. Most of them are available at preprocessed data  [^4] https://huggingface.co/datasets/tdro-llm/finetune_data from tDRO [47] and Sentence Transformers Training Data  [^5] https://huggingface.co/datasets/sentence-transformers/embedding-training-data [59]. Detailed training sets and sampling ratios are listed in the Appendix A.3. BeIR [65] and CMTEB Retrieval [76] benchmarks are used in the evaluation. BeIR is a massive English retrieval benchmark collection with 15 evaluation sets. These heterogeneous datasets cover different domains, tasks, and retrieval granularities. CMTEB Retrieval is a massive Chinese retrieval benchmark collection with 8 sets, also covering a variety of domains. nDCG@10 is reported as the main metric by following [65, 76], accessing both recall and ranking abilities. Recalls (R@{20, 50, 100}) are also reported to assess retrieval ability at large windows.
Models & hyper-parameters
We modify the experiment code based on the implementation  [^6] https://github.com/ma787639046/tdro from tDRO [47]. All experiments are performed with a batch size of 128, 7 hard negatives, a max sequence length of 512, a contrastive temperature $\tau$ of 0.02, and 12k total steps. Following SPLADE [20], the coefficient of the FLOPs regulator is 0.001, and the regulator is quadratically increased to maximum at first 4k steps, which helps reduce the side effects of the regulator at initial steps [56]. LoRA [24] is used to save GPU memory usage during the LLM training, where LoRA r is 16, alpha is 32, and dropout is 0.1. Different LLMs with different sizes are tested as the backbone encoders to verify the generalization abilities, including Mistral-v0.3-7B [26], Llama-3.2-1B, 3B, Llama-3.1-8B [67], and Qwen-2.5-1.5B, 3B, 7B [2]. Direct linear interpolation [72] is used to combine dense and sparse similarity scores as the hybrid retrieval scores. All results are conducted on a DGX node with 8 NVIDIA H800 GPUs, 128 CPU cores of 2 Intel Xeon 8462Y+, and 2 TB of memory.
Baselines
Full-sized symmetric dense retrievers are our main baselines under the same training conditions, as shown in Figure 2. The query inference speeds and retrieval performance gaps are compared as the main results. Additionally, we also include some recent retrievers in the results, including Sentence Transformers static embedding [1], BM25 [60], BGE-m3 [11], E5-Mistral [71], and LLM2Vec [5]. However, they are not our main comparison baselines, because [1] and [60] are parameter-efficient but less performant. LLM-based retrievers [71, 5] are performant but inefficient.
Table 1: Comparison of query encoding time and speedup for full-sized baseline and LightRetriever. Tokenization time is also presented for reference. 65536 queries from MSMARCO [52] are processed. Evaluations are conducted on an NVIDIA DGX node using 1 H800 GPU with Mistral-v0.3-7B.
Tokenzation Baseline LightRetriever
GPU Encode No-GPU Encode GPU Encode
Batch Size Time (s) Time (s) Time (s) speedup Time (s) speedup
16 5.1168 69.9276 3.7065 19 0.1509 463
64 3.8995 42.2991 3.1719 13 0.0510 829
256 2.0713 45.8708 2.9990 15 0.0185 2480
1024 1.4226 57.4545 2.7969 21 0.0107 5370
4096 1.6798 74.7221 2.8218 26 0.0093 8035
8192 OOM 2.8529 0.0090
3.2 Main results
Encoding speeds
The encoding speed comparisons of full-sized baseline and LightRetriever are presented in Table 1. No-GPU and GPU settings are measured for LightRetriever. Because of the efficiency issue, the baseline is only measured on the GPU. As is shown in Table 1, it takes a full-sized Mistral-7B LLM up to 74s for encoding 65536 queries on an H800 GPU across different batch sizes. In comparison, our LightRetriever only consumes less than 0.1509s on the GPU, which achieves over 400-8000x speedup. Even when encoding without GPU acceleration, LightRetriever also achieves over 10-20x speedup compared to the GPU baseline.
Table 2: Retrieval performances of LightRetriever on BeIR and CMTEB Retrieval (CMTEB-R) benchmarks. The nDCG@10 gaps between LightRetriever hybrid and the corresponding baseline are presented. The best metrics of baselines and LightRetrievers are marked in bold.
Benchmark BEIR (15 datasets) CMTEB-R (8 datasets)
nDCG@10 R@20 R@50 R@100 nDCG@10 R@20 R@50 R@100
Static Embed 34.1 44.4 53.1 59.5 31.3 45.5 54.3 60.4
BM25 41.7 48.8 56.5 61.8 50.8 63.9 70.0 74.1
[…]
E5-Mistral 56.9 62.1 68.9 73.6 61.8 75.1 81.6 87.1
Llama3.2-1b
Baseline 53.1 60.2 66.9 72.0 63.8 78.1 84.1 87.9
Light Dense 48.7 55.8 63.0 68.3 58.7 73.1 79.7 83.8
Light Sparse 47.6 54.5 61.5 66.7 53.9 68.8 75.7 80.1
Light Hybrid 52.0-${}_{\text{-1.1}}$ 58.1 65.1 70.1 60.8-${}_{\text{-3.0}}$ 74.8 81.2 85.0
Llama3.2-3b
Baseline 55.6 62.9 69.7 74.7 66.1 81.0 86.6 90.0
Light Dense 51.0 58.6 65.5 70.5 60.4 75.1 81.5 85.3
Light Sparse 48.4 55.5 62.7 67.7 53.0 68.0 75.0 79.6
Light Hybrid 53.5-${}_{\text{-2.1}}$ 59.9 66.8 71.7 61.7-${}_{\text{-4.4}}$ 76.3 82.3 86.3
Llama3.1-8b
Baseline 56.8 64.3 71.1 75.7 67.6 82.0 87.4 90.9
Light Dense 52.0 59.5 66.5 71.4 61.4 75.6 82.3 86.4
Light Sparse 49.5 56.9 64.1 69.1 54.8 69.8 77.3 81.8
Light Hybrid 54.4-${}_{\text{-2.4}}$ 60.9 67.7 72.8 63.0-${}_{\text{-4.6}}$ 77.3 83.6 87.4
Mistral0.3-7b
Baseline 57.5 64.1 71.0 75.7 65.9 80.3 86.5 90.2
Light Dense 52.7 59.2 66.5 71.3 57.4 72.1 79.0 83.6
Light Sparse 48.0 56.1 63.3 68.3 50.7 65.5 72.7 77.8
Light Hybrid 54.5-${}_{\text{-3.0}}$ 60.1 67.4 72.1 59.4-${}_{\text{-6.5}}$ 73.9 80.6 84.8
Qwen2.5-1.5b
Baseline 54.0 60.5 67.5 72.3 65.9 81.0 86.7 90.4
Light Dense 48.9 56.0 63.1 68.1 61.9 77.0 83.2 87.4
Light Sparse 47.3 54.3 61.4 66.4 52.5 67.4 74.9 79.9
Light Hybrid 52.1-${}_{\text{-1.9}}$ 58.2 65.1 70.0 63.8-${}_{\text{-2.1}}$ 78.3 84.5 88.3
Qwen2.5-3b
Baseline 54.9 62.0 68.7 73.5 69.4 83.8 88.9 91.8
Light Dense 50.1 57.2 64.2 69.3 64.7 79.4 85.2 88.8
Light Sparse 47.5 54.9 62.0 67.2 54.1 69.2 76.5 81.1
Light Hybrid 52.8-${}_{\text{-2.1}}$ 59.1 66.0 70.9 65.7-${}_{\text{-3.7}}$ 80.2 86.1 89.3
Qwen2.5-7b
Baseline 56.6 63.6 70.5 75.2 70.1 84.4 89.4 92.4
Light Dense 50.6 58.6 65.6 70.5 65.0 79.8 85.8 89.3
Light Sparse 48.2 56.0 63.5 68.5 55.0 69.7 77.0 82.1
Light Hybrid 53.8-${}_{\text{-2.8}}$ 60.3 67.5 72.5 66.5-${}_{\text{-3.6}}$ 81.1 86.7 89.7
Retrieval performances
The retrieval performances are presented in Table 2. Experiments are conducted with 3 series of LLMs with different sizes. For each base model, we test the full-sized baseline and dense/sparse/hybrid (dense + sparse) retrieval performances of LightRetriever (Light). As is shown in the results, the performance gaps between LightRetriever (Hybrid) and its corresponding baseline are averaged around 5%, ranging from 2.1% (-1.1 point) to 9.9% (-6.5 point). Mistral [26] gets the highest nDCG@10 scores on the English BeIR benchmark, while Llama models [67] are better at recalls of BeIR. Qwen series [2] performs the best on the Chinese CMTEB Retrieval benchmark, attributed to better Chinese pre-training of Qwen.
Figure 5: Performances (Llama-3.1-8b) with different Top-k dimensions or sparsities.
3.3 Controls of vector dimension and sparsity
Retrieval systems usually involve indexing millions of documents into search databases, which consumes large amounts of disk space, especially for dense vectors. For example, indexing 8.8M MS-MARCO passages [52] with Llama-8b model consumes 134.92 GB for Faiss dense retrieval and 11.24 GB for Lucene sparse retrieval. Thus, it’s essential to explore the possibility of controlling index sizes by shrinking vectors. 1) For dense retrieval, we incorporate Matryoshka Representation Learning (MRL) [35] with LightRetriever to support flexible controls of dense vector dimensions. Given a dense vector $v^{den}$ with dimension $H$, MRL derives dimension-shrunk vectors by simply cutting out the Top-k dimensions $v^{den}[:k]$, then applies the same training loss as the original vectors. MRL trains a collection of Top-k dimensions together, enabling flexible choosing of a proper k. 2) For sparse retrieval, the sparse vector directly learns the term impact (i.e., lexical frequency) within a large sparse dimension. Thus, LightRetriever supports controlling sparsity by directly taking the Top-k terms, where the larger impact terms matter more.
We conduct experiments of controlling dense vector dimension and sparse vector sparsity with Top-k of $\{128,256,512,1024,2048,4096\}$ on Llama-3.1-8b. As presented in Figure 5, the performances of LightRetriever decrease marginally when shrinking Top-k to 4x ($4090\rightarrow 1024$). The lowest Top-k setting ($128$) has 7.9% and 8.8% performance gaps on BeIR and CMTEB-R benchmarks compared to the highest Top-k ($4096$), but its disk usages with MS-MARCO are only 4.22 GB for dense trieval and 3.03 GB for sparse retrieval. This shows that our LightRetriever is capable of flexible index controls via simple vector shrinking techniques, with limited effect on retrieval performances.
Table 3: Ablations (nDCG@10, w/Llama-3.1-8b) on BeIR and CMTEB-R about auxiliary KL loss.
Benchmark w/ KL wo/ KL
BeIR CMTEB-R BeIR CMTEB-R
Baseline 56.8 67.6 57.0+0.2 67.9+0.3
Light Dense 52.0 61.4 50.8-1.2 60.1-1.3
Light Sparse 49.5 54.8 48.7-0.8 54.4-0.4
Light Hybrid 54.4 63.0 53.9-0.5 62.1-0.9
3.4 Effect of auxiliary KL loss
Previous works for BERT-based asymmetric retrievers [73, 32] emphasize the importance of alignment between asymmetric students and full symmetric teachers. Our work also uses the KL loss as an auxiliary. Results in Table 3 show that removing such loss incurs limited drops of -0.5 and -0.9 on BeIR and CMTEB-R. Thus, such auxiliary KL loss is not mandatory in our experiment settings.
Table 4: Ablations of dense and sparse retrieval for LightRetriever (Llama-3.2-1b).
Dense Ablation BeIR CMTEB-R
LightRetriever 48.7 58.7
Symmetry
w/ Qry Full - Doc Light 8.7-40.0 9.4-49.3
w/ Qry Light - Doc Light 34.9-13.8 40.1-18.6
Other Ablations
w/ Direct Emb Layer 37.5-11.2 41.3-17.4
wo/ Instruction 47.1-1.6 56.1-2.6
Sparse Ablation BeIR CMTEB-R
LightRetriever 47.6 53.9
Symmetry
w/ Qry Full - Doc Light 16.2-31.4 10.8-43.1
Other Ablations
w/ Value fixed to 1 47.6+0.0 54.7+0.8
4 Ablations
4.1 Symmetry
Our work proposes an asymmetric architecture with an extremely lightweight query (Qry) encoder and a full-sized document (Doc) encoder. Besides, we also perform ablations on other symmetry settings, including lightweight on both query and document, and a full-sized query encoder with a lightweight passage encoder. Results in Table 4 show that a full-sized query encoder with a lightweight passage encoder leads to drastic performance degenerations, maybe caused by embedding collapses. Lightweight on both sides for dense retrieval works, but it does not perform excellently in retrieval tasks, just like Static Embedding [1]. Lightweight on both sides is not suitable for our sparse retrieval, because no learnable parameters are available.
4.2 Other ablations
Do we need a full query model during the dense training?
Dense retrieval of LightRetriever trains token vectors with full Transformers. Fast retrieval is achieved by caching learned dense token vectors and lookups during serving. If we directly use the Embedding Layer within the Transformers, the performances drop 11.2 and 17.4 points on BeIR and CMTEB-R. This verifies that during the training phase, a full-sized model is still needed for learning query token vectors.
Task instructions
Task instructions contribute +1.6 and +2.6 points on BeIR and CMTEB-R of dense retrieval. Similar effects of instructions are also observed in previous LLM retrievers [71, 5].
Value of sparse query vector
The sparse query vector is a key-value type of mapping. In Equation 7, we model the value of sparse query vectors by counting the corresponding token ids that appear in the query. For ablations, we choose a simpler way of modeling by stopping the counting and fixing the value to 1. In Table 4, BeIR results do not change. However, CMTEB-R results improve by +0.8 points. This indicates that counting term frequency in our sparse query vector seems unimportant to some extent, under large amounts of training on LightRetriever. The co-occurrence is more important.
5 Conclusion
Existing LLM-based retrievers use symmetric dual-encoders to model both queries and documents. While documents could be pre-encoded and indexed offline, queries arrive in real-time and need online encoding. The deployment of large LLMs results in inefficient online services. In this paper, we propose LightRetriever, a novel LLM-based hybrid retrieval architecture capable of extremely lightweight query encoding. Our approach achieves over a 1000x speedup for query inference on an H800 GPU, even a 20x speedup without GPU acceleration. Experiments show great robustness and generalization abilities across different foundation LLMs and retrieval tasks.
Appendix A Technical Appendices and Supplementary Material
A.1 Limitations
Our work focuses on online efficiency optimizations of recent LLM-based architecture, whose giant parameters pose a greater need for such optimizations. Other readers may be interested in the application of our methods on non-LLM architectures, e.g., BERT. While such a combination is possible, due to limited time constraints and lesser needs for non-LLM, we leave this to future work.
A.2 Related works
A.2.1 Text retrieval
Text retrieval builds on top of the vector space model [62], which encodes query and document representation vectors, and performs relevance search based on vector similarities.
Dense retrieval
Dense retrieval [29] trains PLM-based encoders to condense queries and passages into low-dimensional dense vectors, then performs relevance search based on the Maximum Inner Product Search (MIPS) [51] algorithm. In recent years, dense retrievers, such as BGE [11] and E5 [70], have gained popularity for strong retrieval abilities across different tasks [50, 65], languages [83], and context granulaties [74, 52], owing to diverse training data [59], retrieval-customized pre-training [21, 75, 46], improved negative-mining [78], and so on. Recent SOTA retrievers [71, 5, 38] start to utilize LLMs [26, 67] as backbone encoders. With larger parameters and pre-training data, LLM-based retrievers enable significantly powerful retrieval capacities and domain adaptation abilities. However, LLMs also incur heavy online workloads when serving as query encoders. Improving the serving efficiency of LLM-based retrievers is still left for exploration.
Sparse retrieval
Traditional statistical sparse retrieval algorithms (e.g., TF-IDF [63], and BM25 [60]) do not involve learnable language models. Instead, they directly analyze lexical frequencies (e.g., TF) and importance (e.g., IDF) in a large vocabulary space. These lexicals are sparsely distributed and require building inverted indexes [45] for feasible similarity metric calculation, such as Apache Lucene [6] and Tantivy [58]. In recent years, sparse retrieval has started to work with pre-trained LMs (PLMs). BGE [11], SPLADE [20], SparTerm [3], and SPARTA [84] use customized XLM-RoBERTa [14] or BERT [18] as backbone encoders to learn term frequencies in an end-to-end manner and directly perform impact search based on the dot product of sparse vectors. Because of sparsity, these works also search with inverted indexes. Recent works [85] also try to incorporate LLMs with sparse retrieval. However, the online efficiency issue still exists.
Hybrid retreval
Hybrid retrieval is a technique that combines retrieval scores [72] or ranks [15] from multiple retrieval systems, such as dense [29] / sparse [60, 20] retrieval, and single [29] / multiple [31, 11] vector retrieval. It enables better retrieval abilities by interpolating results from multiple sources. Linear interpolation [72] of dense and sparse scores is used in our work.
A.2.2 Improving inference efficiency
Numerous efforts endeavor to improve the inference efficiency of text retrieval.
Distilling smaller encoders
Previous works have tried to distill small query encoders from large encoders for better query inference efficiency. KALE [9] prunes the BERT-based query encoder and distills it during post-training. However, when the query encoder is pruned to one Transformers layer, its recall@20 on NQ [36] decreases by 25-30%. [73] proposes asymmetric dual-encoders by pruning BERT layers or using smaller BERT models. Its experiments show that a 2-layer BERT query encoder retains 92.5% of the full-sized performances via distillation and proper student initialization. USTAD [32] also proposes similar asymmetry and distillation strategies, which choose the 12-layer BERT as teachers and the 6-layer DistilBERT or 2-layer BERT-mini as students. Its experiments show that the distilled BERT-mini query encoder achieves 95-97% of the teacher’s performance because of better distillation.
However, these studies have primarily focused on small language models (SLMs), such as BERT [18], and overlooked feasible approaches for improving online inference efficiency of LLM-based retrievers, where such optimization is even more critical. Furthermore, these existing methods typically retain Transformer layers in the query encoders, which limits their efficiency gains. In addition, most of these works rely on SLMs and are trained on relatively narrow datasets, such as MS-MARCO [52], thus unable to compete with LLM-based retrievers.
In contrast, with the rapid advancement of LLM-based text retrievers, driven by larger and more diverse foundation models and datasets, retrieval performance has significantly improved. To address the growing need for efficient online inference in this setting, we propose a novel method that, to our knowledge, is the first to target the online efficiency challenge in LLM-based retrievers. Our approach eliminates the need for Transformers layers for query encoders and does not rely on mandatory distillation. Yet our work only incurs an average performance drop of about 5%. Moreover, we evaluate our method across multiple LLMs of varying sizes and diverse retrieval tasks, demonstrating strong generalization capabilities.
Shrinking index sizes
Previous works have also been dedicated to reducing index sizes for smaller disk space usage. Matryoshka representation learning (MRL) [35] supports adjusting the dense vector dimensions by directly taking the top-k dimension of the original vector as shrink vectors. It aligns and trains the corresponding multi-dimensional vectors in the end-to-end contrastive learning. For example, if $k\in\{64,128,256\}$, query-passage dense vectors with $\{64,128,256\}$-dims could be obtained together after one encoder forward operation. For LM-based sparse retrieval, the similarity scores are directly affected by the impacts of corresponding lexicals [20]. Thus, it naturally supports reducing the index sizes by retaining Top-k terms/lexicals. To explore the potential of controlling index sizes, our work also explores shrinking index sizes with dense MRL and sparse Top-k in the Experiments Section 3.3, which also shows strong generalization capacities.
A.3 Training datasets
Dataset infos
Following the settings of previous LLM-based retrievers [71, 5, 47], our work uses 23 training sets pre-built by tDRO [47] or Sentence Transformers Training Data [59], including 20 English and 3 Chinese datasets. Detailed dataset information is listed in Table 5 as follows.
Table 5: Training Dataset informations.
Dataset Language Category Deduped Size Epoch Ratio
Amazon Review (2018) [53] English Amazon 999999 0.01 0.58%
cMedQA2 [81] Chinese Chinese Medical 99936 1 5.84%
DuReader [57] Chinese Chinese Web Collections 86366 1 5.05%
Eli5 [19] English Reddit 325390 0.05 0.95%
Fever [66] English Wikipedia QA 109808 0.8 5.13%
FiQA [48] English Financial 5498 1 0.32%
GooQA Pairs [30] English Web Collections 3012347 0.025 4.40%
HotpotQA [80] English Wikipedia QA 85000 0.5 2.48%
MSMARCO [52] English Web Collections 502854 1 29.39%
NFCorpus [7] English Medical 2585 1 0.15%
NPR [44] English News 594376 0.05 1.74%
NQ [36] English Wikipedia QA 58800 1 3.44%
PQA Pairs [40] English Wikipedia QA 99999 0.5 2.92%
Quora Duplicates Triples [25] English Forum Duplicates 97011 0.5 2.83%
S2ORC Title-Abstract [42] English Semantic Scholar 99998 0.5 2.92%
SciFact [69] English S2ORC 806 1 0.05%
SPECTER [13] English Semantic Scholar 136642 0.25 2.00%
StackExchange Duplicates (Title-Body) [16] English Forum Duplicates 250516 0.25 3.66%
E5 Synthetic [4] English GPT Synthetic Data 224791 1 13.14%
T2Ranking [77] Chinese Chinese Web Collections 200362 0.5 5.86%
Trivia [28] English Wikipedia QA 60370 0.5 1.76%
Wikihow [33] English WikiHow 128543 0.25 1.88%
Yahoo Answers (Title-Answer) [82] English Yahoo 1198018 0.05 3.50%
Many downstream retrieval benchmarks [65, 76] are evaluated out-of-domain, where no training samples from the downstream sources are available. To avoid the overfitting of training sets, as is shown in the above Table 5, we slightly tune the epoch used in the dataset sampling, by following previous works in [54, 71, 47].
Training set decontamination
Training set decontamination is essential for developing fair, unbiased, and robust retrieval systems. Existing training datasets are derived from multiple sources, whose training data have possible duplicates with the test sets. Following tDRO [47], we perform strict training set decontamination with SimHash [49] to ensure no training queries appear in the BeIR [65] and CMTEB Retrieval [76]. The original sizes, deduplicated sizes, and duplicate numbers as listed in the Table 6.
Table 6: Training set deduplications. We report each dataset with the original size, deduplicated size, and number of duplicates detected in both the train and test sets.
Dataset Original Size Deduped Size Duplicates
Amazon Review (2018) 1000000 999999 1
cMedQA2 100000 99936 64
DuReader 86395 86366 29
Eli5 325475 325390 85
Fever 109810 109808 2
FiQA 5498 5498 0
GooQA Pairs 3012496 3012347 149
HotpotQA 85000 85000 0
MSMARCO 502939 502854 85
NFCorpus 2590 2585 5
NPR 594384 594376 8
NQ 58812 58800 12
PQA Pairs 100000 99999 1
Quora Duplicates Triples 101762 97011 4751
S2ORC Title-Abstract 100000 99998 2
SciFact 809 806 3
SPECTER 136645 136642 3
StackExchange Duplicates (Title-Body) 250519 250516 3
E5 Synthetic 224791 224791 0
T2Ranking 200376 200362 14
Trivia 60380 60370 10
Wikihow 128543 128543 0
Yahoo Answers (Title-Answer) 1198260 1198018 242
Additionally, we also listed some samples of the duplicated training queries, as well as the corresponding test sets in Table 7.
Table 7: Samples of duplicate queries found in both the train and test sets.
Train Set Train qid Train Query Test Set Test qid Test Query
Quora train-164 Is there a framework for auditing social media? Quora 201573 Is there a framework for auditing social media?
Yahoo-Answer train-10046 who is sachin tendulkar? Quora 47702 Who is Sachin Tendulkar?
MSMARCO 968274 where did abraham lincoln died DBPedia QALD2_tr-6 Where did Abraham Lincoln die?
Eli5 train-3088 How does unemployment insurance work? FiQA 2648 How does unemployment insurance work?
A.4 Task instructions
As previously mentioned in Section 2.1, modern LLM-based retrievers [71, 5] introduce task-specific instructions to format input queries, which enables higher retrieval capacities. These LLM-based instruction-tuned retrieval is similar to instruction-based supervised fine-tuning (SFT) in LLM generation [55]. Following Mistral-E5 [71], we reuse most of its task instructions, and format the instruction and query in the following format:
[…]
The detailed instructions are listed below, most of which are directly copied from the previous works E5-Mistral [71] and tDRO [47] for reproducibility. Note that one training set may correspond to multiple instructions to avoid overfitting.
Instructions for training sets
  1. 1.
    Amazon Review (2018)(1): Given a title, retrieve the corresponding reviews from Amazon
  2. 2.
    Amazon Review (2018)(2): Given a title, retrieve a Amazon review
  3. 3.
    cMedQA2: Given a Chinese community medical question, retrieve replies that best answer the question
  4. 4.
    DuReader: Given a Chinese search query, retrieve web passages that answer the question
  5. 5.
    Eli5: Provided a user question, retrieve the highest voted answers on Reddit ELI5 forum
  6. 6.
    GooQA Pairs: Given a web search query, retrieve the corresponding answers from Google
  7. 7.
    HotpotQA: Given a multi-hop question, retrieve documents that can help answer the question
  8. 8.
    MSMARCO: Given a web search query, retrieve relevant passages that answer the query
  9. 9.
    NQ(1): Given a question, retrieve Wikipedia passages that answer the question
  10. 10.
    NQ(2): Retrieve Wikipedia passages that answer the question
  11. 11.
    Quora Duplicates Triples(1): Given a question, retrieve questions that are semantically equivalent to the given question
  12. 12.
    Quora Duplicates Triples(2): Find questions that have the same meaning as the input question
  13. 13.
    S2ORC Title-Abstract(1): Given a title, retrieve the abstract from scientific papers
  14. 14.
    S2ORC Title-Abstract(2): Given a title, retrieve abstracts from scientific papers that match the title
  15. 15.
    SciFact: Given a scientific claim, judge whether the document supports or refutes the claim
  16. 16.
    SPECTER(1): Given a title, retrieve semantic related titles
  17. 17.
    SPECTER(2): Retrieve semantic related titles from scientific publications
  18. 18.
    StackExchange Duplicates (Title-Body): Retrieve duplicate questions and passages from StackOverflow forum
  19. 19.
    T2Ranking: Given a Chinese search query, retrieve web passages that answer the question
  20. 20.
    Trivia(1): Given a question, retrieve Wikipedia passages that answer the question
  21. 21.
    Trivia(2): Retrieve Wikipedia passages that answer the question
  22. 22.
    Wikihow: Given a summary, retrieve Wikipedia passages that match the summary
  23. 23.
    Yahoo Answers (Title-Answer): Given a title, retrieve Yahoo answers that match the title
Instruction for English BeIR benchmarks
  1. 1.
    ArguAna: Given a claim, find documents that refute the claim
  2. 2.
    ClimateFEVER: Given a claim about climate change, retrieve documents that support or refute the claim
  3. 3.
    CQADupStack: Given a question, retrieve detailed question descriptions from Stackexchange that are duplicates to the given question
  4. 4.
    DBPedia: Given a query, retrieve relevant entity descriptions from DBPedia
  5. 5.
    FEVER: Given a claim, retrieve documents that support or refute the claim
  6. 6.
    FiQA2018: Given a financial question, retrieve user replies that best answer the question
  7. 7.
    HotpotQA: Given a multi-hop question, retrieve documents that can help answer the question
  8. 8.
    MSMARCO: Given a web search query, retrieve relevant passages that answer the query
  9. 9.
    NFCorpus: Given a question, retrieve relevant documents that best answer the question
  10. 10.
    NQ: Given a question, retrieve Wikipedia passages that answer the question
  11. 11.
    Quora: Given a question, retrieve questions that are semantically equivalent to the given question
  12. 12.
    SCIDOCS: Given a scientific paper title, retrieve paper abstracts that are cited by the given paper
  13. 13.
    SciFact: Given a scientific claim, retrieve documents that support or refute the claim
  14. 14.
    Touche2020: Given a question, retrieve detailed and persuasive arguments that answer the question
  15. 15.
    TRECCOVID: Given a query on COVID-19, retrieve documents that answer the query
Instruction for Chinese CMTEB Retrieval benchmarks
  1. 1.
    CmedqaRetrieval: Given a Chinese community medical question, retrieve replies that best answer the question
  2. 2.
    CovidRetrieval: Given a question on COVID-19, retrieve news articles that answer the question
  3. 3.
    DuRetrieval: Given a Chinese search query, retrieve web passages that answer the question
  4. 4.
    EcomRetrieval: Given a user query from an e-commerce website, retrieve description sentences of relevant products
  5. 5.
    MMarcoRetrieval: Given a web search query, retrieve relevant passages that answer the query
  6. 6.
    MedicalRetrieval: Given a medical question, retrieve user replies that best answer the question
  7. 7.
    T2Retrieval: Given a Chinese search query, retrieve web passages that answer the question
  8. 8.
    VideoRetrieval: Given a video search query, retrieve the titles of relevant videos
A.5 Efficiency optimizations techniques
Customized causal mask
As defined in the Equation 3, our work computes dense token vectors for each input token with a common prompt. This causes the repeated computation of the common prompt area. We develop a customized causal mask for LLMs to avoid such computation waste. Specifically, given a task instruction $Inst$ and query tokens $T_{q}=\{t_{0},t_{1},...,t_{n-1}\}$, we format the instruction and query tokens as follows,
[…]
Figure 6: An example of the customized attention mask.
The customized causal mask ensures that the token blocks can attend to the common prompt area, but not attend to each other. Let the sequence length be $L$, prompt length be $P$, micro block (each token + <eos>) width be $w$, and let $q\in[0,L-1]$ denote the query index, $k\in[0,L-1]$ the key index. The attention mask (0 means attend, $\infty$- ∞ means masked) is defined as,
$\text{Mask}[q,k]=\begin{cases}0,&\text{if }q<P\text{ and }k\leq q\quad\text{(attn within prompt)}\\ 0,&\text{if }q\geq P\text{ and }k<P\quad\text{(blocks can attend to prompt)}\\ 0,&\text{if }q\geq P\text{ and }\left\lfloor\frac{q-P}{w}\right\rfloor=\left\lfloor\frac{k-P}{w}\right\rfloor\text{ and }k\leq q\quad\text{(attn within current block)}\\ -\infty,&\text{otherwise}\end{cases}$
An example of the customized causal mask is shown in Figure 6.
Optimizations of sparse max aggregation
As stated in Equation 8, the sparse vector is computed via
[…]
, where the vocabulary dimension $V$ is much larger than the hidden dimension $H$. For example, for Llama3.1-8b [67] model, $V=128256$ and $H=4096$, where $\lfloor V/H\rfloor=31$. This means that, given batch size B, sequence length T, hidden size H, vocabulary size V, the intermediate tensors with shape $[B,T,V]$ in the projection, ReLU, and Log are 31x larger than a dense tensor $[B,T,H]$, which consumes huge amounts of GPU memory during training and inference. However, only a tensor with shape $[B,V]$ is needed for the final aggregated representations. Thus, an optimization for computing the sparse vector is needed urgently.
We design an efficient way of computing the sparse vector. Firstly, because ReLU and Log are monotonically increasing functions, we first move the last max aggregation operations to step 2, right after the projection.
[…]
Then we design a customized PyTorch Autograd function to fuse steps 1&2. It computes the projection and maximum operations together along the sequence length dimension $T$. Such sliced computation avoids the creation of a large tensor with a sequence length dimension. Our preliminary observation shows that this helps save the GPU memory usage of the sparse max aggregation up to 100-200x. The forward and backward are presented in the Algorithm 1.
Algorithm 1 Forward and Backward Pass of Efficient Projection then Max aggregation
1:Input tensor $X\in\mathbb{R}^{B\times T\times H}$, weight $W\in\mathbb{R}^{H\times V}$, optional bias $b\in\mathbb{R}^{V}$, optional input attention mask $M\in\{0,1\}^{B\times T}$, where B is batch size, T is sequence length, H is hidden size, V is vocabulary size.
2:Output tensor $Y\in\mathbb{R}^{B\times V}$
3:function Forward($X,W,b,M$)
4:     Initialize $Y\leftarrow-\infty$
5:     for $t=0$ to $T-1$ do
6:         $L_{t}\leftarrow X[:,t,:]\cdot W$
7:         if $b$ is not None then
8:              $L_{t}\leftarrow L_{t}+b$
9:         end if
10:         if $M$ is not None then
11:              for $i=0$ to $B-1$ do
12:                  if $M[i,t]=0$ then
13:                       $L_{t}[i,:]\leftarrow-\infty$
14:                  end if
15:              end for
16:         end if
17:         $Y\leftarrow\max(Y,L_{t})$
18:     end for
19:     Save $\text{argmax}_{t}$ indices as max_indices
20:     return $Y$
21:end function
22:function Backward($\nabla Y,\texttt{max\_indices},X,W$)
23:     Initialize $\nabla X\leftarrow 0$, $\nabla W\leftarrow 0$, $\nabla b\leftarrow 0$
24:     for $t=0$ to $T{-}1$ do
25:         $X_{t}\leftarrow X[:,t,:]\in\mathbb{R}^{B\times H}$
26:         $\nabla X_{t}\leftarrow\nabla X[:,t,:]\in\mathbb{R}^{B\times H}$
27:         $M_{t}\leftarrow\texttt{max\_indices}_{I==t}\in\{0,1\}^{B\times V}$
28:         $\nabla L_{t}\leftarrow\nabla Y\odot M_{t}\in\mathbb{R}^{B\times V}$triangleright Only keep grads where $t$ was max
29:         $\nabla X_{t}\mathrel{+}=\nabla L_{t}\cdot W^{\top}$
30:         $\nabla W\mathrel{+}=X_{t}^{\top}\cdot\nabla L_{t}$
31:         $\nabla b\mathrel{+}=\sum_{i=1}^{B}\nabla L_{t}[i,:]$
32:     end for
33:     return $\nabla X,\nabla W,\nabla b$
34:end function
Other important optimizations
Proper optimizations are important. For example, we found it hard to reproduce quickly with some of the open-sourced baselines [5], which is roughly 6x slower than our implementation when testing with the same size of LLM. Our experiments involve training with 23 train sets and inference with another 23 test sets, which requires proper optimizations urgently. We import other optimizations based on the implementations  [^7] Please refer to [47]’s implementation on https://github.com/ma787639046/tdro. released by tDRO [47]. These optimizations include 1) A distributed multi-node inference framework based on PyTorch RPC [41]; 2) Sequence packing [34] with Flash Attention 2 [17], which gives roughly 50% training speedup due to the removal of the pad tokens.; 3) Fused Kernels from Liger-Kernel [23], which gives roughly 10% speedup to both training and inferencing; 4) Multi-vector gradient accumulation support with GradCache [22] for optionally enlarging batch sizes if necassary.
A.6 Detailed main results
The detailed main results (nDCG@10) on each dataset are shown in the tables below, which are full versions of Table 2.
Table 8: (Part1) Detailed main results (nDCG@10) on BeIR test sets (except MSMARCO, it uses the dev set).
Task / Test, nDCG@10 ArguAna CQADup CFEVER DBPedia FEVER FiQA HotpotQA MSMARCO
Static Embedding 44.4 22.0 20.4 27.4 43.0 20.0 46.8 17.9
BM25 31.5 29.9 21.3 31.3 75.3 23.6 60.3 22.8
BGE-m3-dense 53.9 38.4 26.5 38.5 77.2 40.8 68.6 37.6
BGE-m3-sparse 36.2 30.5 26.7 24.4 85.9 26.9 70.3 17.1
LLM2Vec-llama3-8b 62.8 48.3 34.3 48.3 90.2 55.3 71.8 43.2
E5-Mistral 61.9 42.9 38.3 48.9 87.8 56.6 75.7 43.1
Llama3.2-1b
Baseline 58.0 44.0 41.2 43.2 89.2 45.5 69.3 42.7
Light Dense 49.2 38.1 37.9 34.2 83.6 40.4 60.8 39.2
Light Sparse 49.5 34.5 27.5 35.9 84.7 35.7 65.5 37.2
Light Hybrid 54.1 39.7 39.3 39.5 87.3 42.1 68.2 41.1
Llama3.2-3b
Baseline 58.5 46.8 43.7 46.4 89.9 53.4 73.7 44.3
Light Dense 55.3 40.3 41.7 36.5 84.0 46.2 64.3 41.6
Light Sparse 46.5 34.1 26.9 38.5 86.2 38.9 66.3 39.2
Light Hybrid 56.3 41.4 42.4 41.6 87.6 46.9 69.7 42.9
Llama3.1-8b
Baseline 57.1 49.2 44.1 48.2 90.9 59.0 77.8 46.2
Light Dense 55.8 40.9 41.5 38.7 85.7 49.8 67.6 43.4
Light Sparse 46.4 33.6 32.2 38.8 86.1 41.9 67.9 41.0
Light Hybrid 57.2 41.9 42.4 42.0 87.8 50.7 71.5 44.5
Mistral0.3-7b
Baseline 56.2 49.7 43.6 48.4 91.3 59.6 77.7 46.6
Light Dense 53.5 41.6 40.9 38.9 85.6 50.9 66.9 44.2
Light Sparse 44.0 33.6 21.9 37.0 81.0 43.6 65.4 40.9
Light Hybrid 54.3 42.2 41.4 40.5 86.9 51.7 69.9 44.8
Qwen2.5-1.5b
Baseline 56.6 44.1 40.4 43.9 88.9 47.9 68.1 42.6
Light Dense 51.2 38.3 36.1 33.0 82.1 41.2 59.0 39.1
Light Sparse 46.5 35.0 25.7 36.3 85.1 35.3 63.3 37.3
Light Hybrid 54.1 40.1 37.7 39.5 87.0 42.1 66.5 40.8
Qwen2.5-3b
Baseline 55.5 47.0 44.1 45.8 89.6 52.9 71.6 43.7
Light Dense 49.8 39.6 41.1 35.0 83.9 45.4 62.8 41.0
Light Sparse 45.5 36.2 28.7 36.9 84.9 38.6 64.2 38.7
Light Hybrid 52.8 41.4 42.1 39.8 86.9 46.1 68.2 42.3
Qwen2.5-7b
Baseline 57.1 48.2 43.5 47.9 90.7 56.5 74.7 44.8
Light Dense 53.3 40.0 39.1 36.8 85.1 48.8 64.4 41.8
Light Sparse 47.5 35.6 28.2 36.7 85.4 40.2 65.2 40.1
Light Hybrid 54.4 41.6 42.2 41.0 87.9 49.7 69.4 43.3
Table 9: (Part2) Detailed main results (nDCG@10) on BeIR test sets.
Task / Test, nDCG@10 NFCorpus NQ Quora SCIDOCS SciFact TRECCOVID Touche2020 Avg (# 15)
Static Embedding 30.0 23.1 77.5 13.2 59.3 44.6 22.4 34.1
BM25 32.5 32.9 78.9 15.8 66.5 65.6 36.7 41.7
BGE-m3-dense 31.5 60.1 88.4 15.2 63.6 47.0 20.6 47.2
BGE-m3-sparse 28.3 20.3 73.4 12.2 63.3 52.7 27.8 39.7
LLM2Vec-llama3-8b 41.8 64.2 87.2 23.0 78.2 80.3 20.5 56.6
E5-Mistral 38.6 63.5 89.6 16.3 76.4 87.2 26.4 56.9
Llama3.2-1b
Baseline 35.5 58.6 89.4 19.8 72.3 66.7 20.9 53.1
Light Dense 30.1 52.2 87.0 17.2 66.5 68.8 24.7 48.7
Light Sparse 34.0 52.0 82.2 17.2 69.1 60.6 27.8 47.6
Light Hybrid 34.2 56.6 87.4 18.6 71.5 70.5 30.2 52.0
Llama3.2-3b
Baseline 37.6 62.8 89.5 22.5 75.3 68.7 21.2 55.6
Light Dense 31.7 56.4 87.8 19.0 70.4 68.6 21.6 51.0
Light Sparse 33.3 55.3 83.2 17.7 67.3 64.1 28.5 48.4
Light Hybrid 34.2 59.8 88.1 20.0 71.4 70.9 29.5 53.5
Llama3.1-8b
Baseline 38.6 65.5 89.9 23.8 77.3 62.0 22.2 56.8
Light Dense 31.8 59.0 88.5 19.3 70.6 65.7 21.1 52.0
Light Sparse 33.6 58.4 84.4 18.9 69.7 63.0 27.1 49.5
Light Hybrid 34.6 61.9 88.6 20.7 72.6 71.5 27.8 54.4
Mistral0.3-7b
Baseline 36.8 66.9 90.0 22.7 77.7 71.3 23.6 57.5
Light Dense 31.9 59.8 88.7 18.5 70.3 71.5 26.7 52.7
Light Sparse 33.1 57.9 84.6 18.1 66.7 61.1 31.7 48.0
Light Hybrid 33.8 62.5 88.8 19.8 71.6 76.7 32.4 54.5
Qwen2.5-1.5b
Baseline 36.4 57.1 89.0 20.4 72.7 73.0 28.4 54.0
Light Dense 29.8 50.9 86.7 18.1 66.5 72.6 29.1 48.9
Light Sparse 32.9 51.6 81.3 17.4 66.4 64.8 31.1 47.3
Light Hybrid 34.0 55.8 87.0 19.3 69.2 76.8 32.3 52.1
Qwen2.5-3b
Baseline 35.9 61.4 89.5 22.3 72.9 65.7 26.0 54.9
Light Dense 30.6 54.5 87.5 18.5 64.8 71.2 26.1 50.1
Light Sparse 33.2 54.2 82.2 17.6 67.6 57.0 26.4 47.5
Light Hybrid 34.2 58.5 87.8 19.5 69.1 74.2 29.8 52.8
Qwen2.5-7b
Baseline 37.2 63.7 89.7 23.7 77.2 69.1 25.3 56.6
Light Dense 30.7 55.9 87.6 19.5 67.9 64.1 24.1 50.6
Light Sparse 33.9 55.5 78.7 18.3 69.0 60.5 28.5 48.2
Light Hybrid 34.4 60.1 87.9 21.1 71.1 73.1 29.4 53.8
Table 10: Detailed main results (nDCG@10) on CMTEB Retrieval dev sets.
Task / Dev, nDCG@10 Cmedqa Covid DuReader Ecom MMarco Medical T2 Video Avg (# 8)
Static Embedding 7.5 45.7 37.9 33.9 34.9 15.7 35.8 38.9 31.3
BM25 13.7 86.6 57.1 45.1 48.3 32.1 60.5 62.7 50.8
BGE-m3-dense 31.0 77.2 82.9 57.5 76.9 51.5 80.8 54.4 64.0
BGE-m3-sparse 24.5 76.0 71.4 50.3 59.2 44.0 71.7 58.5 57.0
LLM2Vec-llama3-8b 35.2 16.5 80.7 54.4 76.5 54.6 65.2 52.4 54.4
E5-Mistral 34.2 73.1 87.0 46.0 74.8 52.8 80.7 45.4 61.8
Llama3.2-1b
Baseline 32.3 74.2 84.6 53.8 73.5 50.5 78.0 63.1 63.8
Light Dense 27.2 67.6 80.8 51.0 67.9 43.0 73.4 58.5 58.7
Light Sparse 17.3 64.3 74.4 51.0 63.3 39.4 69.3 52.2 53.9
Light Hybrid 27.2 69.7 81.8 55.5 69.8 44.1 75.9 62.3 60.8
Llama3.2-3b
Baseline 36.8 75.6 87.9 57.7 75.7 53.8 81.1 60.5 66.1
Light Dense 29.8 70.2 84.0 51.7 69.2 46.5 75.7 56.4 60.4
Light Sparse 17.2 65.4 76.1 47.4 65.0 42.5 68.5 42.0 53.0
Light Hybrid 29.9 71.5 84.8 54.5 71.2 47.2 77.1 57.7 61.7
Llama3.1-8b
Baseline 39.5 77.1 89.5 58.9 78.1 57.6 84.1 56.2 67.6
Light Dense 32.1 71.0 85.7 53.0 70.2 49.0 78.5 52.0 61.4
Light Sparse 15.6 70.4 82.9 47.3 63.9 42.2 75.1 41.3 54.8
Light Hybrid 32.1 72.8 86.6 55.3 71.6 49.7 80.1 55.8 63.0
Mistral0.3-7b
Baseline 38.5 74.1 88.6 54.7 75.6 56.6 82.3 56.5 65.9
Light Dense 30.4 64.4 83.6 47.0 64.3 46.3 74.7 48.5 57.4
Light Sparse 11.5 63.2 80.5 43.4 61.7 36.7 71.3 37.4 50.7
Light Hybrid 30.3 67.2 84.5 50.7 67.6 46.9 76.4 51.4 59.4
Qwen2.5-1.5b
Baseline 37.9 78.2 87.6 56.9 75.9 54.7 78.8 57.0 65.9
Light Dense 31.3 73.8 85.3 52.8 71.8 49.5 75.2 55.7 61.9
Light Sparse 14.5 66.7 78.1 49.3 61.6 39.1 63.5 47.0 52.5
Light Hybrid 31.2 74.4 86.0 56.5 73.4 50.5 76.4 61.6 63.8
Qwen2.5-3b
Baseline 39.4 80.1 89.5 61.9 77.8 58.7 84.4 63.6 69.4
Light Dense 32.7 76.0 86.8 56.7 73.2 51.6 80.9 59.5 64.7
Light Sparse 12.8 70.4 80.2 48.8 61.1 39.4 72.4 47.9 54.1
Light Hybrid 32.6 77.0 87.6 58.3 74.0 51.6 82.0 62.8 65.7
Qwen2.5-7b
Baseline 41.3 81.4 91.0 61.1 80.0 59.8 85.5 60.3 70.1
Light Dense 34.3 74.6 88.0 55.7 75.4 53.2 81.6 57.2 65.0
Light Sparse 12.5 72.5 82.4 48.3 65.2 40.4 73.3 45.2 55.0
Light Hybrid 34.0 76.7 88.6 58.0 76.2 53.9 83.0 61.4 66.5

### v3

@@
 Abstract
-Large Language Models (LLMs)-based hybrid retrieval uses LLMs to encode queries and documents into low-dimensional dense or high-dimensional sparse vectors.
-It retrieves documents relevant to search queries based on vector similarities.
+Large Language Models (LLMs)-based text retrieval retrieves documents relevant to search queries based on vector similarities.
 Documents are pre-encoded offline, while queries arrive in real-time, necessitating an efficient online query encoder.
 Although LLMs significantly enhance retrieval capabilities, serving deeply parameterized LLMs slows down query inference throughput and increases demands for online deployment resources.
-In this paper, we propose LightRetriever, a novel LLM-based hybrid retriever with extremely lightweight query encoders.
+In this paper, we propose LightRetriever, a novel LLM-based retriever with extremely lightweight query encoders.
 Our method retains a full-sized LLM for document encoding, but reduces the workload of query encoding to no more than an embedding lookup.
-Compared to serving a full-sized LLM on an H800 GPU, our approach achieves over a 1000× speedup for query inference with GPU acceleration, and even a 20× speedup without GPU.
-Experiments on large-scale retrieval benchmarks demonstrate that our method generalizes well across diverse retrieval tasks, retaining an average of 95% full-sized performance.
+Compared to serving a full LLM on an A800 GPU, our method achieves over 1000x speedup in query encoding and over 10× increase in end-to-end retrieval throughput.
+Extensive experiments on large-scale retrieval benchmarks show that LightRetriever generalizes well across diverse tasks, maintaining an average of 95% retrieval performance.
 1 Introduction
-Given search queries and candidate documents, text retrieval employs dual-encoder architectures to encode representation vectors and search documents relevant to queries based on query-document similarities [62].
+Given search queries and candidate documents, text retrieval employs dual-encoder architectures to encode representation vectors and search documents relevant to queries based on query-document similarities Salton (1975).
 The encoded vectors could be condensed into low-dimensional vectors for dense retrieval or sparsified into high-dimensional vectors for sparse retrieval.
-Dense retrieval [29] compresses textual information into dense vectors with parameterized language models (LMs) and retrieves based on inner product similarity.
-Meanwhile, sparse retrieval [60] characterizes sparse term frequencies on a much broader vocabulary space, and then searches based on lexical overlap and term-based relevance metrics, such as direct impact search (Inner product) [3, 20] and BM25 [60].
-Additionally, hybrid retrieval enhances retrieval abilities by interpolating their similarity scores [72] or ranks [15].
-Text retrieval is broadly applied in various scenarios, such as web search [43], semantic textual similarity [8], fact checking [66], and retrieval augmented generation [39].
-Figure 1: Architecture comparison of LLM-based hybrid retrieval with 1) Full-sized Query Inference (Previous works); and 2) Extreme Lightweight Query Inference (LightRetriever).
+Dense retrieval Karpukhin et al.
+(2020) compresses textual information into dense vectors with parameterized language models (LMs) and retrieves based on inner product similarity.
+Meanwhile, sparse retrieval Robertson et al.
+(1994) characterizes sparse term frequencies on a much broader vocabulary space, and then searches based on lexical overlap and term-based relevance metrics, such as direct impact search (Inner product) Bai et al.
+(2020); Formal et al.
+(2021) and BM25 Robertson et al.
+(1994).
+Additionally, hybrid retrieval enhances retrieval abilities by interpolating their similarity scores Wang et al.
+(2021) or ranks Cormack et al.
+(2009).
+Text retrieval is broadly applied in various scenarios, such as web search Lu et al.
+(2022), semantic textual similarity Bowman et al.
+(2015), fact checking Thorne et al.
+(2018), and retrieval augmented generation Lewis et al.
+(2020).
+Figure 1: LightRetriever targets at extreme query inference speed for LLM-based text retrieval, reducing the workload of query encoding to no more than an embedding lookup.
 Deploying retrieval systems requires different processing workflows for queries and documents.
-As is shown in Figure 1, as search candidates, documents are encoded with a document encoder $Enc_{d}$ and then indexed into search databases [27, 60].
-Because documents are obtained before the search operation, they are pre-computable and distributable, which can be indexed offline and distributed anywhere [79].
+As is shown in Figure 1, as search candidates, documents are encoded with a document encoder $Enc_{d}$ and then indexed into search databases Johnson et al.
+(2019); Robertson et al.
+(1994).
+Because documents are obtained before the search operation, they are pre-computable and distributable, which can be indexed offline and distributed anywhere Yang et al.
+(2018a).
 On the contrary, queries are input and encoded in real-time, which are non-pre-computable and non-distributable.
@@
 Special GPU accelerators are mandatory if the query encoder is a deeply parameterized LM.
-Most existing works of text retrieval [29, 11, 71] utilize symmetric dual-encoder architectures, which share the same LM for both the query and document encoding or have an equal number of parameters on both sides.
-Recent state-of-the-art (SOTA) retrievers [71, 5] employ deeply parameterized LLMs as dual-encoders $Enc_{q}$ and $Enc_{d}$ for enhanced retrieval capabilities and domain adaptation abilities on various languages and tasks.
-However, serving giant LLMs slows down query inference throughput and depletes online deployment resources [37].
+Most existing works of text retrieval Karpukhin et al.
+(2020); Chen et al.
+(2024); Wang et al.
+(2024b) utilize symmetric dual-encoder architectures, which share the same LM for both the query and document encoding or have an equal number of parameters on both sides.
+Recent state-of-the-art (SOTA) retrievers Wang et al.
+(2024b); BehnamGhader et al.
+(2024) employ deeply parameterized LLMs as dual-encoders $Enc_{q}$ and $Enc_{d}$ for enhanced retrieval capabilities and domain adaptation abilities on various languages and tasks.
+However, serving giant LLMs slows down query inference throughput and depletes online deployment resources Kwon et al.
+(2023).
 Nowadays, LLM-based text retrieval incurs non-negligible online workloads, necessitating special optimizations for query inference efficiency.
-Modern Transformers-based LMs handle text modeling with the attention mechanism [68], including the contextual traits and deep interactions of co-occurrences on discrete tokens within certain texts.
-As is the case for existing LLM-based text retrievers [71, 5], queries and documents are modeled equally with the same LLM, enduring the same computational burdens.
-In contrast, traditional statistical retrieval algorithms, such as the popular TF-IDF [63] and BM25 [60], only characterize limited token co-occurrence by analyzing lexical overlaps with customized retrieval metrics.
+Modern Transformers-based LMs handle text modeling with the attention mechanism Vaswani et al.
+(2017), including the contextual traits and deep interactions of co-occurrences on discrete tokens within certain texts.
+As is the case for existing LLM-based text retrievers Wang et al.
+(2024b); BehnamGhader et al.
+(2024), queries and documents are modeled equally with the same LLM, enduring the same computational burdens.
+In contrast, traditional statistical retrieval algorithms, such as the popular BM25 Robertson et al.
+(1994), only characterize limited token co-occurrence by analyzing lexical overlaps with customized retrieval metrics.
 They do not handle the contextual traits and deep co-occurrences, and their encoders do not need to be deeply parametrized during inference.
-Experiments from [65] show that BM25 still performs well in many retrieval tasks, even beating some previous BERT-based retrievers trained on MSMARCO [52].
+Experiments from Thakur et al.
+(2021) show that BM25 still performs well in many retrieval tasks.
 Built on the above observations, to minimize query service cost, we propose LightRetriever, a novel LLM-based hybrid retrieval architecture targeted at extreme lightweight query inference abilities.
-As is shown in Figure 1, we retain a full-sized LLM on the document side ($Enc_{d}$) for strong retrieval capacities, but completely remove the deep modeling of queries.
+Our work retains a full-sized LLM on the document side ($Enc_{d}$) for strong retrieval capacities.
+But as is shown in Figure 1, we completely remove the deep modeling of queries.
 1) For dense retrieval, our method end-to-end trains cacheable token vectors from the LLM-based query encoder.
@@
 Query inference costs are drastically reduced to only one Embedding lookup.
-2) For sparse retrieval, similar to statistical retrieval algorithms, the mapping of a token and its corresponding frequency [^1] Here we directly use the number of a certain token that appeared in the tokenized query as its frequency.
+2) For sparse retrieval, the mapping of a token and its corresponding frequency [^1] Here we directly use the number of a certain token that appeared in the tokenized query as its frequency.
 is directly used as the sparse query vector.
@@
 The dense similarity scores are computed with cosine similarity, while the sparse scores directly use inner product similarity.
-These scores are linearly interpolated [72] for final hybrid retrieval.
-LightRetriever is super fast and lightweight with respect to query inference.
-Compared to serving a full-sized LLM (e.g., Mistral-7B [26]) on an H800 GPU, our method achieves over a 1000× speedup for query inference with GPU acceleration, and even a 20× speedup without GPU.
-Experiments on large-scale retrieval benchmarks demonstrate that our method generalizes well across diverse LLMs (including Mistral-7B, Llama-1B, 3B, 8B, and Qwen-1.5B, 3B, 7B) and retrieval benchmarks (including 15 English tasks from BeIR [65] and 8 Chinese tasks from CMTEB Retrieval [76]), retaining an average of 95% full-sized performance.
+These scores are summed Wang et al.
+(2021) for the final hybrid retrieval.
+LightRetriever is super fast and lightweight for query inference.
+Compared to serving a full-sized LLM on an A800 GPU, our method achieves over a thousand times speedup for query encoding, and over 10x total throughput improvements.
+Experiments on large-scale retrieval benchmarks demonstrate that our method generalizes well across diverse LLMs (including Llama-1B, 3B, 8B, and Qwen-1.5B, 3B, 7B) and retrieval benchmarks (including 15 English tasks from BeIR Thakur et al.
+(2021) and 8 Chinese tasks from CMTEB Retrieval Xiao et al.
+(2023)), maintaining an average of 95% retrieval performance.
 2 Algorithm
-Figure 2: Full-sized symmetric dense retrieval for baseline training.
+Figure 2: Contrastive training of full-sized symmetric dense retrieval.
+A full-sized dual-encoder is used to model both queries and documents.
+Task-specific instructions are added as common practices Su et al.
+(2023); Wang et al.
+(2024b) to promote better domain adaptation abilities.
 2.1 LM-based text retrieval
@@
 $\min_{\theta}\max_{d^{-}\in\mathcal{N}(q)}\mathcal{E}(q,d^{+},d^{-};\theta)\text{, where }\mathcal{E}(q,d^{+},d^{-};\theta)=v_{q}\cdot v_{d^{-}}-v_{q}\cdot v_{d^{+}}$ (1)
-which lowers the upper boundary of ranking similarity errors $\mathcal{E}$ [12].
-The above objective can be effectively optimized by the pairwise/listwise contrastive loss [10], margin loss [61], etc.
-Following [29], listwise contrastive loss ($\ell^{CL}$) is used in our work,
+which lowers the upper boundary of ranking similarity errors $\mathcal{E}$ Chen et al.
+(2009).
+The above objective can be effectively optimized by the pairwise/listwise contrastive loss Cao et al.
+(2007), margin loss Rosset et al.
+(2003), etc.
+Following Karpukhin et al.
+(2020), listwise contrastive loss ($\ell^{CL}$) is used in our work,
 $\ell^{CL}=-\log\frac{e^{v_{q}\cdot v_{d^{+}}/\tau}}{e^{v_{q}\cdot v_{d^{+}}/\tau}+\sum_{d^{-}\in\mathcal{N}(q)}e^{v_{q}\cdot v_{d^{-}}/\tau}}.$ (2)
@@
 Instructed retrieval
-Similar to instructed fine-tuning (SFT) [55], modern LM-based retrievers [64, 76, 70], especially LLM retrievers [71, 5, 38], customize task specific instructions for better domain adaption abilities.
+Similar to instructed fine-tuning (SFT) Ouyang et al.
+(2022), modern LM-based retrievers Su et al.
+(2023); Xiao et al.
+(2023); Wang et al.
+(2024a), especially LLM retrievers Wang et al.
+(2024b); BehnamGhader et al.
+(2024); Lee et al.
+(2024), customize task specific instructions for better domain adaption abilities.
 Specifically, as shown in Figure 2, the task instruction is prepended to the queries before retrieval.
-Although [64, 76] use instructions on both query and passage sides, most LLM-based retrievers [71, 5, 38] do not use passage instructions to reuse the pre-built indexes across different tasks.
-In our work, dense retrieval follows the query instructions of Mistral-E5 [71] and does not involve passage instructions.
+Although Su et al.
+(2023); Xiao et al.
+(2023) use instructions on both query and passage sides, most LLM-based retrievers Wang et al.
+(2024b); BehnamGhader et al.
+(2024); Lee et al.
+(2024) do not use passage instructions to reuse the pre-built indexes across different tasks.
+In our work, dense retrieval follows the query instructions of Mistral-E5 Wang et al.
+(2024b) and does not involve passage instructions.
 The sparse retrieval is incompatible with instructions because it does not involve learnable query encoders.
@@
 Figure 3: Dense retrieval of LightRetriever.
+Three stages are involved for efficient query modeling.
+1) Training: A shared prompt concatenated with a single query token is passed through a full encoder independently to obtain token-level representations.
+The full query representation is then computed by averaging the token vectors corresponding to each query token.
+2) Caching: Prior to serving, the trained encoder is used to precompute token embeddings for the entire LLM’s vocabulary, which are stored in a single embedding lookup matrix.
+3) Online Serving: At inference time, query embeddings are efficiently generated by lookup and averaging the cached token embeddings, eliminating the need for deep model inference.
 Training
 Given a task instruction $Inst$ and query tokens $T_{q}=\{t_{0},t_{1},...,t_{n-1}\}$, each query token $t_{i}$ ($i\in\{0,...,n-1\}$) are prepended with the instruction.
-Then, the corresponding query token vectors $v_{t_{i}}^{\text{den}}$ are obtained by encoding [^2] Specifically, EOS pooling is used in our work by following Mistral-E5 [71].
+Then, the corresponding query token vectors $v_{t_{i}}^{\text{den}}$ are obtained by encoding [^2] Specifically, EOS pooling is used in our work by following Mistral-E5 Wang et al.
+(2024b).
 with the query encoder [^3] A customized causal mask is designed to avoid repeated prompt computations, detailed in Appendix A.7.,
@@
 Contrastive loss is the main objective for dense training, as defined in Equation 2.
-Following [9, 32], an auxiliary KL loss is also used to align the asymmetric similarity scores $S^{student}=v_{q}^{den}\cdot v_{d}^{den}$ to full-sized symmetric similarity scores $S^{teacher}=Enc_{q}(q)\cdot Enc_{d}(d)$,
-$\ell^{Align}=\texttt{KL-Div}(S^{student},S^{teacher})$ (5)
-Full-sized symmetric baselines could be trained with asymmetric encoders together.
-Thus, the computation of auxiliary loss is economical.
-Experiments show that this helps further narrow the performance gaps, but it is not mandatory.
 Caching
@@
 Figure 4: Sparse retrieval of LightRetriever.
+The query representation is a term-based, unlearnable sparse vector, which is obtained from tokenization.
+The document representation is end-to-end trained by projecting last layer hidden states of LLM to vocabulary space, then optimized via contrastive learning.
 Training
@@
 These hidden states are then projected into the vocabulary space to get logits $w$ via the model’s own language modeling head, parameterized by a projection matrix $P$.
-Following SPLADE [20], to induce sparsity and suppress dominant term frequencies, we apply a ReLU activation followed by a log-saturation function, and then perform a max-pooling operation of the logits $w$ over the sequence length dimension,
+Following SPLADE Formal et al.
+(2021), to induce sparsity and suppress dominant term frequencies, we apply a ReLU activation followed by a log-saturation function, and then perform a max-pooling operation of the logits $w$ over the sequence length dimension,
 $v_{d}^{\text{spr}}=\max(w)\text{, where }w=\text{logits}=\ln(\max(h_{\text{last}}\cdot P,0)+1)$ (7)
@@
 Thus, the sparsify regulator is needed during the training phase.
-Following SPLADE [20], given a batch of documents $d=\{d_{0},...,d_{i}\},i<\text{batch size }N$, the FLOPs regulator [56] is used for sparsification.
+Following SPLADE Formal et al.
+(2021), given a batch of documents $d=\{d_{0},...,d_{i}\},i<\text{batch size }N$, the FLOPs regulator Paria et al.
+(2020) is used for sparsification.
 This regulator first averages the logits across different documents within a certain batch, then computes the squared sum to get the final regulator.
@@
 The sparse query vector $v_{q}^{\text{spr}}$ during online serving follows the same rule in Equation 6. No more query encoder for sparse retrieval is needed.
+Table 1: Comparison of retrieval speed and effectiveness.
+We compare the time consumption of retrieving 65,536 Bing queries over 1 million passages from the MS MARCO dataset Nguyen et al.
+(2016), evaluating both full symmetric retrievers and the query-lightweight LightRetriever.
+To assess the trade-off between efficiency and effectiveness, we also report nDCG@10 scores on English (BeIR) and Chinese (CMTEB-Retrieval) benchmarks.
+For retrieval efficiency, we report the consumed time of: query tokenization, model encoding, maximum search time of Faiss/Lucene, and total end-to-end retrieval time.
+The overall throughput is measured in Queries Per Second (QPS).
+Benchmark / nDCG@10 Time consumption / s & Throughput / QPS
+Model BeIR CMTEB-R Tokenize Encode Search Total QPS
+Full-Llama8b 56.8 67.6 1.3746 109.4853 8.5133 119.3730 549
+Full-Llama1b 53.1 63.8 1.2564 19.9030 8.4897 29.6490 2210
+LightRetriever-Llama8b 54.4 63.0 0.8209 0.0412 8.5010 9.3630 6999
+Full-Qwen7b 56.6 70.1 1.4144 100.6716 8.5279 110.6140 592
+Full-Qwen1.5b 54.0 65.9 1.4210 53.0858 8.5299 63.0370 1040
+LightRetriever-Qwen7b 53.8 66.5 0.8300 0.0420 8.5150 9.3870 6982
 3 Experiments
 3.1 Settings
-Datasets
-A total of 23 training sets (including 20 English and 3 Chinese) are used in our experiments by following [47, 71].
-Most of them are available at preprocessed data  [^4] https://huggingface.co/datasets/tdro-llm/finetune_data from tDRO [47] and Sentence Transformers Training Data  [^5] https://huggingface.co/datasets/sentence-transformers/embedding-training-data [59].
-Detailed training sets and sampling ratios are listed in the Appendix A.3.
-BeIR [65] and CMTEB Retrieval [76] benchmarks are used in the evaluation.
+Training data
+Our work finetunes on a large existing collection of 20 English and 3 Chinese datasets with 8.38M samples to ensure broad domain coverage and diversity.
+To ensure reproducibility, our work reuses the preprocessed fine-tuning data from tDRO Ma et al.
+(2025) and Sentence Transformers Training Data Reimers (2019).
+English datasets include Amazon Review Ni et al.
+(2019), Eli5 Fan et al.
+(2019), FEVER Thorne et al.
+(2018), FiQA Maia et al.
+(2018), GooQA Khashabi et al.
+(2021), HotpotQA Yang et al.
+(2018b), MSMARCO Nguyen et al.
+(2016), NFCorpus Boteva et al.
+(2016), NPR Lucas et al.
+(2023), NQ Kwiatkowski et al.
+(2019), PAQ Lewis et al.
+(2021), Quora Duplicates Iyer et al.
+(2012), S2ORC Lo et al.
+(2020), Scifact Wadden et al.
+(2020), Specter Cohan et al.
+(2020), StackExchange Duplicates da Silva et al.
+(2018), E5 synthetic BeastyZ (2024), Trivia Joshi et al.
+(2017), WikiHow Koupaee and Wang (2018), and Yahoo Answers Zhang et al.
+(2015).
+Chinese datasets include cMedQA2 Zhang et al.
+(2018), DuReader Qiu et al.
+(2022), and T2Ranking Xie et al.
+(2023).
+Benchmarks
+BeIR Thakur et al.
+(2021) and CMTEB Retrieval Xiao et al.
+(2023) benchmarks are used in the evaluation.
 BeIR is a massive English retrieval benchmark collection with 15 evaluation sets.
@@
 CMTEB Retrieval is a massive Chinese retrieval benchmark collection with 8 sets, also covering a variety of domains.
-nDCG@10 is reported as the main metric by following [65, 76], accessing both recall and ranking abilities.
+nDCG@10 is reported as the main metric by following Thakur et al.
+(2021); Xiao et al.
+(2023), accessing both recall and ranking abilities.
 Recalls (R@{20, 50, 100}) are also reported to assess retrieval ability at large windows.
-Models & hyper-parameters
-We modify the experiment code based on the implementation  [^6] https://github.com/ma787639046/tdro from tDRO [47].
+Training hyper-parameters
+Our experiments are conducted based on the implementation of tDRO Ma et al.
+(2025).
 All experiments are performed with a batch size of 128, 7 hard negatives, a max sequence length of 512, a contrastive temperature $\tau$ of 0.02, and 12k total steps.
-Following SPLADE [20], the coefficient of the FLOPs regulator is 0.001, and the regulator is quadratically increased to maximum at first 4k steps, which helps reduce the side effects of the regulator at initial steps [56].
-LoRA [24] is used to save GPU memory usage during the LLM training, where LoRA r is 16, alpha is 32, and dropout is 0.1.
-Different LLMs with different sizes are tested as the backbone encoders to verify the generalization abilities, including Mistral-v0.3-7B [26], Llama-3.2-1B, 3B, Llama-3.1-8B [67], and Qwen-2.5-1.5B, 3B, 7B [2].
-Direct linear interpolation [72] is used to combine dense and sparse similarity scores as the hybrid retrieval scores.
-All results are conducted on a DGX node with 8 NVIDIA H800 GPUs, 128 CPU cores of 2 Intel Xeon 8462Y+, and 2 TB of memory.
-Baselines
-Full-sized symmetric dense retrievers are our main baselines under the same training conditions, as shown in Figure 2.
-The query inference speeds and retrieval performance gaps are compared as the main results.
-Additionally, we also include some recent retrievers in the results, including Sentence Transformers static embedding [1], BM25 [60], BGE-m3 [11], E5-Mistral [71], and LLM2Vec [5].
-However, they are not our main comparison baselines, because [1] and [60] are parameter-efficient but less performant.
-LLM-based retrievers [71, 5] are performant but inefficient.
-Table 1: Comparison of query encoding time and speedup for full-sized baseline and LightRetriever.
-Tokenization time is also presented for reference.
-65536 queries from MSMARCO [52] are processed.
-Evaluations are conducted on an NVIDIA DGX node using 1 H800 GPU with Mistral-v0.3-7B.
-Tokenzation Baseline LightRetriever
-GPU Encode No-GPU Encode GPU Encode
-Batch Size Time (s) Time (s) Time (s) speedup Time (s) speedup
-16 5.1168 69.9276 3.7065 19 0.1509 463
-64 3.8995 42.2991 3.1719 13 0.0510 829
-256 2.0713 45.8708 2.9990 15 0.0185 2480
-1024 1.4226 57.4545 2.7969 21 0.0107 5370
-4096 1.6798 74.7221 2.8218 26 0.0093 8035
-8192 OOM 2.8529 0.0090
-3.2 Main results
-Encoding speeds
-The encoding speed comparisons of full-sized baseline and LightRetriever are presented in Table 1.
-No-GPU and GPU settings are measured for LightRetriever.
-Because of the efficiency issue, the baseline is only measured on the GPU.
-As is shown in Table 1, it takes a full-sized Mistral-7B LLM up to 74s for encoding 65536 queries on an H800 GPU across different batch sizes.
-In comparison, our LightRetriever only consumes less than 0.1509s on the GPU, which achieves over 400-8000x speedup.
-Even when encoding without GPU acceleration, LightRetriever also achieves over 10-20x speedup compared to the GPU baseline.
-Table 2: Retrieval performances of LightRetriever on BeIR and CMTEB Retrieval (CMTEB-R) benchmarks.
-The nDCG@10 gaps between LightRetriever hybrid and the corresponding baseline are presented.
-The best metrics of baselines and LightRetrievers are marked in bold.
+Following SPLADE Formal et al.
+(2021), the coefficient of the FLOPs regulator is 0.001, and the regulator is quadratically increased to the maximum at the first 4k steps, which helps reduce the side effects of the regulator at initial steps Paria et al.
+(2020).
+LoRA Hu et al.
+(2022) is used to save training GPU memory, where LoRA r is 16, alpha is 32, and dropout is 0.1.
+Different LLMs with different sizes are tested as the backbone encoders to verify the generalization abilities, including Llama-3.2-1B, 3B, Llama-3.1-8B Touvron et al.
+(2023), and Qwen-2.5-1.5B, 3B, 7B Bai et al.
+(2023).
+The hybrid similarity scores are linearly summed from normalized dense and sparse scores Wang et al.
+(2021).
+All trainings are conducted on 8 NVIDIA H800 or A800 GPUs.
+Speed comparisons
+To evaluate the online speedup of LightRetriever, we compare the online time consumptions for retrieving 65536 Bing queries over 1M passages from MSMARCONguyen et al.
+(2016).
+All speed tests are conducted on a single A800 GPU, with a batch size of 256, a dimension of 1k for Faiss exact search, and 64 threads for Anserini (Lucene) sparse search.
+Faiss and Anserini search in parallel, where Faiss searches more slowly in our cases.
+Table 2: Performance comparisons on BeIR and CMTEB Retrieval (CMTEB-R) benchmarks.
+The best metrics of full symmetric retrievers and LightRetrievers are marked in bold.
+Their nDCG@10 gaps are also presented.
 Benchmark BEIR (15 datasets) CMTEB-R (8 datasets)
@@
 BM25 41.7 48.8 56.5 61.8 50.8 63.9 70.0 74.1
+Static Embed + BM25 44.7 51.9 59.8 65.1 52.1 65.8 71.9 76.1
 […]
@@
 […]
-E5-Mistral 56.9 62.1 68.9 73.6 61.8 75.1 81.6 87.1
-Llama3.2-1b
-Baseline 53.1 60.2 66.9 72.0 63.8 78.1 84.1 87.9
-Light Dense 48.7 55.8 63.0 68.3 58.7 73.1 79.7 83.8
-Light Sparse 47.6 54.5 61.5 66.7 53.9 68.8 75.7 80.1
-Light Hybrid 52.0-${}_{\text{-1.1}}$ 58.1 65.1 70.1 60.8-${}_{\text{-3.0}}$ 74.8 81.2 85.0
-Llama3.2-3b
-Baseline 55.6 62.9 69.7 74.7 66.1 81.0 86.6 90.0
-Light Dense 51.0 58.6 65.5 70.5 60.4 75.1 81.5 85.3
-Light Sparse 48.4 55.5 62.7 67.7 53.0 68.0 75.0 79.6
-Light Hybrid 53.5-${}_{\text{-2.1}}$ 59.9 66.8 71.7 61.7-${}_{\text{-4.4}}$ 76.3 82.3 86.3
-Llama3.1-8b
-Baseline 56.8 64.3 71.1 75.7 67.6 82.0 87.4 90.9
-Light Dense 52.0 59.5 66.5 71.4 61.4 75.6 82.3 86.4
-Light Sparse 49.5 56.9 64.1 69.1 54.8 69.8 77.3 81.8
-Light Hybrid 54.4-${}_{\text{-2.4}}$ 60.9 67.7 72.8 63.0-${}_{\text{-4.6}}$ 77.3 83.6 87.4
-Mistral0.3-7b
-Baseline 57.5 64.1 71.0 75.7 65.9 80.3 86.5 90.2
-Light Dense 52.7 59.2 66.5 71.3 57.4 72.1 79.0 83.6
-Light Sparse 48.0 56.1 63.3 68.3 50.7 65.5 72.7 77.8
-Light Hybrid 54.5-${}_{\text{-3.0}}$ 60.1 67.4 72.1 59.4-${}_{\text{-6.5}}$ 73.9 80.6 84.8
-Qwen2.5-1.5b
-Baseline 54.0 60.5 67.5 72.3 65.9 81.0 86.7 90.4
-Light Dense 48.9 56.0 63.1 68.1 61.9 77.0 83.2 87.4
-Light Sparse 47.3 54.3 61.4 66.4 52.5 67.4 74.9 79.9
-Light Hybrid 52.1-${}_{\text{-1.9}}$ 58.2 65.1 70.0 63.8-${}_{\text{-2.1}}$ 78.3 84.5 88.3
-Qwen2.5-3b
-Baseline 54.9 62.0 68.7 73.5 69.4 83.8 88.9 91.8
-Light Dense 50.1 57.2 64.2 69.3 64.7 79.4 85.2 88.8
-Light Sparse 47.5 54.9 62.0 67.2 54.1 69.2 76.5 81.1
-Light Hybrid 52.8-${}_{\text{-2.1}}$ 59.1 66.0 70.9 65.7-${}_{\text{-3.7}}$ 80.2 86.1 89.3
-Qwen2.5-7b
-Baseline 56.6 63.6 70.5 75.2 70.1 84.4 89.4 92.4
-Light Dense 50.6 58.6 65.6 70.5 65.0 79.8 85.8 89.3
-Light Sparse 48.2 56.0 63.5 68.5 55.0 69.7 77.0 82.1
-Light Hybrid 53.8-${}_{\text{-2.8}}$ 60.3 67.5 72.5 66.5-${}_{\text{-3.6}}$ 81.1 86.7 89.7
-Retrieval performances
-The retrieval performances are presented in Table 2.
-Experiments are conducted with 3 series of LLMs with different sizes.
-For each base model, we test the full-sized baseline and dense/sparse/hybrid (dense + sparse) retrieval performances of LightRetriever (Light).
-As is shown in the results, the performance gaps between LightRetriever (Hybrid) and its corresponding baseline are averaged around 5%, ranging from 2.1% (-1.1 point) to 9.9% (-6.5 point).
-Mistral [26] gets the highest nDCG@10 scores on the English BeIR benchmark, while Llama models [67] are better at recalls of BeIR.
-Qwen series [2] performs the best on the Chinese CMTEB Retrieval benchmark, attributed to better Chinese pre-training of Qwen.
-Figure 5: Performances (Llama-3.1-8b) with different Top-k dimensions or sparsities.
-3.3 Controls of vector dimension and sparsity
-Retrieval systems usually involve indexing millions of documents into search databases, which consumes large amounts of disk space, especially for dense vectors.
-For example, indexing 8.8M MS-MARCO passages [52] with Llama-8b model consumes 134.92 GB for Faiss dense retrieval and 11.24 GB for Lucene sparse retrieval.
-Thus, it’s essential to explore the possibility of controlling index sizes by shrinking vectors.
-1) For dense retrieval, we incorporate Matryoshka Representation Learning (MRL) [35] with LightRetriever to support flexible controls of dense vector dimensions.
-Given a dense vector $v^{den}$ with dimension $H$, MRL derives dimension-shrunk vectors by simply cutting out the Top-k dimensions $v^{den}[:k]$, then applies the same training loss as the original vectors.
-MRL trains a collection of Top-k dimensions together, enabling flexible choosing of a proper k.
-2) For sparse retrieval, the sparse vector directly learns the term impact (i.e., lexical frequency) within a large sparse dimension.
-Thus, LightRetriever supports controlling sparsity by directly taking the Top-k terms, where the larger impact terms matter more.
-We conduct experiments of controlling dense vector dimension and sparse vector sparsity with Top-k of $\{128,256,512,1024,2048,4096\}$ on Llama-3.1-8b.
-As presented in Figure 5, the performances of LightRetriever decrease marginally when shrinking Top-k to 4x ($4090\rightarrow 1024$).
-The lowest Top-k setting ($128$) has 7.9% and 8.8% performance gaps on BeIR and CMTEB-R benchmarks compared to the highest Top-k ($4096$), but its disk usages with MS-MARCO are only 4.22 GB for dense trieval and 3.03 GB for sparse retrieval.
-This shows that our LightRetriever is capable of flexible index controls via simple vector shrinking techniques, with limited effect on retrieval performances.
-Table 3: Ablations (nDCG@10, w/Llama-3.1-8b) on BeIR and CMTEB-R about auxiliary KL loss.
-Benchmark w/ KL wo/ KL
-BeIR CMTEB-R BeIR CMTEB-R
-Baseline 56.8 67.6 57.0+0.2 67.9+0.3
-Light Dense 52.0 61.4 50.8-1.2 60.1-1.3
-Light Sparse 49.5 54.8 48.7-0.8 54.4-0.4
-Light Hybrid 54.4 63.0 53.9-0.5 62.1-0.9
-3.4 Effect of auxiliary KL loss
-Previous works for BERT-based asymmetric retrievers [73, 32] emphasize the importance of alignment between asymmetric students and full symmetric teachers.
-Our work also uses the KL loss as an auxiliary.
-Results in Table 3 show that removing such loss incurs limited drops of -0.5 and -0.9 on BeIR and CMTEB-R.
-Thus, such auxiliary KL loss is not mandatory in our experiment settings.
-Table 4: Ablations of dense and sparse retrieval for LightRetriever (Llama-3.2-1b).
-Dense Ablation BeIR CMTEB-R
-LightRetriever 48.7 58.7
-Symmetry
-w/ Qry Full - Doc Light 8.7-40.0 9.4-49.3
-w/ Qry Light - Doc Light 34.9-13.8 40.1-18.6
-Other Ablations
-w/ Direct Emb Layer 37.5-11.2 41.3-17.4
-wo/ Instruction 47.1-1.6 56.1-2.6
-Sparse Ablation BeIR CMTEB-R
-LightRetriever 47.6 53.9
-Symmetry
-w/ Qry Full - Doc Light 16.2-31.4 10.8-43.1
-Other Ablations
-w/ Value fixed to 1 47.6+0.0 54.7+0.8
+[…]
+E5-Mistral7b 56.9 62.1 68.9 73.6 61.8 75.1 81.6 87.1
+Full Symmetric Retrievers
+Llama3.2-1b 53.1 60.2 66.9 72.0 63.8 78.1 84.1 87.9
+Llama3.2-3b 55.6 62.9 69.7 74.7 66.1 81.0 86.6 90.0
+Llama3.1-8b 56.8 64.3 71.1 75.7 67.6 82.0 87.4 90.9
+Qwen2.5-1.5b 54.0 60.5 67.5 72.3 65.9 81.0 86.7 90.4
+Qwen2.5-3b 54.9 62.0 68.7 73.5 69.4 83.8 88.9 91.8
+Qwen2.5-7b 56.6 63.6 70.5 75.2 70.1 84.4 89.4 92.4
+LightRetriever
+Llama3.2-1b 52.0-$_{\text{-1.1}}$ 58.1 65.1 70.1 60.8-$_{\text{-3.0}}$ 74.8 81.2 85.0
+Llama3.2-3b 53.5-$_{\text{-2.1}}$ 59.9 66.8 71.7 61.7-$_{\text{-4.4}}$ 76.3 82.3 86.3
+Llama3.1-8b 54.4-$_{\text{-2.4}}$ 60.9 67.7 72.8 63.0-$_{\text{-4.6}}$ 77.3 83.6 87.4
+Qwen2.5-1.5b 52.1-$_{\text{-1.9}}$ 58.2 65.1 70.0 63.8-$_{\text{-2.1}}$ 78.3 84.5 88.3
+Qwen2.5-3b 52.8-$_{\text{-2.1}}$ 59.1 66.0 70.9 65.7-$_{\text{-3.7}}$ 80.2 86.1 89.3
+Qwen2.5-7b 53.8-$_{\text{-2.8}}$ 60.3 67.5 72.5 66.5-$_{\text{-3.6}}$ 81.1 86.7 89.7
+Performance comparisons
+[…]
+Notably, Static Embedding performs retrieval using a single shared embedding layer for both queries and documents, offering low retrieval performance but highly efficient baseline.
+BM25, as a classical term-frequency-based method, remains a widely adopted sparse retriever.
+The combination of them thus provides a reference for efficient retrieval on both sides.
+In contrast, BGE-m3 is a Transformer-based retriever trained with multi-stage pretraining and fine-tuning, and represents state-of-the-art performance in hybrid retrieval.
+Except for Static Embedding and BM25, all compared retrievers rely on deep Transformer encoders on both query and document sides.
+LightRetriever is, to our knowledge, the first method to address the online efficiency bottleneck on the query side of LLM-based retrievers, striking a balance between retrieval quality and inference efficiency.
+3.2 Main results
+Encoding speeds
+Table 1 presents a detailed breakdown of retrieval time for processing 65,536 Bing queries over 1 million MS MARCO passages, highlighting the efficiency advantages of LightRetriever over full symmetric retrievers.
+Among all time consumptions, model encoding is the dominant contributor to latency in LLM-based retrievers.
+For example, Full-Llama8b and Full-Qwen7b require over 100 seconds solely for query encoding (109.5s and 100.7s, respectively), making them a bottleneck for large-scale, real-time retrieval.
+In contrast, LightRetriever reduces encoding time to below 50 ms (0.0412–0.0420s), achieving over a 2500× speedup in the encoding phase.
+This drastic reduction is made possible by replacing full forward passes of large Transformer models with a simple embedding lookup.
+And the overall throughput achieves over 10x QPS speedup.
+Notably, this lightweight encoding does not significantly compromise retrieval performance.
+LightRetriever maintains competitive nDCG@10 scores on both English (BeIR) and Chinese (CMTEB-Retrieval) benchmarks.
+These results prove the effectiveness of LightRetriever’s decoupled query modeling paradigm, which drastically improves online query encoding efficiency while preserving good retrieval quality, making it suitable for latency-critical applications.
+Performance comparisons
+Table 2 compares retrieval effectiveness across a wide range of retrievers on both English (BeIR) and Chinese (CMTEB-Retrieval) benchmarks.
+The main focus of this comparison is to evaluate the performance degradation introduced by LightRetriever’s lightweight query encoding.
+Overall, by leveraging deep Transformer encoders on both query and document sides, full symmetric retrievers achieve the highest retrieval scores.
+For instance, Llama3.1-8b performs best on BeIR (56.8 nDCG@10), while Qwen2.5-7b reaches the best performance on CMTEB-R (70.1 nDCG@10).
+However, such performance comes at the cost of significant computational overhead, as discussed above.
+[…]
+These results demonstrate that LightRetriever effectively balances retrieval quality and efficiency.
+By avoiding expensive query-side encoding and maintaining high retrieval performance, it establishes a promising new paradigm for scalable LLM-based retrieval systems, especially in latency-sensitive or high-throughput scenarios.
 4 Ablations
-4.1 Symmetry
-Our work proposes an asymmetric architecture with an extremely lightweight query (Qry) encoder and a full-sized document (Doc) encoder.
-Besides, we also perform ablations on other symmetry settings, including lightweight on both query and document, and a full-sized query encoder with a lightweight passage encoder.
-Results in Table 4 show that a full-sized query encoder with a lightweight passage encoder leads to drastic performance degenerations, maybe caused by embedding collapses.
-Lightweight on both sides for dense retrieval works, but it does not perform excellently in retrieval tasks, just like Static Embedding [1].
-Lightweight on both sides is not suitable for our sparse retrieval, because no learnable parameters are available.
-4.2 Other ablations
-Do we need a full query model during the dense training?
-Dense retrieval of LightRetriever trains token vectors with full Transformers.
-Fast retrieval is achieved by caching learned dense token vectors and lookups during serving.
-If we directly use the Embedding Layer within the Transformers, the performances drop 11.2 and 17.4 points on BeIR and CMTEB-R.
-This verifies that during the training phase, a full-sized model is still needed for learning query token vectors.
-Task instructions
-Task instructions contribute +1.6 and +2.6 points on BeIR and CMTEB-R of dense retrieval. Similar effects of instructions are also observed in previous LLM retrievers [71, 5].
-Value of sparse query vector
-The sparse query vector is a key-value type of mapping.
-In Equation 7, we model the value of sparse query vectors by counting the corresponding token ids that appear in the query.
-For ablations, we choose a simpler way of modeling by stopping the counting and fixing the value to 1.
-In Table 4, BeIR results do not change.
-However, CMTEB-R results improve by +0.8 points.
-This indicates that counting term frequency in our sparse query vector seems unimportant to some extent, under large amounts of training on LightRetriever.
-The co-occurrence is more important.
+Our core design adopts an asymmetric architecture, with a lightweight query MLP Embedding and a full-sized document model for inference.
+Specifically, we utilize a full-sized model on both sides at training time, then cache all query token embeddings in one Embedding Lookup for fast query-side inference.
+In ablation studies below (A1-A3), we are interested in other potential symmetric/asymmetric architectures, which are described in the following Table 3.
+Table 3: Symmetry Ablation settings (A1-A3).
+Train Inference
+Query Doc Query Doc
+Ours Full Full MLP Full
+A1 Full Full MLP MLP
+A2 MLP Full MLP Full
+A3 1-layer Transformers Full 1-layer Transformers Full
+A1. Is symmetric lightweight inference effective?
+To test the necessity of the above asymmetry, we evaluate a symmetric lightweight setup where both sides use simplified dense encoders.
+As shown in Table 4, this leads to severe performance drops on both BeIR (–13.8) and CMTEB-R (–18.6), confirming that full document representations are crucial for maintaining retrieval effectiveness.
+This symmetric setting is not functional for sparse retrieval due to the absence of learnable parameters.
+The best symmetric lightweight retriever is Static Embedding (one MLP) + BM25 (term-based) in Table 2, with nDCG@10 of 44.7 on BeIR and 52.1 on CMTEB-R.
+However, due to the lack of deep modeling on documents, such a combination is still outperformed.
+We also tested a reversed asymmetry with deep encoding for queries and lightweight documents. However, we observed training instability and degraded performance, likely caused by embedding collapse.
+A2. Can we use only a query MLP for training?
+To confirm that a deep query encoder is still needed during training, we replaced the Transformer-based query encoder with a simple MLP that reuses only the LLM’s embedding layer.
+This setup significantly degrades performance (–11.2 on BeIR, –17.4 on CMTEB-R), indicating that the full-sized modeling for query side training remains essential during training to learn effective token-level representations.
+Table 4: Ablations results of LightRetriever
+Ablations / nDCG@10 BeIR CMTEB-R
+Llama-1b 48.7 58.7
+w/ Both-side Lightweight 34.9-13.8 40.1-18.6
+w/ Encq use a MLP 37.5-11.2 41.3-17.4
+Llama-8b 54.4 63.0
+w/ 1st-layer on query 50.1-4.3 54.4-8.6
+A3. Is adding one Transformers layer better on the query side?
+We also evaluated using only the first Transformer layer on the query side.
+While this reduces model size during training, it still underperforms the LightRetriever (–4.3 on BeIR, –8.6 on CMTEB-R) and introduces inference-time overhead, making it incompatible with the design goals of LightRetriever.
+This again proves the need for full-sized query modeling during the training phase.
+Figure 5: Ablations of retrieval performances (nDCG@10) of LightRetriever on BeIR and CMTEB-R benchmarks.
+A4. Breakdown Ablation of Retrieval Performance
+Figure 5 reports LightRetriever’s performance under dense, sparse, and hybrid configurations.
+Dense retrieval consistently outperforms sparse across all backbones (e.g., LLaMA3, Qwen2.5) and model sizes (1B–8B).
+While sparse retrieval alone is less effective, the hybrid approach recovers most of the performance gap with minimal overhead.
+For instance, LLaMA3-1B achieves 52.0 and 60.8 nDCG@10 on BeIR and CMTEB-R in the hybrid setting, with only 1.1 and 3.0 points below the full model, but with a much smaller query encoder.
+These results confirm the complementarity of both signals and the effectiveness of LightRetriever in balancing efficiency and accuracy.
 5 Conclusion
@@
 A.2.1 Text retrieval
-Text retrieval builds on top of the vector space model [62], which encodes query and document representation vectors, and performs relevance search based on vector similarities.
+Text retrieval builds on top of the vector space model Salton [1975], which encodes query and document representation vectors, and performs relevance search based on vector similarities.
 Dense retrieval
-Dense retrieval [29] trains PLM-based encoders to condense queries and passages into low-dimensional dense vectors, then performs relevance search based on the Maximum Inner Product Search (MIPS) [51] algorithm.
-In recent years, dense retrievers, such as BGE [11] and E5 [70], have gained popularity for strong retrieval abilities across different tasks [50, 65], languages [83], and context granulaties [74, 52], owing to diverse training data [59], retrieval-customized pre-training [21, 75, 46], improved negative-mining [78], and so on.
-Recent SOTA retrievers [71, 5, 38] start to utilize LLMs [26, 67] as backbone encoders.
+Dense retrieval Karpukhin et al.
+[2020] trains PLM-based encoders to condense queries and passages into low-dimensional dense vectors, then performs relevance search based on the Maximum Inner Product Search (MIPS) Mussmann and Ermon [2016] algorithm.
+In recent years, dense retrievers, such as BGE Chen et al.
+[2024] and E5 Wang et al.
+[2024a], have gained popularity for strong retrieval abilities across different tasks Muennighoff et al.
+[2023], Thakur et al.
+[2021], languages Zhang et al.
+[2023], and context granulaties Williams et al.
+[2018], Nguyen et al.
+[2016], owing to diverse training data Reimers [2019], retrieval-customized pre-training Gao and Callan [2021], Xiao et al.
+[2022], Ma et al.
+[2024], improved negative-mining Xiong et al.
+[2021], and so on.
+Recent SOTA retrievers Wang et al.
+[2024b], BehnamGhader et al.
+[2024], Lee et al.
+[2024] start to utilize LLMs Jiang et al.
+[2023], Touvron et al.
+[2023] as backbone encoders.
 With larger parameters and pre-training data, LLM-based retrievers enable significantly powerful retrieval capacities and domain adaptation abilities.
@@
 Sparse retrieval
-Traditional statistical sparse retrieval algorithms (e.g., TF-IDF [63], and BM25 [60]) do not involve learnable language models.
+Traditional statistical sparse retrieval algorithms (e.g., TF-IDF Salton and Buckley [1988], and BM25 Robertson et al.
+[1994]) do not involve learnable language models.
 Instead, they directly analyze lexical frequencies (e.g., TF) and importance (e.g., IDF) in a large vocabulary space.
-These lexicals are sparsely distributed and require building inverted indexes [45] for feasible similarity metric calculation, such as Apache Lucene [6] and Tantivy [58].
+These lexicals are sparsely distributed and require building inverted indexes Luhn [1957] for feasible similarity metric calculation, such as Apache Lucene Bialecki et al.
+[2012] and Tantivy Quickwit Inc.
+[2024].
 In recent years, sparse retrieval has started to work with pre-trained LMs (PLMs).
-BGE [11], SPLADE [20], SparTerm [3], and SPARTA [84] use customized XLM-RoBERTa [14] or BERT [18] as backbone encoders to learn term frequencies in an end-to-end manner and directly perform impact search based on the dot product of sparse vectors.
+BGE Chen et al.
+[2024], SPLADE Formal et al.
+[2021], SparTerm Bai et al.
+[2020], and SPARTA Zhao et al.
+[2021] use customized XLM-RoBERTa Conneau et al.
+[2020] or BERT Devlin et al.
+[2019] as backbone encoders to learn term frequencies in an end-to-end manner and directly perform impact search based on the dot product of sparse vectors.
 Because of sparsity, these works also search with inverted indexes.
-Recent works [85] also try to incorporate LLMs with sparse retrieval.
+Recent works Zhuang et al.
+[2024] also try to incorporate LLMs with sparse retrieval.
 However, the online efficiency issue still exists.
 Hybrid retreval
-Hybrid retrieval is a technique that combines retrieval scores [72] or ranks [15] from multiple retrieval systems, such as dense [29] / sparse [60, 20] retrieval, and single [29] / multiple [31, 11] vector retrieval.
+Hybrid retrieval is a technique that combines retrieval scores Wang et al.
+[2021] or ranks Cormack et al.
+[2009] from multiple retrieval systems, such as dense Karpukhin et al.
+[2020] / sparse Robertson et al.
+[1994], Formal et al.
+[2021] retrieval, and single Karpukhin et al.
+[2020] / multiple Khattab and Zaharia [2020], Chen et al.
+[2024] vector retrieval.
 It enables better retrieval abilities by interpolating results from multiple sources.
-Linear interpolation [72] of dense and sparse scores is used in our work.
+Linear interpolation Wang et al.
+[2021] of dense and sparse scores is used in our work.
 A.2.2 Improving inference efficiency
@@
 Previous works have tried to distill small query encoders from large encoders for better query inference efficiency.
-KALE [9] prunes the BERT-based query encoder and distills it during post-training.
-However, when the query encoder is pruned to one Transformers layer, its recall@20 on NQ [36] decreases by 25-30%.
-[73] proposes asymmetric dual-encoders by pruning BERT layers or using smaller BERT models.
+KALE Campos et al.
+[2023] prunes the BERT-based query encoder and distills it during post-training.
+However, when the query encoder is pruned to one Transformers layer, its recall@20 on NQ Kwiatkowski et al.
+[2019] decreases by 25-30%.
+Wang and Lyu [2023] proposes asymmetric dual-encoders by pruning BERT layers or using smaller BERT models.
 Its experiments show that a 2-layer BERT query encoder retains 92.5% of the full-sized performances via distillation and proper student initialization.
-USTAD [32] also proposes similar asymmetry and distillation strategies, which choose the 12-layer BERT as teachers and the 6-layer DistilBERT or 2-layer BERT-mini as students.
+USTAD Kim et al.
+[2024] also proposes similar asymmetry and distillation strategies, which choose the 12-layer BERT as teachers and the 6-layer DistilBERT or 2-layer BERT-mini as students.
 Its experiments show that the distilled BERT-mini query encoder achieves 95-97% of the teacher’s performance because of better distillation.
-However, these studies have primarily focused on small language models (SLMs), such as BERT [18], and overlooked feasible approaches for improving online inference efficiency of LLM-based retrievers, where such optimization is even more critical.
+However, these studies have primarily focused on small language models (SLMs), such as BERT Devlin et al.
+[2019], and overlooked feasible approaches for improving online inference efficiency of LLM-based retrievers, where such optimization is even more critical.
 Furthermore, these existing methods typically retain Transformer layers in the query encoders, which limits their efficiency gains.
-In addition, most of these works rely on SLMs and are trained on relatively narrow datasets, such as MS-MARCO [52], thus unable to compete with LLM-based retrievers.
+In addition, most of these works rely on SLMs and are trained on relatively narrow datasets, such as MS-MARCO Nguyen et al.
+[2016], thus unable to compete with LLM-based retrievers.
 In contrast, with the rapid advancement of LLM-based text retrievers, driven by larger and more diverse foundation models and datasets, retrieval performance has significantly improved.
@@
 Previous works have also been dedicated to reducing index sizes for smaller disk space usage.
-Matryoshka representation learning (MRL) [35] supports adjusting the dense vector dimensions by directly taking the top-k dimension of the original vector as shrink vectors.
+Matryoshka representation learning (MRL) Kusupati et al.
+[2022] supports adjusting the dense vector dimensions by directly taking the top-k dimension of the original vector as shrink vectors.
 It aligns and trains the corresponding multi-dimensional vectors in the end-to-end contrastive learning.
 For example, if $k\in\{64,128,256\}$, query-passage dense vectors with $\{64,128,256\}$-dims could be obtained together after one encoder forward operation.
-For LM-based sparse retrieval, the similarity scores are directly affected by the impacts of corresponding lexicals [20].
+For LM-based sparse retrieval, the similarity scores are directly affected by the impacts of corresponding lexicals Formal et al.
+[2021].
 Thus, it naturally supports reducing the index sizes by retaining Top-k terms/lexicals.
-To explore the potential of controlling index sizes, our work also explores shrinking index sizes with dense MRL and sparse Top-k in the Experiments Section 3.3, which also shows strong generalization capacities.
-A.3 Training datasets
+To explore the potential of controlling index sizes, our work also explores shrinking index sizes with dense MRL and sparse Top-k in the Section A.3, which also shows strong generalization capacities.
+Figure 6: Performances (Llama-3.1-8b) with different Top-k dimensions or sparsities.
+A.3 Additional results about controls of vector dimension and sparsity
+Retrieval systems usually involve indexing millions of documents into search databases, which consumes large amounts of disk space, especially for dense vectors.
+For example, indexing 8.8M MS-MARCO passages Nguyen et al.
+[2016] with Llama-8b model consumes 134.92 GB for Faiss dense retrieval and 11.24 GB for Lucene sparse retrieval.
+Thus, it’s essential to explore the possibility of controlling index sizes by shrinking vectors.
+1) For dense retrieval, we incorporate Matryoshka Representation Learning (MRL) Kusupati et al.
+[2022] with LightRetriever to support flexible controls of dense vector dimensions.
+Given a dense vector $v^{den}$ with dimension $H$, MRL derives dimension-shrunk vectors by simply cutting out the Top-k dimensions $v^{den}[:k]$, then applies the same training loss as the original vectors.
+MRL trains a collection of Top-k dimensions together, enabling flexible choosing of a proper k.
+2) For sparse retrieval, the sparse vector directly learns the term impact (i.e., lexical frequency) within a large sparse dimension.
+Thus, LightRetriever supports controlling sparsity by directly taking the Top-k terms, where the larger impact terms matter more.
+We conduct experiments of controlling dense vector dimension and sparse vector sparsity with Top-k of $\{128,256,512,1024,2048,4096\}$ on Llama-3.1-8b.
+As presented in Figure 6, the performances of LightRetriever decrease marginally when shrinking Top-k to 4x ($4090\rightarrow 1024$).
+The lowest Top-k setting ($128$) has 7.9% and 8.8% performance gaps on BeIR and CMTEB-R benchmarks compared to the highest Top-k ($4096$), but its disk usages with MS-MARCO are only 4.22 GB for dense trieval and 3.03 GB for sparse retrieval.
+This shows that our LightRetriever is capable of flexible index controls via simple vector shrinking techniques, with limited effect on retrieval performances.
+A.4 Additional ablation about the effect of auxiliary KL loss
+Table 5: Ablations (nDCG@10, w/Llama-3.1-8b) on BeIR and CMTEB-R about auxiliary KL loss.
+Benchmark w/ KL wo/ KL
+BeIR CMTEB-R BeIR CMTEB-R
+Dense 52.0 61.4 50.8-1.2 60.1-1.3
+Sparse 49.5 54.8 48.7-0.8 54.4-0.4
+Hybrid 54.4 63.0 53.9-0.5 62.1-0.9
+Previous works for BERT-based asymmetric retrievers Wang and Lyu [2023], Kim et al.
+[2024] emphasize the importance of alignment between asymmetric students and full symmetric teachers.
+To explore the effect of such alignment, besides contrastive loss as the main loss function, we also explore to adopt the KL loss as an auxiliary to align the asymmetric similarity scores $S^{student}=v_{q}^{den}\cdot v_{d}^{den}$ to full-sized symmetric similarity scores $S^{teacher}=Enc_{q}(q)\cdot Enc_{d}(d)$,
+$\ell^{Align}=\texttt{KL-Div}(S^{student},S^{teacher})$ (9)
+Results in Table 5 show that removing such loss incurs limited drops of -0.5 and -0.9 on BeIR and CMTEB-R. Thus, such auxiliary KL loss is not mandatory needed in our experiments.
+A.5 Training datasets
 Dataset infos
-Following the settings of previous LLM-based retrievers [71, 5, 47], our work uses 23 training sets pre-built by tDRO [47] or Sentence Transformers Training Data [59], including 20 English and 3 Chinese datasets.
+Following the settings of previous LLM-based retrievers Wang et al.
+[2024b], BehnamGhader et al.
+[2024], Ma et al.
+[2025], our work uses 23 training sets pre-built by tDRO Ma et al.
+[2025] or Sentence Transformers Training Data Reimers [2019], including 20 English and 3 Chinese datasets.
 Detailed dataset information is listed in Table 6 as follows.
@@
 Dataset Language Category Deduped Size Epoch Ratio
-Amazon Review (2018) [53] English Amazon 999999 0.01 0.58%
-cMedQA2 [81] Chinese Chinese Medical 99936 1 5.84%
-DuReader [57] Chinese Chinese Web Collections 86366 1 5.05%
-Eli5 [19] English Reddit 325390 0.05 0.95%
-Fever [66] English Wikipedia QA 109808 0.8 5.13%
-FiQA [48] English Financial 5498 1 0.32%
-GooQA Pairs [30] English Web Collections 3012347 0.025 4.40%
-HotpotQA [80] English Wikipedia QA 85000 0.5 2.48%
-MSMARCO [52] English Web Collections 502854 1 29.39%
-NFCorpus [7] English Medical 2585 1 0.15%
-NPR [44] English News 594376 0.05 1.74%
-NQ [36] English Wikipedia QA 58800 1 3.44%
-PQA Pairs [40] English Wikipedia QA 99999 0.5 2.92%
-Quora Duplicates Triples [25] English Forum Duplicates 97011 0.5 2.83%
-S2ORC Title-Abstract [42] English Semantic Scholar 99998 0.5 2.92%
-SciFact [69] English S2ORC 806 1 0.05%
-SPECTER [13] English Semantic Scholar 136642 0.25 2.00%
-StackExchange Duplicates (Title-Body) [16] English Forum Duplicates 250516 0.25 3.66%
-E5 Synthetic [4] English GPT Synthetic Data 224791 1 13.14%
-T2Ranking [77] Chinese Chinese Web Collections 200362 0.5 5.86%
-Trivia [28] English Wikipedia QA 60370 0.5 1.76%
-Wikihow [33] English WikiHow 128543 0.25 1.88%
-Yahoo Answers (Title-Answer) [82] English Yahoo 1198018 0.05 3.50%
-Many downstream retrieval benchmarks [65, 76] are evaluated out-of-domain, where no training samples from the downstream sources are available.
-To avoid the overfitting of training sets, as is shown in the above Table 5, we slightly tune the epoch used in the dataset sampling, by following previous works in [54, 71, 47].
+Amazon Review (2018) Ni et al. [2019] English Amazon $999$ 0.01 0.58%
+cMedQA2 Zhang et al. [2018] Chinese Chinese Medical 99936 1 5.84%
+DuReader Qiu et al. [2022] Chinese Chinese Web Collections 86366 1 5.05%
+Eli5 Fan et al. [2019] English Reddit 325390 0.05 0.95%
+Fever Thorne et al. [2018] English Wikipedia QA 109808 0.8 5.13%
+FiQA Maia et al. [2018] English Financial 5498 1 0.32%
+GooQA Pairs Khashabi et al. [2021] English Web Collections 3012347 0.025 4.40%
+HotpotQA Yang et al. [2018b] English Wikipedia QA 85000 0.5 2.48%
+MSMARCO Nguyen et al. [2016] English Web Collections 502854 1 29.39%
+NFCorpus Boteva et al. [2016] English Medical 2585 1 0.15%
+NPR Lucas et al. [2023] English News 594376 0.05 1.74%
+NQ Kwiatkowski et al. [2019] English Wikipedia QA 58800 1 3.44%
+PQA Pairs Lewis et al. [2021] English Wikipedia QA 99999 0.5 2.92%
+Quora Duplicates Triples Iyer et al. [2012] English Forum Duplicates 97011 0.5 2.83%
+S2ORC Title-Abstract Lo et al. [2020] English Semantic Scholar 99998 0.5 2.92%
+SciFact Wadden et al. [2020] English S2ORC 806 1 0.05%
+SPECTER Cohan et al. [2020] English Semantic Scholar 136642 0.25 2.00%
+StackExchange Duplicates (Title-Body) da Silva et al. [2018] English Forum Duplicates 250516 0.25 3.66%
+E5 Synthetic BeastyZ [2024] English GPT Synthetic Data 224791 1 13.14%
+T2Ranking Xie et al. [2023] Chinese Chinese Web Collections 200362 0.5 5.86%
+Trivia Joshi et al. [2017] English Wikipedia QA 60370 0.5 1.76%
+Wikihow Koupaee and Wang [2018] English WikiHow 128543 0.25 1.88%
+Yahoo Answers (Title-Answer) Zhang et al. [2015] English Yahoo 1198018 0.05 3.50%
+Many downstream retrieval benchmarks Thakur et al.
+[2021], Xiao et al.
+[2023] are evaluated out-of-domain, where no training samples from the downstream sources are available.
+To avoid the overfitting of training sets, as is shown in the above Table 6, we slightly tune the epoch used in the dataset sampling, by following previous works in Ni et al.
+[2022], Wang et al.
+[2024b], Ma et al.
+[2025].
 Training set decontamination
@@
 Existing training datasets are derived from multiple sources, whose training data have possible duplicates with the test sets.
-Following tDRO [47], we perform strict training set decontamination with SimHash [49] to ensure no training queries appear in the BeIR [65] and CMTEB Retrieval [76].
+Following tDRO Ma et al.
+[2025], we perform strict training set decontamination with SimHash Manku et al.
+[2007] to ensure no training queries appear in the BeIR Thakur et al.
+[2021] and CMTEB Retrieval Xiao et al.
+[2023].
 The original sizes, deduplicated sizes, and duplicate numbers as listed in the Table 7.
@@
 Dataset Original Size Deduped Size Duplicates
-Amazon Review (2018) 1000000 999999 1
+Amazon Review (2018) 1000000 $999$ 1
 cMedQA2 100000 99936 64
@@
 Eli5 train-3088 How does unemployment insurance work? FiQA 2648 How does unemployment insurance work?
-A.4 Task instructions
-As previously mentioned in Section 2.1, modern LLM-based retrievers [71, 5] introduce task-specific instructions to format input queries, which enables higher retrieval capacities.
-These LLM-based instruction-tuned retrieval is similar to instruction-based supervised fine-tuning (SFT) in LLM generation [55].
-Following Mistral-E5 [71], we reuse most of its task instructions, and format the instruction and query in the following format:
+A.6 Task instructions
+As previously mentioned in Section 2.1, modern LLM-based retrievers Wang et al.
+[2024b], BehnamGhader et al.
+[2024] introduce task-specific instructions to format input queries, which enables higher retrieval capacities.
+These LLM-based instruction-tuned retrieval is similar to instruction-based supervised fine-tuning (SFT) in LLM generation Ouyang et al.
+[2022].
+Following Mistral-E5 Wang et al.
+[2024b], we reuse most of its task instructions, and format the instruction and query in the following format:
 […]
-The detailed instructions are listed below, most of which are directly copied from the previous works E5-Mistral [71] and tDRO [47] for reproducibility.
+The detailed instructions are listed below, most of which are directly copied from the previous works E5-Mistral Wang et al.
+[2024b] and tDRO Ma et al.
+[2025] for reproducibility.
 Note that one training set may correspond to multiple instructions to avoid overfitting.
@@
     VideoRetrieval: Given a video search query, retrieve the titles of relevant videos
-A.5 Efficiency optimizations techniques
+A.7 Efficiency optimizations techniques
 Customized causal mask
@@
 Specifically, given a task instruction $Inst$ and query tokens $T_{q}=\{t_{0},t_{1},...,t_{n-1}\}$, we format the instruction and query tokens as follows,
-[…]
+$\text{Input IDs}=\texttt{<bos>}~\text{Inst}~t_{0}~\texttt{<eos>}~\dots~t_{n-1}~\texttt{<eos>}$ (10)
 Figure 7: An example of the customized attention mask.
@@
 Let the sequence length be $L$, prompt length be $P$, micro block (each token + <eos>) width be $w$, and let $q\in[0,L-1]$ denote the query index, $k\in[0,L-1]$ the key index.
-The attention mask (0 means attend, $\infty$- ∞ means masked) is defined as,
-$\text{Mask}[q,k]=\begin{cases}0,&\text{if }q<P\text{ and }k\leq q\quad\text{(attn within prompt)}\\ 0,&\text{if }q\geq P\text{ and }k<P\quad\text{(blocks can attend to prompt)}\\ 0,&\text{if }q\geq P\text{ and }\left\lfloor\frac{q-P}{w}\right\rfloor=\left\lfloor\frac{k-P}{w}\right\rfloor\text{ and }k\leq q\quad\text{(attn within current block)}\\ -\infty,&\text{otherwise}\end{cases}$
+The attention mask (0 means attend, -$\infty$ means masked) is defined as,
+[…]
 An example of the customized causal mask is shown in Figure 7.
@@
 , where the vocabulary dimension $V$ is much larger than the hidden dimension $H$.
-For example, for Llama3.1-8b [67] model, $V=128256$ and $H=4096$, where $\lfloor V/H\rfloor=31$.
+For example, for Llama3.1-8b Touvron et al.
+[2023] model, $V=128256$ and $H=4096$, where $\lfloor V/H\rfloor=31$.
 This means that, given batch size B, sequence length T, hidden size H, vocabulary size V, the intermediate tensors with shape $[B,T,V]$ in the projection, ReLU, and Log are 31x larger than a dense tensor $[B,T,H]$, which consumes huge amounts of GPU memory during training and inference.
@@
 4:  Initialize $Y\leftarrow$-$\infty$
-5:     for $t=0$ to $T-1$ do
+5:  for $t=0$ to T1T-1 do
 6:   $L_{t}\leftarrow X[:,t,:]\cdot W$
@@
 10:   if $M$ is not None then
-11:              for $i=0$ to $B-1$ do
+11:     for $i=0$ to B1B-1 do
 12:      if $M[i,t]=0$ then
@@
 23:  Initialize $\nabla X\leftarrow 0$, $\nabla W\leftarrow 0$, $\nabla b\leftarrow 0$
-24:     for $t=0$ to $T{-}1$ do
+24:  for $t=0$ to T1T{-}1 do
 25:   $X_{t}\leftarrow X[:,t,:]\in\mathbb{R}^{B\times H}$
@@
 27:   $M_{t}\leftarrow\texttt{max\_indices}_{I==t}\in\{0,1\}^{B\times V}$
-28:         $\nabla L_{t}\leftarrow\nabla Y\odot M_{t}\in\mathbb{R}^{B\times V}$triangleright Only keep grads where $t$ was max
+28:   $\nabla L_{t}\leftarrow\nabla Y\odot M_{t}\in\mathbb{R}^{B\times V}$ $\triangleright$ Only keep grads where $t$ was max
 29:   $\nabla X_{t}\mathrel{+}=\nabla L_{t}\cdot W^{\top}$
@@
 Proper optimizations are important.
-For example, we found it hard to reproduce quickly with some of the open-sourced baselines [5], which is roughly 6x slower than our implementation when testing with the same size of LLM.
+For example, we found it hard to reproduce quickly with some of the open-sourced baselines BehnamGhader et al.
+[2024], which is roughly 6x slower than our implementation when testing with the same size of LLM.
 Our experiments involve training with 23 train sets and inference with another 23 test sets, which requires proper optimizations urgently.
-We import other optimizations based on the implementations  [^7] Please refer to [47]’s implementation on https://github.com/ma787639046/tdro.
-released by tDRO [47].
-These optimizations include 1) A distributed multi-node inference framework based on PyTorch RPC [41]; 2) Sequence packing [34] with Flash Attention 2 [17], which gives roughly 50% training speedup due to the removal of the pad tokens.; 3) Fused Kernels from Liger-Kernel [23], which gives roughly 10% speedup to both training and inferencing; 4) Multi-vector gradient accumulation support with GradCache [22] for optionally enlarging batch sizes if necassary.
-A.6 Detailed main results
+We import other optimizations based on the implementations  [^4] Please refer to Ma et al.
+[2025]’s implementation on https://github.com/ma787639046/tdro.
+released by tDRO Ma et al.
+[2025].
+These optimizations include 1) A distributed multi-node inference framework based on PyTorch RPC Li et al.
+[2023]; 2) Sequence packing Kundu et al.
+[2024] with Flash Attention 2 Dao [2023], which gives roughly 50% training speedup due to the removal of the pad tokens.; 3) Fused Kernels from Liger-Kernel Hsu et al.
+[2024], which gives roughly 10% speedup to both training and inferencing; 4) Multi-vector gradient accumulation support with GradCache Gao et al.
+[2021] for optionally enlarging batch sizes if necassary.
+A.8 Detailed main results
 The detailed main results (nDCG@10) on each dataset are shown in the tables below, which are full versions of Table 2.
@@
 Llama3.2-1b
-Baseline 58.0 44.0 41.2 43.2 89.2 45.5 69.3 42.7
-Light Dense 49.2 38.1 37.9 34.2 83.6 40.4 60.8 39.2
-Light Sparse 49.5 34.5 27.5 35.9 84.7 35.7 65.5 37.2
-Light Hybrid 54.1 39.7 39.3 39.5 87.3 42.1 68.2 41.1
+Full Symmetric 58.0 44.0 41.2 43.2 89.2 45.5 69.3 42.7
+LightRetriever Dense 49.2 38.1 37.9 34.2 83.6 40.4 60.8 39.2
+LightRetriever Sparse 49.5 34.5 27.5 35.9 84.7 35.7 65.5 37.2
+LightRetriever Hybrid 54.1 39.7 39.3 39.5 87.3 42.1 68.2 41.1
 Llama3.2-3b
-Baseline 58.5 46.8 43.7 46.4 89.9 53.4 73.7 44.3
-Light Dense 55.3 40.3 41.7 36.5 84.0 46.2 64.3 41.6
-Light Sparse 46.5 34.1 26.9 38.5 86.2 38.9 66.3 39.2
-Light Hybrid 56.3 41.4 42.4 41.6 87.6 46.9 69.7 42.9
+Full Symmetric 58.5 46.8 43.7 46.4 89.9 53.4 73.7 44.3
+LightRetriever Dense 55.3 40.3 41.7 36.5 84.0 46.2 64.3 41.6
+LightRetriever Sparse 46.5 34.1 26.9 38.5 86.2 38.9 66.3 39.2
+LightRetriever Hybrid 56.3 41.4 42.4 41.6 87.6 46.9 69.7 42.9
 Llama3.1-8b
-Baseline 57.1 49.2 44.1 48.2 90.9 59.0 77.8 46.2
-Light Dense 55.8 40.9 41.5 38.7 85.7 49.8 67.6 43.4
-Light Sparse 46.4 33.6 32.2 38.8 86.1 41.9 67.9 41.0
-Light Hybrid 57.2 41.9 42.4 42.0 87.8 50.7 71.5 44.5
+Full Symmetric 57.1 49.2 44.1 48.2 90.9 59.0 77.8 46.2
+LightRetriever Dense 55.8 40.9 41.5 38.7 85.7 49.8 67.6 43.4
+LightRetriever Sparse 46.4 33.6 32.2 38.8 86.1 41.9 67.9 41.0
+LightRetriever Hybrid 57.2 41.9 42.4 42.0 87.8 50.7 71.5 44.5
 Mistral0.3-7b
-Baseline 56.2 49.7 43.6 48.4 91.3 59.6 77.7 46.6
-Light Dense 53.5 41.6 40.9 38.9 85.6 50.9 66.9 44.2
-Light Sparse 44.0 33.6 21.9 37.0 81.0 43.6 65.4 40.9
-Light Hybrid 54.3 42.2 41.4 40.5 86.9 51.7 69.9 44.8
+Full Symmetric 56.2 49.7 43.6 48.4 91.3 59.6 77.7 46.6
+LightRetriever Dense 53.5 41.6 40.9 38.9 85.6 50.9 66.9 44.2
+LightRetriever Sparse 44.0 33.6 21.9 37.0 81.0 43.6 65.4 40.9
+LightRetriever Hybrid 54.3 42.2 41.4 40.5 86.9 51.7 69.9 44.8
 Qwen2.5-1.5b
-Baseline 56.6 44.1 40.4 43.9 88.9 47.9 68.1 42.6
-Light Dense 51.2 38.3 36.1 33.0 82.1 41.2 59.0 39.1
-Light Sparse 46.5 35.0 25.7 36.3 85.1 35.3 63.3 37.3
-Light Hybrid 54.1 40.1 37.7 39.5 87.0 42.1 66.5 40.8
+Full Symmetric 56.6 44.1 40.4 43.9 88.9 47.9 68.1 42.6
+LightRetriever Dense 51.2 38.3 36.1 33.0 82.1 41.2 59.0 39.1
+LightRetriever Sparse 46.5 35.0 25.7 36.3 85.1 35.3 63.3 37.3
+LightRetriever Hybrid 54.1 40.1 37.7 39.5 87.0 42.1 66.5 40.8
 Qwen2.5-3b
-Baseline 55.5 47.0 44.1 45.8 89.6 52.9 71.6 43.7
-Light Dense 49.8 39.6 41.1 35.0 83.9 45.4 62.8 41.0
-Light Sparse 45.5 36.2 28.7 36.9 84.9 38.6 64.2 38.7
-Light Hybrid 52.8 41.4 42.1 39.8 86.9 46.1 68.2 42.3
+Full Symmetric 55.5 47.0 44.1 45.8 89.6 52.9 71.6 43.7
+LightRetriever Dense 49.8 39.6 41.1 35.0 83.9 45.4 62.8 41.0
+LightRetriever Sparse 45.5 36.2 28.7 36.9 84.9 38.6 64.2 38.7
+LightRetriever Hybrid 52.8 41.4 42.1 39.8 86.9 46.1 68.2 42.3
 Qwen2.5-7b
-Baseline 57.1 48.2 43.5 47.9 90.7 56.5 74.7 44.8
-Light Dense 53.3 40.0 39.1 36.8 85.1 48.8 64.4 41.8
-Light Sparse 47.5 35.6 28.2 36.7 85.4 40.2 65.2 40.1
-Light Hybrid 54.4 41.6 42.2 41.0 87.9 49.7 69.4 43.3
+Full Symmetric 57.1 48.2 43.5 47.9 90.7 56.5 74.7 44.8
+LightRetriever Dense 53.3 40.0 39.1 36.8 85.1 48.8 64.4 41.8
+LightRetriever Sparse 47.5 35.6 28.2 36.7 85.4 40.2 65.2 40.1
+LightRetriever Hybrid 54.4 41.6 42.2 41.0 87.9 49.7 69.4 43.3
 Table 10: (Part2) Detailed main results (nDCG@10) on BeIR test sets.
@@
 Llama3.2-1b
-Baseline 35.5 58.6 89.4 19.8 72.3 66.7 20.9 53.1
-Light Dense 30.1 52.2 87.0 17.2 66.5 68.8 24.7 48.7
-Light Sparse 34.0 52.0 82.2 17.2 69.1 60.6 27.8 47.6
-Light Hybrid 34.2 56.6 87.4 18.6 71.5 70.5 30.2 52.0
+Full Symmetric 35.5 58.6 89.4 19.8 72.3 66.7 20.9 53.1
+LightRetriever Dense 30.1 52.2 87.0 17.2 66.5 68.8 24.7 48.7
+LightRetriever Sparse 34.0 52.0 82.2 17.2 69.1 60.6 27.8 47.6
+LightRetriever Hybrid 34.2 56.6 87.4 18.6 71.5 70.5 30.2 52.0
 Llama3.2-3b
-Baseline 37.6 62.8 89.5 22.5 75.3 68.7 21.2 55.6
-Light Dense 31.7 56.4 87.8 19.0 70.4 68.6 21.6 51.0
-Light Sparse 33.3 55.3 83.2 17.7 67.3 64.1 28.5 48.4
-Light Hybrid 34.2 59.8 88.1 20.0 71.4 70.9 29.5 53.5
+Full Symmetric 37.6 62.8 89.5 22.5 75.3 68.7 21.2 55.6
+LightRetriever Dense 31.7 56.4 87.8 19.0 70.4 68.6 21.6 51.0
+LightRetriever Sparse 33.3 55.3 83.2 17.7 67.3 64.1 28.5 48.4
+LightRetriever Hybrid 34.2 59.8 88.1 20.0 71.4 70.9 29.5 53.5
 Llama3.1-8b
-Baseline 38.6 65.5 89.9 23.8 77.3 62.0 22.2 56.8
-Light Dense 31.8 59.0 88.5 19.3 70.6 65.7 21.1 52.0
-Light Sparse 33.6 58.4 84.4 18.9 69.7 63.0 27.1 49.5
-Light Hybrid 34.6 61.9 88.6 20.7 72.6 71.5 27.8 54.4
+Full Symmetric 38.6 65.5 89.9 23.8 77.3 62.0 22.2 56.8
+LightRetriever Dense 31.8 59.0 88.5 19.3 70.6 65.7 21.1 52.0
+LightRetriever Sparse 33.6 58.4 84.4 18.9 69.7 63.0 27.1 49.5
+LightRetriever Hybrid 34.6 61.9 88.6 20.7 72.6 71.5 27.8 54.4
 Mistral0.3-7b
-Baseline 36.8 66.9 90.0 22.7 77.7 71.3 23.6 57.5
-Light Dense 31.9 59.8 88.7 18.5 70.3 71.5 26.7 52.7
-Light Sparse 33.1 57.9 84.6 18.1 66.7 61.1 31.7 48.0
-Light Hybrid 33.8 62.5 88.8 19.8 71.6 76.7 32.4 54.5
+Full Symmetric 36.8 66.9 90.0 22.7 77.7 71.3 23.6 57.5
+LightRetriever Dense 31.9 59.8 88.7 18.5 70.3 71.5 26.7 52.7
+LightRetriever Sparse 33.1 57.9 84.6 18.1 66.7 61.1 31.7 48.0
+LightRetriever Hybrid 33.8 62.5 88.8 19.8 71.6 76.7 32.4 54.5
 Qwen2.5-1.5b
-Baseline 36.4 57.1 89.0 20.4 72.7 73.0 28.4 54.0
-Light Dense 29.8 50.9 86.7 18.1 66.5 72.6 29.1 48.9
-Light Sparse 32.9 51.6 81.3 17.4 66.4 64.8 31.1 47.3
-Light Hybrid 34.0 55.8 87.0 19.3 69.2 76.8 32.3 52.1
+Full Symmetric 36.4 57.1 89.0 20.4 72.7 73.0 28.4 54.0
+LightRetriever Dense 29.8 50.9 86.7 18.1 66.5 72.6 29.1 48.9
+LightRetriever Sparse 32.9 51.6 81.3 17.4 66.4 64.8 31.1 47.3
+LightRetriever Hybrid 34.0 55.8 87.0 19.3 69.2 76.8 32.3 52.1
 Qwen2.5-3b
-Baseline 35.9 61.4 89.5 22.3 72.9 65.7 26.0 54.9
-Light Dense 30.6 54.5 87.5 18.5 64.8 71.2 26.1 50.1
-Light Sparse 33.2 54.2 82.2 17.6 67.6 57.0 26.4 47.5
-Light Hybrid 34.2 58.5 87.8 19.5 69.1 74.2 29.8 52.8
+Full Symmetric 35.9 61.4 89.5 22.3 72.9 65.7 26.0 54.9
+LightRetriever Dense 30.6 54.5 87.5 18.5 64.8 71.2 26.1 50.1
+LightRetriever Sparse 33.2 54.2 82.2 17.6 67.6 57.0 26.4 47.5
+LightRetriever Hybrid 34.2 58.5 87.8 19.5 69.1 74.2 29.8 52.8
 Qwen2.5-7b
-Baseline 37.2 63.7 89.7 23.7 77.2 69.1 25.3 56.6
-Light Dense 30.7 55.9 87.6 19.5 67.9 64.1 24.1 50.6
-Light Sparse 33.9 55.5 78.7 18.3 69.0 60.5 28.5 48.2
-Light Hybrid 34.4 60.1 87.9 21.1 71.1 73.1 29.4 53.8
+Full Symmetric 37.2 63.7 89.7 23.7 77.2 69.1 25.3 56.6
+LightRetriever Dense 30.7 55.9 87.6 19.5 67.9 64.1 24.1 50.6
+LightRetriever Sparse 33.9 55.5 78.7 18.3 69.0 60.5 28.5 48.2
+LightRetriever Hybrid 34.4 60.1 87.9 21.1 71.1 73.1 29.4 53.8
 Table 11: Detailed main results (nDCG@10) on CMTEB Retrieval dev sets.
@@
 Llama3.2-1b
-Baseline 32.3 74.2 84.6 53.8 73.5 50.5 78.0 63.1 63.8
-Light Dense 27.2 67.6 80.8 51.0 67.9 43.0 73.4 58.5 58.7
-Light Sparse 17.3 64.3 74.4 51.0 63.3 39.4 69.3 52.2 53.9
-Light Hybrid 27.2 69.7 81.8 55.5 69.8 44.1 75.9 62.3 60.8
+Full Symmetric 32.3 74.2 84.6 53.8 73.5 50.5 78.0 63.1 63.8
+LightRetriever Dense 27.2 67.6 80.8 51.0 67.9 43.0 73.4 58.5 58.7
+LightRetriever Sparse 17.3 64.3 74.4 51.0 63.3 39.4 69.3 52.2 53.9
+LightRetriever Hybrid 27.2 69.7 81.8 55.5 69.8 44.1 75.9 62.3 60.8
 Llama3.2-3b
-Baseline 36.8 75.6 87.9 57.7 75.7 53.8 81.1 60.5 66.1
-Light Dense 29.8 70.2 84.0 51.7 69.2 46.5 75.7 56.4 60.4
-Light Sparse 17.2 65.4 76.1 47.4 65.0 42.5 68.5 42.0 53.0
-Light Hybrid 29.9 71.5 84.8 54.5 71.2 47.2 77.1 57.7 61.7
+Full Symmetric 36.8 75.6 87.9 57.7 75.7 53.8 81.1 60.5 66.1
+LightRetriever Dense 29.8 70.2 84.0 51.7 69.2 46.5 75.7 56.4 60.4
+LightRetriever Sparse 17.2 65.4 76.1 47.4 65.0 42.5 68.5 42.0 53.0
+LightRetriever Hybrid 29.9 71.5 84.8 54.5 71.2 47.2 77.1 57.7 61.7
 Llama3.1-8b
-Baseline 39.5 77.1 89.5 58.9 78.1 57.6 84.1 56.2 67.6
-Light Dense 32.1 71.0 85.7 53.0 70.2 49.0 78.5 52.0 61.4
-Light Sparse 15.6 70.4 82.9 47.3 63.9 42.2 75.1 41.3 54.8
-Light Hybrid 32.1 72.8 86.6 55.3 71.6 49.7 80.1 55.8 63.0
+Full Symmetric 39.5 77.1 89.5 58.9 78.1 57.6 84.1 56.2 67.6
+LightRetriever Dense 32.1 71.0 85.7 53.0 70.2 49.0 78.5 52.0 61.4
+LightRetriever Sparse 15.6 70.4 82.9 47.3 63.9 42.2 75.1 41.3 54.8
+LightRetriever Hybrid 32.1 72.8 86.6 55.3 71.6 49.7 80.1 55.8 63.0
 Mistral0.3-7b
-Baseline 38.5 74.1 88.6 54.7 75.6 56.6 82.3 56.5 65.9
-Light Dense 30.4 64.4 83.6 47.0 64.3 46.3 74.7 48.5 57.4
-Light Sparse 11.5 63.2 80.5 43.4 61.7 36.7 71.3 37.4 50.7
-Light Hybrid 30.3 67.2 84.5 50.7 67.6 46.9 76.4 51.4 59.4
+Full Symmetric 38.5 74.1 88.6 54.7 75.6 56.6 82.3 56.5 65.9
+LightRetriever Dense 30.4 64.4 83.6 47.0 64.3 46.3 74.7 48.5 57.4
+LightRetriever Sparse 11.5 63.2 80.5 43.4 61.7 36.7 71.3 37.4 50.7
+LightRetriever Hybrid 30.3 67.2 84.5 50.7 67.6 46.9 76.4 51.4 59.4
 Qwen2.5-1.5b
-Baseline 37.9 78.2 87.6 56.9 75.9 54.7 78.8 57.0 65.9
-Light Dense 31.3 73.8 85.3 52.8 71.8 49.5 75.2 55.7 61.9
-Light Sparse 14.5 66.7 78.1 49.3 61.6 39.1 63.5 47.0 52.5
-Light Hybrid 31.2 74.4 86.0 56.5 73.4 50.5 76.4 61.6 63.8
+Full Symmetric 37.9 78.2 87.6 56.9 75.9 54.7 78.8 57.0 65.9
+LightRetriever Dense 31.3 73.8 85.3 52.8 71.8 49.5 75.2 55.7 61.9
+LightRetriever Sparse 14.5 66.7 78.1 49.3 61.6 39.1 63.5 47.0 52.5
+LightRetriever Hybrid 31.2 74.4 86.0 56.5 73.4 50.5 76.4 61.6 63.8
 Qwen2.5-3b
-Baseline 39.4 80.1 89.5 61.9 77.8 58.7 84.4 63.6 69.4
-Light Dense 32.7 76.0 86.8 56.7 73.2 51.6 80.9 59.5 64.7
-Light Sparse 12.8 70.4 80.2 48.8 61.1 39.4 72.4 47.9 54.1
-Light Hybrid 32.6 77.0 87.6 58.3 74.0 51.6 82.0 62.8 65.7
+Full Symmetric 39.4 80.1 89.5 61.9 77.8 58.7 84.4 63.6 69.4
+LightRetriever Dense 32.7 76.0 86.8 56.7 73.2 51.6 80.9 59.5 64.7
+LightRetriever Sparse 12.8 70.4 80.2 48.8 61.1 39.4 72.4 47.9 54.1
+LightRetriever Hybrid 32.6 77.0 87.6 58.3 74.0 51.6 82.0 62.8 65.7
 Qwen2.5-7b
-Baseline 41.3 81.4 91.0 61.1 80.0 59.8 85.5 60.3 70.1
-Light Dense 34.3 74.6 88.0 55.7 75.4 53.2 81.6 57.2 65.0
-Light Sparse 12.5 72.5 82.4 48.3 65.2 40.4 73.3 45.2 55.0
-Light Hybrid 34.0 76.7 88.6 58.0 76.2 53.9 83.0 61.4 66.5
+Full Symmetric 41.3 81.4 91.0 61.1 80.0 59.8 85.5 60.3 70.1
+LightRetriever Dense 34.3 74.6 88.0 55.7 75.4 53.2 81.6 57.2 65.0
+LightRetriever Sparse 12.5 72.5 82.4 48.3 65.2 40.4 73.3 45.2 55.0
+LightRetriever Hybrid 34.0 76.7 88.6 58.0 76.2 53.9 83.0 61.4 66.5

### v4

@@
-LightRetriever: A LLM-based Hybrid Retrieval Architecture with 1000× Faster Query Inference
+LightRetriever: A LLM-based Text Retrieval Architecture with Extremely Faster Query Inference
 Abstract
@@
 1 Introduction
-Given search queries and candidate documents, text retrieval employs dual-encoder architectures to encode representation vectors and search documents relevant to queries based on query-document similarities Salton (1975).
+Given search queries and candidate documents, text retrieval employs dual-encoder architectures to encode representation vectors and search documents relevant to queries based on query-document similarities [63].
 The encoded vectors could be condensed into low-dimensional vectors for dense retrieval or sparsified into high-dimensional vectors for sparse retrieval.
-Dense retrieval Karpukhin et al.
-(2020) compresses textual information into dense vectors with parameterized language models (LMs) and retrieves based on inner product similarity.
-Meanwhile, sparse retrieval Robertson et al.
-(1994) characterizes sparse term frequencies on a much broader vocabulary space, and then searches based on lexical overlap and term-based relevance metrics, such as direct impact search (Inner product) Bai et al.
-(2020); Formal et al.
-(2021) and BM25 Robertson et al.
-(1994).
-Additionally, hybrid retrieval enhances retrieval abilities by interpolating their similarity scores Wang et al.
-(2021) or ranks Cormack et al.
-(2009).
-Text retrieval is broadly applied in various scenarios, such as web search Lu et al.
-(2022), semantic textual similarity Bowman et al.
-(2015), fact checking Thorne et al.
-(2018), and retrieval augmented generation Lewis et al.
-(2020).
+Dense retrieval [30] compresses textual information into dense vectors with parameterized language models (LMs) and retrieves based on inner product similarity.
+Meanwhile, sparse retrieval [61] characterizes sparse term frequencies on a much broader vocabulary space, and then searches based on lexical overlap and term-based relevance metrics, such as direct impact search (Inner product) [3, 21] and BM25 [61].
+Additionally, hybrid retrieval enhances retrieval abilities by interpolating their similarity scores [73] or ranks [15].
+Text retrieval is broadly applied in various scenarios, such as web search [44], semantic textual similarity [8], fact checking [67], and retrieval augmented generation [40].
 Figure 1: LightRetriever targets at extreme query inference speed for LLM-based text retrieval, reducing the workload of query encoding to no more than an embedding lookup.
 Deploying retrieval systems requires different processing workflows for queries and documents.
-As is shown in Figure 1, as search candidates, documents are encoded with a document encoder $Enc_{d}$ and then indexed into search databases Johnson et al.
-(2019); Robertson et al.
-(1994).
-Because documents are obtained before the search operation, they are pre-computable and distributable, which can be indexed offline and distributed anywhere Yang et al.
-(2018a).
+As is shown in Figure 1, as search candidates, documents are encoded with a document encoder $Enc_{d}$ and then indexed into search databases [28, 61].
+Because documents are obtained before the search operation, they are pre-computable and distributable, which can be indexed offline and distributed anywhere [80].
 On the contrary, queries are input and encoded in real-time, which are non-pre-computable and non-distributable.
@@
 Special GPU accelerators are mandatory if the query encoder is a deeply parameterized LM.
-Most existing works of text retrieval Karpukhin et al.
-(2020); Chen et al.
-(2024); Wang et al.
-(2024b) utilize symmetric dual-encoder architectures, which share the same LM for both the query and document encoding or have an equal number of parameters on both sides.
-Recent state-of-the-art (SOTA) retrievers Wang et al.
-(2024b); BehnamGhader et al.
-(2024) employ deeply parameterized LLMs as dual-encoders $Enc_{q}$ and $Enc_{d}$ for enhanced retrieval capabilities and domain adaptation abilities on various languages and tasks.
-However, serving giant LLMs slows down query inference throughput and depletes online deployment resources Kwon et al.
-(2023).
-Nowadays, LLM-based text retrieval incurs non-negligible online workloads, necessitating special optimizations for query inference efficiency.
-Modern Transformers-based LMs handle text modeling with the attention mechanism Vaswani et al.
-(2017), including the contextual traits and deep interactions of co-occurrences on discrete tokens within certain texts.
-As is the case for existing LLM-based text retrievers Wang et al.
-(2024b); BehnamGhader et al.
-(2024), queries and documents are modeled equally with the same LLM, enduring the same computational burdens.
-In contrast, traditional statistical retrieval algorithms, such as the popular BM25 Robertson et al.
-(1994), only characterize limited token co-occurrence by analyzing lexical overlaps with customized retrieval metrics.
-They do not handle the contextual traits and deep co-occurrences, and their encoders do not need to be deeply parametrized during inference.
-Experiments from Thakur et al.
-(2021) show that BM25 still performs well in many retrieval tasks.
-Built on the above observations, to minimize query service cost, we propose LightRetriever, a novel LLM-based hybrid retrieval architecture targeted at extreme lightweight query inference abilities.
-Our work retains a full-sized LLM on the document side ($Enc_{d}$) for strong retrieval capacities.
-But as is shown in Figure 1, we completely remove the deep modeling of queries.
-1) For dense retrieval, our method end-to-end trains cacheable token vectors from the LLM-based query encoder.
-The dense query representations are simply aggregated by averaging corresponding token vectors shown in the texts.
-All query token vectors are pre-computed and cached into a simple Embedding layer.
-Query inference costs are drastically reduced to only one Embedding lookup.
-2) For sparse retrieval, the mapping of a token and its corresponding frequency [^1] Here we directly use the number of a certain token that appeared in the tokenized query as its frequency.
-is directly used as the sparse query vector.
-Thus, no more encoding stage for sparse retrieval is needed.
-The dense similarity scores are computed with cosine similarity, while the sparse scores directly use inner product similarity.
-These scores are summed Wang et al.
-(2021) for the final hybrid retrieval.
+Most existing works of text retrieval [30, 11, 72] utilize symmetric dual-encoder architectures, which share the same LM for both the query and document encoding or have an equal number of parameters on both sides.
+Recent state-of-the-art (SOTA) retrievers [72, 5] employ deeply parameterized LLMs as dual-encoders $Enc_{q}$ and $Enc_{d}$ for enhanced retrieval capabilities and domain adaptation abilities.
+However, serving giant LLMs slows down query inference throughput and depletes online deployment resources [38].
+Despite the rapid progress on improving retrieval quality, little prior work has investigated the efficiency challenges of LLM-based retrievers, necessitating special optimizations for query inference efficiency.
+Modern Transformer-based LMs empower text retrievers with rich contextual modeling through the attention mechanism [69], enabling deep interactions across contextual tokens.
+This capability explains the success of recent LLM-based dual-encoders [72, 5], but it also means that both queries and documents incur the same computational burden.
+In contrast, traditional statistical methods such as BM25 [61] rely purely on lexical overlap and token frequency, requiring negligible inference cost while still remaining competitive in many retrieval tasks [66].
+This contrast highlights a practical tension: While document encoders can afford heavy pre-computation offline, query encoders must operate under strict online latency and resource constraints.
+Therefore, it is natural to ask: do queries truly need the same degree of deep contextual modeling as documents?
+To strike a balance between performances and efficiencies, we propose LightRetriever, a new hybrid retrieval architecture that explicitly breaks the symmetry between document and query encoders.
+Our key insight is that while documents benefit from the full modeling power of LLMs, queries do not require equally heavy processing to achieve strong retrieval performance.
+To this end, LightRetriever preserves a full-sized LLM encoder on the document side, but entirely removes the deep modeling on the query side.
+1) For dense retrieval, we end-to-end train cacheable token embeddings from the LLM-based query encoder and aggregate them through simple averaging, reducing query inference to a single embedding lookup; 2) For sparse retrieval, we directly map token counts to sparse vectors without any encoding.
+Dense and sparse scores are summed [73] for the final hybrid retrieval.
+LightRetriever achieves retrieval quality comparable to full dual-encoder LLMs, but with orders-of-magnitude faster query inference.
 LightRetriever is super fast and lightweight for query inference.
 Compared to serving a full-sized LLM on an A800 GPU, our method achieves over a thousand times speedup for query encoding, and over 10x total throughput improvements.
-Experiments on large-scale retrieval benchmarks demonstrate that our method generalizes well across diverse LLMs (including Llama-1B, 3B, 8B, and Qwen-1.5B, 3B, 7B) and retrieval benchmarks (including 15 English tasks from BeIR Thakur et al.
-(2021) and 8 Chinese tasks from CMTEB Retrieval Xiao et al.
-(2023)), maintaining an average of 95% retrieval performance.
+Experiments on large-scale retrieval benchmarks demonstrate that our method generalizes well across diverse LLMs (including Llama-1B, 3B, 8B, and Qwen-1.5B, 3B, 7B) and retrieval benchmarks (including 15 English tasks from BeIR [66] and 8 Chinese tasks from CMTEB Retrieval [77]), maintaining an average of 95% retrieval performance.
 2 Algorithm
@@
 A full-sized dual-encoder is used to model both queries and documents.
-Task-specific instructions are added as common practices Su et al.
-(2023); Wang et al.
-(2024b) to promote better domain adaptation abilities.
+Task-specific instructions are added as common practices [65, 72] to promote better domain adaptation abilities.
 2.1 LM-based text retrieval
@@
 $\min_{\theta}\max_{d^{-}\in\mathcal{N}(q)}\mathcal{E}(q,d^{+},d^{-};\theta)\text{, where }\mathcal{E}(q,d^{+},d^{-};\theta)=v_{q}\cdot v_{d^{-}}-v_{q}\cdot v_{d^{+}}$ (1)
-which lowers the upper boundary of ranking similarity errors $\mathcal{E}$ Chen et al.
-(2009).
-The above objective can be effectively optimized by the pairwise/listwise contrastive loss Cao et al.
-(2007), margin loss Rosset et al.
-(2003), etc.
-Following Karpukhin et al.
-(2020), listwise contrastive loss ($\ell^{CL}$) is used in our work,
+which lowers the upper boundary of ranking similarity errors $\mathcal{E}$ [12].
+The above objective can be effectively optimized by the pairwise/listwise contrastive loss [10], margin loss [62], etc.
+Following [30], listwise contrastive loss ($\ell^{CL}$) is used in our work,
 $\ell^{CL}=-\log\frac{e^{v_{q}\cdot v_{d^{+}}/\tau}}{e^{v_{q}\cdot v_{d^{+}}/\tau}+\sum_{d^{-}\in\mathcal{N}(q)}e^{v_{q}\cdot v_{d^{-}}/\tau}}.$ (2)
@@
 Instructed retrieval
-Similar to instructed fine-tuning (SFT) Ouyang et al.
-(2022), modern LM-based retrievers Su et al.
-(2023); Xiao et al.
-(2023); Wang et al.
-(2024a), especially LLM retrievers Wang et al.
-(2024b); BehnamGhader et al.
-(2024); Lee et al.
-(2024), customize task specific instructions for better domain adaption abilities.
+Similar to instructed fine-tuning (SFT) [56], modern LM-based retrievers [65, 77, 71], especially LLM retrievers [72, 5, 39], customize task specific instructions for better domain adaption abilities.
 Specifically, as shown in Figure 2, the task instruction is prepended to the queries before retrieval.
-Although Su et al.
-(2023); Xiao et al.
-(2023) use instructions on both query and passage sides, most LLM-based retrievers Wang et al.
-(2024b); BehnamGhader et al.
-(2024); Lee et al.
-(2024) do not use passage instructions to reuse the pre-built indexes across different tasks.
-In our work, dense retrieval follows the query instructions of Mistral-E5 Wang et al.
-(2024b) and does not involve passage instructions.
+Although [65, 77] use instructions on both query and passage sides, most LLM-based retrievers [72, 5, 39] do not use passage instructions to reuse the pre-built indexes across different tasks.
+In our work, dense retrieval follows the query instructions of Mistral-E5 [72] and does not involve passage instructions.
 The sparse retrieval is incompatible with instructions because it does not involve learnable query encoders.
@@
 Given a task instruction $Inst$ and query tokens $T_{q}=\{t_{0},t_{1},...,t_{n-1}\}$, each query token $t_{i}$ ($i\in\{0,...,n-1\}$) are prepended with the instruction.
-Then, the corresponding query token vectors $v_{t_{i}}^{\text{den}}$ are obtained by encoding [^2] Specifically, EOS pooling is used in our work by following Mistral-E5 Wang et al.
-(2024b).
+Then, the corresponding query token vectors $v_{t_{i}}^{\text{den}}$ are obtained by encoding [^1] Specifically, EOS pooling is used in our work by following Mistral-E5 [72].
 with the query encoder [^2] A customized causal mask is designed to avoid repeated prompt computations, detailed in Appendix A.7.,
-$v_{t_{i}}^{\text{den}}=Enc_{q}(Inst;t_{i})$ (3)
-Then, the dense query vector $v_{q}^{\text{den}}$ of the whole query text is aggregated by averaging token vectors,
-$v_{q}^{\text{den}}=\frac{1}{n}\sum_{i=0}^{n-1}v_{t_{i}}^{\text{den}}$ (4)
+$v_{q}^{\text{den}}=\frac{1}{n}\sum_{i=0}^{n-1}v_{t_{i}}^{\text{den}}\text{, where }v_{t_{i}}^{\text{den}}=Enc_{q}(Inst;t_{i})$ (3)
 Contrastive loss is the main objective for dense training, as defined in Equation 2.
@@
 These hidden states are then projected into the vocabulary space to get logits $w$ via the model’s own language modeling head, parameterized by a projection matrix $P$.
-Following SPLADE Formal et al.
-(2021), to induce sparsity and suppress dominant term frequencies, we apply a ReLU activation followed by a log-saturation function, and then perform a max-pooling operation of the logits $w$ over the sequence length dimension,
+To induce sparsity and suppress dominant term frequencies, we apply a ReLU activation followed by a log-saturation function, and then perform a max-pooling operation of the logits $w$ over the sequence length dimension by following [21],
 $v_{d}^{\text{spr}}=\max(w)\text{, where }w=\text{logits}=\ln(\max(h_{\text{last}}\cdot P,0)+1)$ (6)
 Here, $v_{d}^{\text{spr}}\in\mathbb{R}^{V}$ denotes the final sparse representation vector of the document.
-Given sparse vectors $v_{q}^{\text{spr}}$ and $v_{d}^{\text{spr}}$, the same contrastive loss in Equation 2 and auxiliary loss in Equation 9 are applied for representation learning.
+Given sparse vectors $v_{q}^{\text{spr}}$ and $v_{d}^{\text{spr}}$, the same contrastive loss in Equation 2 is applied for representation learning.
 $v_{d}^{\text{spr}}$ is not naturally sparse enough, because LM tends to generate dense distributions.
 Thus, the sparsify regulator is needed during the training phase.
-Following SPLADE Formal et al.
-(2021), given a batch of documents $d=\{d_{0},...,d_{i}\},i<\text{batch size }N$, the FLOPs regulator Paria et al.
-(2020) is used for sparsification.
+Given a batch of documents $d=\{d_{0},...,d_{i}\},i<\text{batch size }N$, the FLOPs regulator [57] is used for sparsification.
 This regulator first averages the logits across different documents within a certain batch, then computes the squared sum to get the final regulator.
@@
 Table 1: Comparison of retrieval speed and effectiveness.
-We compare the time consumption of retrieving 65,536 Bing queries over 1 million passages from the MS MARCO dataset Nguyen et al.
-(2016), evaluating both full symmetric retrievers and the query-lightweight LightRetriever.
+We compare the time consumption of retrieving 65,536 Bing queries over 1 million passages from the MS MARCO dataset [53], evaluating both full symmetric retrievers and the query-lightweight LightRetriever.
 To assess the trade-off between efficiency and effectiveness, we also report nDCG@10 scores on English (BeIR) and Chinese (CMTEB-Retrieval) benchmarks.
@@
 Our work finetunes on a large existing collection of 20 English and 3 Chinese datasets with 8.38M samples to ensure broad domain coverage and diversity.
-To ensure reproducibility, our work reuses the preprocessed fine-tuning data from tDRO Ma et al.
-(2025) and Sentence Transformers Training Data Reimers (2019).
-English datasets include Amazon Review Ni et al.
-(2019), Eli5 Fan et al.
-(2019), FEVER Thorne et al.
-(2018), FiQA Maia et al.
-(2018), GooQA Khashabi et al.
-(2021), HotpotQA Yang et al.
-(2018b), MSMARCO Nguyen et al.
-(2016), NFCorpus Boteva et al.
-(2016), NPR Lucas et al.
-(2023), NQ Kwiatkowski et al.
-(2019), PAQ Lewis et al.
-(2021), Quora Duplicates Iyer et al.
-(2012), S2ORC Lo et al.
-(2020), Scifact Wadden et al.
-(2020), Specter Cohan et al.
-(2020), StackExchange Duplicates da Silva et al.
-(2018), E5 synthetic BeastyZ (2024), Trivia Joshi et al.
-(2017), WikiHow Koupaee and Wang (2018), and Yahoo Answers Zhang et al.
-(2015).
-Chinese datasets include cMedQA2 Zhang et al.
-(2018), DuReader Qiu et al.
-(2022), and T2Ranking Xie et al.
-(2023).
+To ensure reproducibility, our work reuses the preprocessed fine-tuning data from tDRO [48] and Sentence Transformers Training Data [60].
+English datasets include Amazon Review [54], Eli5 [20], FEVER [67], FiQA [49], GooQA [31], HotpotQA [81], MSMARCO [53], NFCorpus [7], NPR [45], NQ [37], PAQ [41], Quora Duplicates [26], S2ORC [43], Scifact [70], Specter [13], StackExchange Duplicates [17], E5 synthetic [4], Trivia [29], WikiHow [34], and Yahoo Answers [83].
+Chinese datasets include cMedQA2 [82], DuReader [58], and T2Ranking [78].
 Benchmarks
-BeIR Thakur et al.
-(2021) and CMTEB Retrieval Xiao et al.
-(2023) benchmarks are used in the evaluation.
+BeIR [66] and CMTEB Retrieval [77] benchmarks are used in the evaluation.
 BeIR is a massive English retrieval benchmark collection with 15 evaluation sets.
@@
 CMTEB Retrieval is a massive Chinese retrieval benchmark collection with 8 sets, also covering a variety of domains.
-nDCG@10 is reported as the main metric by following Thakur et al.
-(2021); Xiao et al.
-(2023), accessing both recall and ranking abilities.
+nDCG@10 is reported as the main metric by following [66, 77], accessing both recall and ranking abilities.
 Recalls (R@{20, 50, 100}) are also reported to assess retrieval ability at large windows.
 Training hyper-parameters
-Our experiments are conducted based on the implementation of tDRO Ma et al.
-(2025).
+Our experiments are conducted based on the implementation of tDRO [48].
 All experiments are performed with a batch size of 128, 7 hard negatives, a max sequence length of 512, a contrastive temperature $\tau$ of 0.02, and 12k total steps.
-Following SPLADE Formal et al.
-(2021), the coefficient of the FLOPs regulator is 0.001, and the regulator is quadratically increased to the maximum at the first 4k steps, which helps reduce the side effects of the regulator at initial steps Paria et al.
-(2020).
-LoRA Hu et al.
-(2022) is used to save training GPU memory, where LoRA r is 16, alpha is 32, and dropout is 0.1.
-Different LLMs with different sizes are tested as the backbone encoders to verify the generalization abilities, including Llama-3.2-1B, 3B, Llama-3.1-8B Touvron et al.
-(2023), and Qwen-2.5-1.5B, 3B, 7B Bai et al.
-(2023).
-The hybrid similarity scores are linearly summed from normalized dense and sparse scores Wang et al.
-(2021).
+Following SPLADE [21], the coefficient of the FLOPs regulator is 0.001, and the regulator is quadratically increased to the maximum at the first 4k steps, which helps reduce the side effects of the regulator at initial steps [57].
+LoRA [25] is used to save training GPU memory, where LoRA r is 16, alpha is 32, and dropout is 0.1.
+Different LLMs with different sizes are tested as the backbone encoders to verify the generalization abilities, including Llama-3.2-1B, 3B, Llama-3.1-8B [68], and Qwen-2.5-1.5B, 3B, 7B [2].
+The hybrid similarity scores are linearly summed from normalized dense and sparse scores [73].
 All trainings are conducted on 8 NVIDIA H800 or A800 GPUs.
 Speed comparisons
-To evaluate the online speedup of LightRetriever, we compare the online time consumptions for retrieving 65536 Bing queries over 1M passages from MSMARCONguyen et al.
-(2016).
+To evaluate the online speedup of LightRetriever, we compare the online time consumptions for retrieving 65536 Bing queries over 1M passages from MSMARCO[53].
 All speed tests are conducted on a single A800 GPU, with a batch size of 256, a dimension of 1k for Faiss exact search, and 64 threads for Anserini (Lucene) sparse search.
@@
 Performance comparisons
-Table 2 compares retrieval effectiveness across a wide range of retrievers on both English (BeIR) and Chinese (CMTEB-Retrieval) benchmarks.
+Table 2 compares retrieval effectiveness on English (BeIR) and Chinese (CMTEB-Retrieval) benchmarks.
 The main focus of this comparison is to evaluate the performance degradation introduced by LightRetriever’s lightweight query encoding.
-Overall, by leveraging deep Transformer encoders on both query and document sides, full symmetric retrievers achieve the highest retrieval scores.
-For instance, Llama3.1-8b performs best on BeIR (56.8 nDCG@10), while Qwen2.5-7b reaches the best performance on CMTEB-R (70.1 nDCG@10).
-However, such performance comes at the cost of significant computational overhead, as discussed above.
+Overall, by leveraging deep Transformer encoders on both query and document sides, full symmetric retrievers with deep LLM encoders achieve the best scores (e.g., Llama3.1-8b on BeIR with 56.8 nDCG@10, Qwen2.5-7b on CMTEB-R with 70.1 nDCG@10), but at significant computational cost.
+LightRetriever achieves promising effectiveness while drastically reducing query-side computational burden.
+Across different backbones, it incurs only modest degradations (typically around 5% with 1–5 absolute points).
+For example, LightRetriever-Qwen2.5-7b achieves 53.8 nDCG@10 on BeIR and 66.5 nDCG@10 on CMTEB-R, only 2.8 and 3.6 lower than its full counterpart.
+Detailed task-level results are shown in Appendix Tables 911.
+Notably, LightRetriever retains strong performance on large-scale open-domain QA, fact-checking, and web search benchmarks such as FEVER [67], SciFact [70], TRECCOVID [16], and MSMARCO [53].
+On smaller, domain-specific datasets (e.g., FiQA [49] and CmedQA [82]), drops can be moderately larger (around 5 points) but remain tolerable, suggesting good generalization with room for future task-specific adaptation.
 […]
-These results demonstrate that LightRetriever effectively balances retrieval quality and efficiency.
-By avoiding expensive query-side encoding and maintaining high retrieval performance, it establishes a promising new paradigm for scalable LLM-based retrieval systems, especially in latency-sensitive or high-throughput scenarios.
 4 Ablations
@@
 In ablation studies below (A1-A3), we are interested in other potential symmetric/asymmetric architectures, which are described in the following Table 4.
-Table 3: Symmetry Ablation settings (A1-A3).
+Table 3: Symmetry Ablation settings (A1-A3). 1L-Trans denotes a single-layer Transformer.
 Train Inference
@@
 A2 MLP Full MLP Full
-A3 1-layer Transformers Full 1-layer Transformers Full
+A3 1L-Trans Full 1L-Trans Full
+Table 4: Ablations (A1-3) results (nDCG@10) of LightRetriever.
+Ablations BeIR CMTEB-R
+Llama-1b 48.7 58.7
+A1: Both-side Light 34.9-13.8 40.1-18.6
+A2: Encq use a MLP 37.5-11.2 41.3-17.4
+Llama-8b 54.4 63.0
+A3: Encq use 1L-Trans 50.1-4.3 54.4-8.6
 A1. Is symmetric lightweight inference effective?
@@
 This setup significantly degrades performance (–11.2 on BeIR, –17.4 on CMTEB-R), indicating that the full-sized modeling for query side training remains essential during training to learn effective token-level representations.
-Table 4: Ablations results of LightRetriever
-Ablations / nDCG@10 BeIR CMTEB-R
-Llama-1b 48.7 58.7
-w/ Both-side Lightweight 34.9-13.8 40.1-18.6
-w/ Encq use a MLP 37.5-11.2 41.3-17.4
-Llama-8b 54.4 63.0
-w/ 1st-layer on query 50.1-4.3 54.4-8.6
 A3. Is adding one Transformers layer better on the query side?
@@
 A.2.1 Text retrieval
-Text retrieval builds on top of the vector space model Salton [1975], which encodes query and document representation vectors, and performs relevance search based on vector similarities.
+Text retrieval builds on top of the vector space model [63], which encodes query and document representation vectors, and performs relevance search based on vector similarities.
 Dense retrieval
-Dense retrieval Karpukhin et al.
-[2020] trains PLM-based encoders to condense queries and passages into low-dimensional dense vectors, then performs relevance search based on the Maximum Inner Product Search (MIPS) Mussmann and Ermon [2016] algorithm.
-In recent years, dense retrievers, such as BGE Chen et al.
-[2024] and E5 Wang et al.
-[2024a], have gained popularity for strong retrieval abilities across different tasks Muennighoff et al.
-[2023], Thakur et al.
-[2021], languages Zhang et al.
-[2023], and context granulaties Williams et al.
-[2018], Nguyen et al.
-[2016], owing to diverse training data Reimers [2019], retrieval-customized pre-training Gao and Callan [2021], Xiao et al.
-[2022], Ma et al.
-[2024], improved negative-mining Xiong et al.
-[2021], and so on.
-Recent SOTA retrievers Wang et al.
-[2024b], BehnamGhader et al.
-[2024], Lee et al.
-[2024] start to utilize LLMs Jiang et al.
-[2023], Touvron et al.
-[2023] as backbone encoders.
+Dense retrieval [30] trains PLM-based encoders to condense queries and passages into low-dimensional dense vectors, then performs relevance search based on the Maximum Inner Product Search (MIPS) [52] algorithm.
+In recent years, dense retrievers, such as BGE [11] and E5 [71], have gained popularity for strong retrieval abilities across different tasks [51, 66], languages [84], and context granulaties [75, 53], owing to diverse training data [60], retrieval-customized pre-training [22, 76, 47], improved negative-mining [79], and so on.
+Recent SOTA retrievers [72, 5, 39] start to utilize LLMs [27, 68] as backbone encoders.
 With larger parameters and pre-training data, LLM-based retrievers enable significantly powerful retrieval capacities and domain adaptation abilities.
@@
 Sparse retrieval
-Traditional statistical sparse retrieval algorithms (e.g., TF-IDF Salton and Buckley [1988], and BM25 Robertson et al.
-[1994]) do not involve learnable language models.
+Traditional statistical sparse retrieval algorithms (e.g., TF-IDF [64], and BM25 [61]) do not involve learnable language models.
 Instead, they directly analyze lexical frequencies (e.g., TF) and importance (e.g., IDF) in a large vocabulary space.
-These lexicals are sparsely distributed and require building inverted indexes Luhn [1957] for feasible similarity metric calculation, such as Apache Lucene Bialecki et al.
-[2012] and Tantivy Quickwit Inc.
-[2024].
+These lexicals are sparsely distributed and require building inverted indexes [46] for feasible similarity metric calculation, such as Apache Lucene [6] and Tantivy [59].
 In recent years, sparse retrieval has started to work with pre-trained LMs (PLMs).
-BGE Chen et al.
-[2024], SPLADE Formal et al.
-[2021], SparTerm Bai et al.
-[2020], and SPARTA Zhao et al.
-[2021] use customized XLM-RoBERTa Conneau et al.
-[2020] or BERT Devlin et al.
-[2019] as backbone encoders to learn term frequencies in an end-to-end manner and directly perform impact search based on the dot product of sparse vectors.
+BGE [11], SPLADE [21], SparTerm [3], and SPARTA [85] use customized XLM-RoBERTa [14] or BERT [19] as backbone encoders to learn term frequencies in an end-to-end manner and directly perform impact search based on the dot product of sparse vectors.
 Because of sparsity, these works also search with inverted indexes.
-Recent works Zhuang et al.
-[2024] also try to incorporate LLMs with sparse retrieval.
+Recent works [86] also try to incorporate LLMs with sparse retrieval.
 However, the online efficiency issue still exists.
 Hybrid retreval
-Hybrid retrieval is a technique that combines retrieval scores Wang et al.
-[2021] or ranks Cormack et al.
-[2009] from multiple retrieval systems, such as dense Karpukhin et al.
-[2020] / sparse Robertson et al.
-[1994], Formal et al.
-[2021] retrieval, and single Karpukhin et al.
-[2020] / multiple Khattab and Zaharia [2020], Chen et al.
-[2024] vector retrieval.
+Hybrid retrieval is a technique that combines retrieval scores [73] or ranks [15] from multiple retrieval systems, such as dense [30] / sparse [61, 21] retrieval, and single [30] / multiple [32, 11] vector retrieval.
 It enables better retrieval abilities by interpolating results from multiple sources.
-Linear interpolation Wang et al.
-[2021] of dense and sparse scores is used in our work.
+Linear interpolation [73] of dense and sparse scores is used in our work.
 A.2.2 Improving inference efficiency
@@
 Previous works have tried to distill small query encoders from large encoders for better query inference efficiency.
-KALE Campos et al.
-[2023] prunes the BERT-based query encoder and distills it during post-training.
-However, when the query encoder is pruned to one Transformers layer, its recall@20 on NQ Kwiatkowski et al.
-[2019] decreases by 25-30%.
-Wang and Lyu [2023] proposes asymmetric dual-encoders by pruning BERT layers or using smaller BERT models.
+KALE [9] prunes the BERT-based query encoder and distills it during post-training.
+However, when the query encoder is pruned to one Transformers layer, its recall@20 on NQ [37] decreases by 25-30%.
+[74] proposes asymmetric dual-encoders by pruning BERT layers or using smaller BERT models.
 Its experiments show that a 2-layer BERT query encoder retains 92.5% of the full-sized performances via distillation and proper student initialization.
-USTAD Kim et al.
-[2024] also proposes similar asymmetry and distillation strategies, which choose the 12-layer BERT as teachers and the 6-layer DistilBERT or 2-layer BERT-mini as students.
+USTAD [33] also proposes similar asymmetry and distillation strategies, which choose the 12-layer BERT as teachers and the 6-layer DistilBERT or 2-layer BERT-mini as students.
 Its experiments show that the distilled BERT-mini query encoder achieves 95-97% of the teacher’s performance because of better distillation.
-However, these studies have primarily focused on small language models (SLMs), such as BERT Devlin et al.
-[2019], and overlooked feasible approaches for improving online inference efficiency of LLM-based retrievers, where such optimization is even more critical.
+However, these studies have primarily focused on small language models (SLMs), such as BERT [19], and overlooked feasible approaches for improving online inference efficiency of LLM-based retrievers, where such optimization is even more critical.
 Furthermore, these existing methods typically retain Transformer layers in the query encoders, which limits their efficiency gains.
-In addition, most of these works rely on SLMs and are trained on relatively narrow datasets, such as MS-MARCO Nguyen et al.
-[2016], thus unable to compete with LLM-based retrievers.
+In addition, most of these works rely on SLMs and are trained on relatively narrow datasets, such as MS-MARCO [53], thus unable to compete with LLM-based retrievers.
 In contrast, with the rapid advancement of LLM-based text retrievers, driven by larger and more diverse foundation models and datasets, retrieval performance has significantly improved.
@@
 Previous works have also been dedicated to reducing index sizes for smaller disk space usage.
-Matryoshka representation learning (MRL) Kusupati et al.
-[2022] supports adjusting the dense vector dimensions by directly taking the top-k dimension of the original vector as shrink vectors.
+Matryoshka representation learning (MRL) [36] supports adjusting the dense vector dimensions by directly taking the top-k dimension of the original vector as shrink vectors.
 It aligns and trains the corresponding multi-dimensional vectors in the end-to-end contrastive learning.
 For example, if $k\in\{64,128,256\}$, query-passage dense vectors with $\{64,128,256\}$-dims could be obtained together after one encoder forward operation.
-For LM-based sparse retrieval, the similarity scores are directly affected by the impacts of corresponding lexicals Formal et al.
-[2021].
+For LM-based sparse retrieval, the similarity scores are directly affected by the impacts of corresponding lexicals [21].
 Thus, it naturally supports reducing the index sizes by retaining Top-k terms/lexicals.
@@
 Retrieval systems usually involve indexing millions of documents into search databases, which consumes large amounts of disk space, especially for dense vectors.
-For example, indexing 8.8M MS-MARCO passages Nguyen et al.
-[2016] with Llama-8b model consumes 134.92 GB for Faiss dense retrieval and 11.24 GB for Lucene sparse retrieval.
+For example, indexing 8.8M MS-MARCO passages [53] with Llama-8b model consumes 134.92 GB for Faiss dense retrieval and 11.24 GB for Lucene sparse retrieval.
 Thus, it’s essential to explore the possibility of controlling index sizes by shrinking vectors.
-1) For dense retrieval, we incorporate Matryoshka Representation Learning (MRL) Kusupati et al.
-[2022] with LightRetriever to support flexible controls of dense vector dimensions.
+1) For dense retrieval, we incorporate Matryoshka Representation Learning (MRL) [36] with LightRetriever to support flexible controls of dense vector dimensions.
 Given a dense vector $v^{den}$ with dimension $H$, MRL derives dimension-shrunk vectors by simply cutting out the Top-k dimensions $v^{den}[:k]$, then applies the same training loss as the original vectors.
@@
 Hybrid 54.4 63.0 53.9-0.5 62.1-0.9
-Previous works for BERT-based asymmetric retrievers Wang and Lyu [2023], Kim et al.
-[2024] emphasize the importance of alignment between asymmetric students and full symmetric teachers.
+Previous works for BERT-based asymmetric retrievers [74, 33] emphasize the importance of alignment between asymmetric students and full symmetric teachers.
 To explore the effect of such alignment, besides contrastive loss as the main loss function, we also explore to adopt the KL loss as an auxiliary to align the asymmetric similarity scores $S^{student}=v_{q}^{den}\cdot v_{d}^{den}$ to full-sized symmetric similarity scores $S^{teacher}=Enc_{q}(q)\cdot Enc_{d}(d)$,
@@
 Dataset infos
-Following the settings of previous LLM-based retrievers Wang et al.
-[2024b], BehnamGhader et al.
-[2024], Ma et al.
-[2025], our work uses 23 training sets pre-built by tDRO Ma et al.
-[2025] or Sentence Transformers Training Data Reimers [2019], including 20 English and 3 Chinese datasets.
+Following the settings of previous LLM-based retrievers [72, 5, 48], our work uses 23 training sets pre-built by tDRO [48] or Sentence Transformers Training Data [60], including 20 English and 3 Chinese datasets.
 Detailed dataset information is listed in Table 6 as follows.
@@
 Dataset Language Category Deduped Size Epoch Ratio
-Amazon Review (2018) Ni et al. [2019] English Amazon $999$ 0.01 0.58%
-cMedQA2 Zhang et al. [2018] Chinese Chinese Medical 99936 1 5.84%
-DuReader Qiu et al. [2022] Chinese Chinese Web Collections 86366 1 5.05%
-Eli5 Fan et al. [2019] English Reddit 325390 0.05 0.95%
-Fever Thorne et al. [2018] English Wikipedia QA 109808 0.8 5.13%
-FiQA Maia et al. [2018] English Financial 5498 1 0.32%
-GooQA Pairs Khashabi et al. [2021] English Web Collections 3012347 0.025 4.40%
-HotpotQA Yang et al. [2018b] English Wikipedia QA 85000 0.5 2.48%
-MSMARCO Nguyen et al. [2016] English Web Collections 502854 1 29.39%
-NFCorpus Boteva et al. [2016] English Medical 2585 1 0.15%
-NPR Lucas et al. [2023] English News 594376 0.05 1.74%
-NQ Kwiatkowski et al. [2019] English Wikipedia QA 58800 1 3.44%
-PQA Pairs Lewis et al. [2021] English Wikipedia QA 99999 0.5 2.92%
-Quora Duplicates Triples Iyer et al. [2012] English Forum Duplicates 97011 0.5 2.83%
-S2ORC Title-Abstract Lo et al. [2020] English Semantic Scholar 99998 0.5 2.92%
-SciFact Wadden et al. [2020] English S2ORC 806 1 0.05%
-SPECTER Cohan et al. [2020] English Semantic Scholar 136642 0.25 2.00%
-StackExchange Duplicates (Title-Body) da Silva et al. [2018] English Forum Duplicates 250516 0.25 3.66%
-E5 Synthetic BeastyZ [2024] English GPT Synthetic Data 224791 1 13.14%
-T2Ranking Xie et al. [2023] Chinese Chinese Web Collections 200362 0.5 5.86%
-Trivia Joshi et al. [2017] English Wikipedia QA 60370 0.5 1.76%
-Wikihow Koupaee and Wang [2018] English WikiHow 128543 0.25 1.88%
-Yahoo Answers (Title-Answer) Zhang et al. [2015] English Yahoo 1198018 0.05 3.50%
-Many downstream retrieval benchmarks Thakur et al.
-[2021], Xiao et al.
-[2023] are evaluated out-of-domain, where no training samples from the downstream sources are available.
-To avoid the overfitting of training sets, as is shown in the above Table 6, we slightly tune the epoch used in the dataset sampling, by following previous works in Ni et al.
-[2022], Wang et al.
-[2024b], Ma et al.
-[2025].
+Amazon Review (2018) [54] English Amazon $999$ 0.01 0.58%
+cMedQA2 [82] Chinese Chinese Medical 99936 1 5.84%
+DuReader [58] Chinese Chinese Web Collections 86366 1 5.05%
+Eli5 [20] English Reddit 325390 0.05 0.95%
+Fever [67] English Wikipedia QA 109808 0.8 5.13%
+FiQA [49] English Financial 5498 1 0.32%
+GooQA Pairs [31] English Web Collections 3012347 0.025 4.40%
+HotpotQA [81] English Wikipedia QA 85000 0.5 2.48%
+MSMARCO [53] English Web Collections 502854 1 29.39%
+NFCorpus [7] English Medical 2585 1 0.15%
+NPR [45] English News 594376 0.05 1.74%
+NQ [37] English Wikipedia QA 58800 1 3.44%
+PQA Pairs [41] English Wikipedia QA 99999 0.5 2.92%
+Quora Duplicates Triples [26] English Forum Duplicates 97011 0.5 2.83%
+S2ORC Title-Abstract [43] English Semantic Scholar 99998 0.5 2.92%
+SciFact [70] English S2ORC 806 1 0.05%
+SPECTER [13] English Semantic Scholar 136642 0.25 2.00%
+StackExchange Duplicates (Title-Body) [17] English Forum Duplicates 250516 0.25 3.66%
+E5 Synthetic [4] English GPT Synthetic Data 224791 1 13.14%
+T2Ranking [78] Chinese Chinese Web Collections 200362 0.5 5.86%
+Trivia [29] English Wikipedia QA 60370 0.5 1.76%
+Wikihow [34] English WikiHow 128543 0.25 1.88%
+Yahoo Answers (Title-Answer) [83] English Yahoo 1198018 0.05 3.50%
+Many downstream retrieval benchmarks [66, 77] are evaluated out-of-domain, where no training samples from the downstream sources are available.
+To avoid the overfitting of training sets, as is shown in the above Table 6, we slightly tune the epoch used in the dataset sampling, by following previous works in [55, 72, 48].
 Training set decontamination
@@
 Existing training datasets are derived from multiple sources, whose training data have possible duplicates with the test sets.
-Following tDRO Ma et al.
-[2025], we perform strict training set decontamination with SimHash Manku et al.
-[2007] to ensure no training queries appear in the BeIR Thakur et al.
-[2021] and CMTEB Retrieval Xiao et al.
-[2023].
+Following tDRO [48], we perform strict training set decontamination with SimHash [50] to ensure no training queries appear in the BeIR [66] and CMTEB Retrieval [77].
 The original sizes, deduplicated sizes, and duplicate numbers as listed in the Table 7.
@@
 A.6 Task instructions
-As previously mentioned in Section 2.1, modern LLM-based retrievers Wang et al.
-[2024b], BehnamGhader et al.
-[2024] introduce task-specific instructions to format input queries, which enables higher retrieval capacities.
-These LLM-based instruction-tuned retrieval is similar to instruction-based supervised fine-tuning (SFT) in LLM generation Ouyang et al.
-[2022].
-Following Mistral-E5 Wang et al.
-[2024b], we reuse most of its task instructions, and format the instruction and query in the following format:
+As previously mentioned in Section 2.1, modern LLM-based retrievers [72, 5] introduce task-specific instructions to format input queries, which enables higher retrieval capacities.
+These LLM-based instruction-tuned retrieval is similar to instruction-based supervised fine-tuning (SFT) in LLM generation [56].
+Following Mistral-E5 [72], we reuse most of its task instructions, and format the instruction and query in the following format:
 […]
-The detailed instructions are listed below, most of which are directly copied from the previous works E5-Mistral Wang et al.
-[2024b] and tDRO Ma et al.
-[2025] for reproducibility.
+The detailed instructions are listed below, most of which are directly copied from the previous works E5-Mistral [72] and tDRO [48] for reproducibility.
 Note that one training set may correspond to multiple instructions to avoid overfitting.
@@
 , where the vocabulary dimension $V$ is much larger than the hidden dimension $H$.
-For example, for Llama3.1-8b Touvron et al.
-[2023] model, $V=128256$ and $H=4096$, where $\lfloor V/H\rfloor=31$.
+For example, for Llama3.1-8b [68] model, $V=128256$ and $H=4096$, where $\lfloor V/H\rfloor=31$.
 This means that, given batch size B, sequence length T, hidden size H, vocabulary size V, the intermediate tensors with shape $[B,T,V]$ in the projection, ReLU, and Log are 31x larger than a dense tensor $[B,T,H]$, which consumes huge amounts of GPU memory during training and inference.
@@
 Proper optimizations are important.
-For example, we found it hard to reproduce quickly with some of the open-sourced baselines BehnamGhader et al.
-[2024], which is roughly 6x slower than our implementation when testing with the same size of LLM.
+For example, we found it hard to reproduce quickly with some of the open-sourced baselines [5], which is roughly 6x slower than our implementation when testing with the same size of LLM.
 Our experiments involve training with 23 train sets and inference with another 23 test sets, which requires proper optimizations urgently.
-We import other optimizations based on the implementations  [^4] Please refer to Ma et al.
-[2025]’s implementation on https://github.com/ma787639046/tdro.
-released by tDRO Ma et al.
-[2025].
-These optimizations include 1) A distributed multi-node inference framework based on PyTorch RPC Li et al.
-[2023]; 2) Sequence packing Kundu et al.
-[2024] with Flash Attention 2 Dao [2023], which gives roughly 50% training speedup due to the removal of the pad tokens.; 3) Fused Kernels from Liger-Kernel Hsu et al.
-[2024], which gives roughly 10% speedup to both training and inferencing; 4) Multi-vector gradient accumulation support with GradCache Gao et al.
-[2021] for optionally enlarging batch sizes if necassary.
+We import other optimizations based on the implementations  [^3] Please refer to [48]’s implementation on https://github.com/ma787639046/tdro.
+released by tDRO [48].
+These optimizations include 1) A distributed multi-node inference framework based on PyTorch RPC [42]; 2) Sequence packing [35] with Flash Attention 2 [18], which gives roughly 50% training speedup due to the removal of the pad tokens.; 3) Fused Kernels from Liger-Kernel [24], which gives roughly 10% speedup to both training and inferencing; 4) Multi-vector gradient accumulation support with GradCache [23] for optionally enlarging batch sizes if necassary.
 A.8 Detailed main results

### v5

@@
 1 Introduction
-Given search queries and candidate documents, text retrieval employs dual-encoder architectures to encode representation vectors and search documents relevant to queries based on query-document similarities [63].
+Given search queries and candidate documents, text retrieval employs dual-encoder architectures to encode representation vectors and search documents relevant to queries based on query-document similarities (Salton et al., 1975).
 The encoded vectors could be condensed into low-dimensional vectors for dense retrieval or sparsified into high-dimensional vectors for sparse retrieval.
-Dense retrieval [30] compresses textual information into dense vectors with parameterized language models (LMs) and retrieves based on inner product similarity.
-Meanwhile, sparse retrieval [61] characterizes sparse term frequencies on a much broader vocabulary space, and then searches based on lexical overlap and term-based relevance metrics, such as direct impact search (Inner product) [3, 21] and BM25 [61].
-Additionally, hybrid retrieval enhances retrieval abilities by interpolating their similarity scores [73] or ranks [15].
-Text retrieval is broadly applied in various scenarios, such as web search [44], semantic textual similarity [8], fact checking [67], and retrieval augmented generation [40].
+Dense retrieval (Karpukhin et al., 2020) compresses textual information into dense vectors with parameterized language models (LMs) and retrieves based on inner product similarity.
+Meanwhile, sparse retrieval (Robertson et al., 1994) characterizes sparse term frequencies on a much broader vocabulary space, and then searches based on lexical overlap and term-based relevance metrics, such as direct TF-based search (Inner product) (Bai et al., 2020; Formal et al., 2021) and BM25 (Robertson et al., 1994).
+Additionally, hybrid retrieval enhances retrieval abilities by interpolating their similarity scores (Wang et al., 2021) or ranks (Cormack et al., 2009).
+Text retrieval is broadly applied in various scenarios, such as web search (Lu et al., 2022), semantic textual similarity (Bowman et al., 2015), fact checking (Thorne et al., 2018), and retrieval augmented generation (Lewis et al., 2020).
 Figure 1: LightRetriever targets at extreme query inference speed for LLM-based text retrieval, reducing the workload of query encoding to no more than an embedding lookup.
 Deploying retrieval systems requires different processing workflows for queries and documents.
-As is shown in Figure 1, as search candidates, documents are encoded with a document encoder $Enc_{d}$ and then indexed into search databases [28, 61].
-Because documents are obtained before the search operation, they are pre-computable and distributable, which can be indexed offline and distributed anywhere [80].
+As shown in Figure 1, as search candidates, documents are encoded with a document encoder $Enc_{d}$ and then indexed into search databases (Johnson et al., 2021; Robertson et al., 1994).
+Because documents are obtained before the search operation, they are pre-computable and distributable, which can be indexed offline and distributed anywhere (Yang et al., 2018a).
 On the contrary, queries are input and encoded in real-time, which are non-pre-computable and non-distributable.
@@
 Special GPU accelerators are mandatory if the query encoder is a deeply parameterized LM.
-Most existing works of text retrieval [30, 11, 72] utilize symmetric dual-encoder architectures, which share the same LM for both the query and document encoding or have an equal number of parameters on both sides.
-Recent state-of-the-art (SOTA) retrievers [72, 5] employ deeply parameterized LLMs as dual-encoders $Enc_{q}$ and $Enc_{d}$ for enhanced retrieval capabilities and domain adaptation abilities.
-However, serving giant LLMs slows down query inference throughput and depletes online deployment resources [38].
+Recent state-of-the-art (SOTA) retrievers (Wang et al., 2024a; BehnamGhader et al., 2024) employ deeply parameterized LLMs as symmetric dual-encoders $Enc_{q}$ and $Enc_{d}$ for enhanced retrieval capabilities.
+However, serving giant LLMs slows down query inference throughput and depletes online deployment resources (Kwon et al., 2023).
 Despite the rapid progress on improving retrieval quality, little prior work has investigated the efficiency challenges of LLM-based retrievers, necessitating special optimizations for query inference efficiency.
-Modern Transformer-based LMs empower text retrievers with rich contextual modeling through the attention mechanism [69], enabling deep interactions across contextual tokens.
-This capability explains the success of recent LLM-based dual-encoders [72, 5], but it also means that both queries and documents incur the same computational burden.
-In contrast, traditional statistical methods such as BM25 [61] rely purely on lexical overlap and token frequency, requiring negligible inference cost while still remaining competitive in many retrieval tasks [66].
+Modern Transformer-based LMs empower text retrievers with rich contextual modeling through the attention mechanism (Vaswani et al., 2017), enabling deep interactions across contextual tokens.
+This capability explains the success of recent LLM-based dual-encoders (Wang et al., 2024a; BehnamGhader et al., 2024), but it also means that both queries and documents incur the same computational burden.
+In contrast, traditional statistical methods such as BM25 (Robertson et al., 1994) rely purely on lexical overlap and token frequency, requiring negligible inference cost while still remaining competitive in many retrieval tasks (Thakur et al., 2021).
 This contrast highlights a practical tension: While document encoders can afford heavy pre-computation offline, query encoders must operate under strict online latency and resource constraints.
@@
 1) For dense retrieval, we end-to-end train cacheable token embeddings from the LLM-based query encoder and aggregate them through simple averaging, reducing query inference to a single embedding lookup; 2) For sparse retrieval, we directly map token counts to sparse vectors without any encoding.
-Dense and sparse scores are summed [73] for the final hybrid retrieval.
-LightRetriever achieves retrieval quality comparable to full dual-encoder LLMs, but with orders-of-magnitude faster query inference.
-LightRetriever is super fast and lightweight for query inference.
+Dense and sparse scores are linearly interpolated (Wang et al., 2021) for the final hybrid retrieval.
+LightRetriever achieves retrieval quality mostly comparable to full dual-encoder LLMs, but with orders-of-magnitude faster query inference.
 Compared to serving a full-sized LLM on an A800 GPU, our method achieves over a thousand times speedup for query encoding, and over 10x total throughput improvements.
-Experiments on large-scale retrieval benchmarks demonstrate that our method generalizes well across diverse LLMs (including Llama-1B, 3B, 8B, and Qwen-1.5B, 3B, 7B) and retrieval benchmarks (including 15 English tasks from BeIR [66] and 8 Chinese tasks from CMTEB Retrieval [77]), maintaining an average of 95% retrieval performance.
+Importantly, our experiments show that adopting a lightweight query encoder does not mean queries are deprived of deep semantic understanding.
+Our work tries to transfer the major modeling cost from the query side to the document side, while still preserving the benefits of LLM-powered relevance-based similarity search with query-document interactions.
+Experiments on large-scale retrieval benchmarks demonstrate that our method generalizes well across diverse LLMs (including Llama-1B, 3B, 8B, and Qwen-1.5B, 3B, 7B) and retrieval benchmarks (including 15 English tasks from BeIR (Thakur et al., 2021) and 8 Chinese tasks from CMTEB Retrieval (Xiao et al., 2024)), maintaining an average of 95% retrieval performance.
 2 Algorithm
@@
 A full-sized dual-encoder is used to model both queries and documents.
-Task-specific instructions are added as common practices [65, 72] to promote better domain adaptation abilities.
+Task-specific instructions are added as common practices (Su et al., 2023; Wang et al., 2024a) to promote better domain adaptation abilities.
 2.1 LM-based text retrieval
@@
 $\min_{\theta}\max_{d^{-}\in\mathcal{N}(q)}\mathcal{E}(q,d^{+},d^{-};\theta)\text{, where }\mathcal{E}(q,d^{+},d^{-};\theta)=v_{q}\cdot v_{d^{-}}-v_{q}\cdot v_{d^{+}}$ (1)
-which lowers the upper boundary of ranking similarity errors $\mathcal{E}$ [12].
-The above objective can be effectively optimized by the pairwise/listwise contrastive loss [10], margin loss [62], etc.
-Following [30], listwise contrastive loss ($\ell^{CL}$) is used in our work,
+which lowers the upper boundary of ranking similarity errors $\mathcal{E}$ (Chen et al., 2009).
+The above objective can be effectively optimized by contrastive loss (Cao et al., 2007), margin loss (Rosset et al., 2003), etc.
+Following (Karpukhin et al., 2020), listwise contrastive loss ($\ell^{CL}$) is used in our work,
 $\ell^{CL}=-\log\frac{e^{v_{q}\cdot v_{d^{+}}/\tau}}{e^{v_{q}\cdot v_{d^{+}}/\tau}+\sum_{d^{-}\in\mathcal{N}(q)}e^{v_{q}\cdot v_{d^{-}}/\tau}}.$ (2)
@@
 Instructed retrieval
-Similar to instructed fine-tuning (SFT) [56], modern LM-based retrievers [65, 77, 71], especially LLM retrievers [72, 5, 39], customize task specific instructions for better domain adaption abilities.
+Similar to instructed fine-tuning (SFT) (Ouyang et al., 2022), modern LM-based retrievers (Su et al., 2023; Xiao et al., 2024; Wang et al., 2024b), especially LLM retrievers (Wang et al., 2024a; BehnamGhader et al., 2024; Lee et al., 2024), customize task specific instructions for better domain adaption abilities.
 Specifically, as shown in Figure 2, the task instruction is prepended to the queries before retrieval.
-Although [65, 77] use instructions on both query and passage sides, most LLM-based retrievers [72, 5, 39] do not use passage instructions to reuse the pre-built indexes across different tasks.
-In our work, dense retrieval follows the query instructions of Mistral-E5 [72] and does not involve passage instructions.
+Although (Su et al., 2023; Xiao et al., 2024) use instructions on both query and passage sides, most LLM-based retrievers (Wang et al., 2024a; BehnamGhader et al., 2024; Lee et al., 2024) do not use passage instructions to reuse the pre-built indexes across different tasks.
+In our work, dense retrieval follows the query instructions of Mistral-E5 (Wang et al., 2024a) and does not involve passage instructions.
 The sparse retrieval is incompatible with instructions because it does not involve learnable query encoders.
@@
 Given a task instruction $Inst$ and query tokens $T_{q}=\{t_{0},t_{1},...,t_{n-1}\}$, each query token $t_{i}$ ($i\in\{0,...,n-1\}$) are prepended with the instruction.
-Then, the corresponding query token vectors $v_{t_{i}}^{\text{den}}$ are obtained by encoding [^1] Specifically, EOS pooling is used in our work by following Mistral-E5 [72].
-with the query encoder [^2] A customized causal mask is designed to avoid repeated prompt computations, detailed in Appendix A.7.,
+Then, the corresponding query token vectors $v_{t_{i}}^{\text{den}}$ are obtained via last token pooling with the query encoder [^1] A customized causal mask is designed to avoid repeated prompt computations, detailed in Appendix A.8..
+Then, the dense query vector $v_{q}^{\text{den}}$ of the whole query text is aggregated by averaging token vectors,
 $v_{q}^{\text{den}}=\frac{1}{n}\sum_{i=0}^{n-1}v_{t_{i}}^{\text{den}}\text{, where }v_{t_{i}}^{\text{den}}=Enc_{q}(Inst;t_{i})$ (3)
@@
 Given a certain vocabulary with size $V$, all token vectors $v_{t}^{den},t\in\{0,...,V-1\}$ are precomputed as in Equation 3 and stored in a lookup Embedding $E$.
+Typically, caching such a lookup Embedding on 8 H800 with a Llama-8b model consumes less than 20s, which is a one-go and cheap operation.
+In Appendix A.4, LightRetriever also shows good ability to balance the serving Embedding size via simple dimension truncations (Kusupati et al., 2022).
 Online serving
@@
 $H$ is the dimension of LLM’s hidden states, which is also the dimension of dense vectors.
-The dense query vector is obtained with a simple lookup-then-average operation,
+The dense query vector is obtained with a simple lookup-then-average operation, which is super efficient with or without GPU acceleration.
 $v_{q}^{\text{den}}=\frac{1}{n}\sum_{i=0}^{n-1}v_{t_{i}}^{den}=\frac{1}{n}\sum_{i=0}^{n-1}E[t_{i}]$ (4)
-The online serving of our dense retriever is extremely lightweight, because such Embedding lookup is super efficient with or without GPU hardware.
 2.3 Sparse retrieval of LightRetriever
@@
 These hidden states are then projected into the vocabulary space to get logits $w$ via the model’s own language modeling head, parameterized by a projection matrix $P$.
-To induce sparsity and suppress dominant term frequencies, we apply a ReLU activation followed by a log-saturation function, and then perform a max-pooling operation of the logits $w$ over the sequence length dimension by following [21],
+To induce sparsity and suppress dominant term frequencies, we apply a ReLU activation followed by a log-saturation function, and then perform a max-pooling operation of the logits $w$ over the sequence length dimension by following (Formal et al., 2021),
 $v_{d}^{\text{spr}}=\max(w)\text{, where }w=\text{logits}=\ln(\max(h_{\text{last}}\cdot P,0)+1)$ (6)
@@
 Thus, the sparsify regulator is needed during the training phase.
-Given a batch of documents $d=\{d_{0},...,d_{i}\},i<\text{batch size }N$, the FLOPs regulator [57] is used for sparsification.
+Given a batch of documents $d=\{d_{0},...,d_{i}\},i<\text{batch size }N$, the FLOPs regulator (Paria et al., 2020) is used for sparsification.
 This regulator first averages the logits across different documents within a certain batch, then computes the squared sum to get the final regulator.
@@
 The sparse query vector $v_{q}^{\text{spr}}$ during online serving follows the same rule in Equation 5. No more query encoder for sparse retrieval is needed.
-Table 1: Comparison of retrieval speed and effectiveness.
-We compare the time consumption of retrieving 65,536 Bing queries over 1 million passages from the MS MARCO dataset [53], evaluating both full symmetric retrievers and the query-lightweight LightRetriever.
-To assess the trade-off between efficiency and effectiveness, we also report nDCG@10 scores on English (BeIR) and Chinese (CMTEB-Retrieval) benchmarks.
-For retrieval efficiency, we report the consumed time of: query tokenization, model encoding, maximum search time of Faiss/Lucene, and total end-to-end retrieval time.
-The overall throughput is measured in Queries Per Second (QPS).
-Benchmark / nDCG@10 Time consumption / s & Throughput / QPS
-Model BeIR CMTEB-R Tokenize Encode Search Total QPS
-Full-Llama8b 56.8 67.6 1.3746 109.4853 8.5133 119.3730 549
-Full-Llama1b 53.1 63.8 1.2564 19.9030 8.4897 29.6490 2210
-LightRetriever-Llama8b 54.4 63.0 0.8209 0.0412 8.5010 9.3630 6999
-Full-Qwen7b 56.6 70.1 1.4144 100.6716 8.5279 110.6140 592
-Full-Qwen1.5b 54.0 65.9 1.4210 53.0858 8.5299 63.0370 1040
-LightRetriever-Qwen7b 53.8 66.5 0.8300 0.0420 8.5150 9.3870 6982
 3 Experiments
@@
 Our work finetunes on a large existing collection of 20 English and 3 Chinese datasets with 8.38M samples to ensure broad domain coverage and diversity.
-To ensure reproducibility, our work reuses the preprocessed fine-tuning data from tDRO [48] and Sentence Transformers Training Data [60].
-English datasets include Amazon Review [54], Eli5 [20], FEVER [67], FiQA [49], GooQA [31], HotpotQA [81], MSMARCO [53], NFCorpus [7], NPR [45], NQ [37], PAQ [41], Quora Duplicates [26], S2ORC [43], Scifact [70], Specter [13], StackExchange Duplicates [17], E5 synthetic [4], Trivia [29], WikiHow [34], and Yahoo Answers [83].
-Chinese datasets include cMedQA2 [82], DuReader [58], and T2Ranking [78].
+To ensure reproducibility, our work reuses the preprocessed fine-tuning data from tDRO (Ma et al., 2025) and Sentence Transformers Training Data.
+English datasets include Amazon Review (Ni et al., 2019), Eli5 (Fan et al., 2019), FEVER (Thorne et al., 2018), FiQA (Maia et al., 2018), GooAQ (Khashabi et al., 2021), HotpotQA (Yang et al., 2018b), MSMARCO (Nguyen et al., 2016), NFCorpus (Boteva et al., 2016), NPR (Lucas et al., 2023), NQ (Kwiatkowski et al., 2019), PAQ (Lewis et al., 2021), Quora Duplicates (Iyer et al., 2017), S2ORC (Lo et al., 2020), Scifact (Wadden et al., 2020), Specter (Cohan et al., 2020), StackExchange Duplicates (da Silva et al., 2018), E5 synthetic (BeastyZ, 2024), Trivia (Joshi et al., 2017), WikiHow (Koupaee and Wang, 2018), and Yahoo Answers (Zhang et al., 2015).
+Chinese datasets include cMedQA2 (Zhang et al., 2018), DuReader (Qiu et al., 2022), and T2Ranking (Xie et al., 2023).
 Benchmarks
-BeIR [66] and CMTEB Retrieval [77] benchmarks are used in the evaluation.
+BeIR (Thakur et al., 2021) and CMTEB Retrieval (Xiao et al., 2024) benchmarks are used in the evaluation.
 BeIR is a massive English retrieval benchmark collection with 15 evaluation sets.
@@
 CMTEB Retrieval is a massive Chinese retrieval benchmark collection with 8 sets, also covering a variety of domains.
-nDCG@10 is reported as the main metric by following [66, 77], accessing both recall and ranking abilities.
+nDCG@10 is reported as the main metric by following (Thakur et al., 2021; Xiao et al., 2024), accessing both recall and ranking abilities.
 Recalls (R@{20, 50, 100}) are also reported to assess retrieval ability at large windows.
 Training hyper-parameters
-Our experiments are conducted based on the implementation of tDRO [48].
+Our experiments are conducted based on the implementation of tDRO (Ma et al., 2025).
 All experiments are performed with a batch size of 128, 7 hard negatives, a max sequence length of 512, a contrastive temperature $\tau$ of 0.02, and 12k total steps.
-Following SPLADE [21], the coefficient of the FLOPs regulator is 0.001, and the regulator is quadratically increased to the maximum at the first 4k steps, which helps reduce the side effects of the regulator at initial steps [57].
-LoRA [25] is used to save training GPU memory, where LoRA r is 16, alpha is 32, and dropout is 0.1.
-Different LLMs with different sizes are tested as the backbone encoders to verify the generalization abilities, including Llama-3.2-1B, 3B, Llama-3.1-8B [68], and Qwen-2.5-1.5B, 3B, 7B [2].
-The hybrid similarity scores are linearly summed from normalized dense and sparse scores [73].
+Following SPLADE (Formal et al., 2021), the coefficient of the FLOPs regulator is 0.001, and the regulator is quadratically increased to the maximum at the first 4k steps, which helps reduce the side effects of the regulator at initial steps (Paria et al., 2020).
+LoRA (Hu et al., 2022) is used to save training GPU memory, where LoRA r is 16, alpha is 32, and dropout is 0.1.
+Different LLMs with different sizes are tested as the backbone encoders to verify the generalization abilities, including Llama-3.2-1B, 3B, Llama-3.1-8B (Touvron et al., 2023), and Qwen-2.5-1.5B, 3B, 7B (Bai et al., 2023).
+The hybrid similarity scores are linearly summed from normalized dense and sparse scores (Wang et al., 2021).
 All trainings are conducted on 8 NVIDIA H800 or A800 GPUs.
+Specifically, training a LightRetriver-Llama8b on 8 H800 consumes about 15 hours.
 Speed comparisons
-To evaluate the online speedup of LightRetriever, we compare the online time consumptions for retrieving 65536 Bing queries over 1M passages from MSMARCO[53].
+To evaluate the online speedup of LightRetriever, we compare the online time consumptions for retrieving 65536 Bing queries over 1M passages from MSMARCO (Nguyen et al., 2016).
 All speed tests are conducted on a single A800 GPU, with a batch size of 256, a dimension of 1k for Faiss exact search, and 64 threads for Anserini (Lucene) sparse search.
 Faiss and Anserini search in parallel, where Faiss searches more slowly in our cases.
+Performance comparisons
+[…]
+Notably, Static Embedding performs retrieval using a single shared embedding layer for both queries and documents, offering low retrieval performance but a highly efficient baseline.
+BM25, as a classical term-frequency-based method, remains a widely adopted sparse retriever.
+The combination of them thus provides a reference for efficient retrieval on both sides.
+In contrast, BGE-m3 is a Transformer-based retriever trained with multi-stage pretraining and fine-tuning, and represents state-of-the-art performance in hybrid retrieval.
+USTAD (Kim et al., 2024) reports a small 6-layer DistilBERT retriever, offering a SOTA distillation baseline for small language models (SLM).
+In Appendix A.5, applications of similar distillation techniques are also discussed in our works.
+Except for Static Embedding and BM25, all compared retrievers, especially LLM retrievers, rely on deep Transformer encoders on both query and document sides.
+LightRetriever is, to our knowledge, the first method to address the online efficiency bottleneck on the query side of LLM-based retrievers, striking a balance between retrieval quality and inference efficiency.
+3.2 Main results
+Encoding speeds
+Table 1 presents a detailed breakdown of retrieval time for processing 65,536 Bing queries over 1 million MSMARCO passages, highlighting the efficiency advantages of LightRetriever over full symmetric retrievers.
+Among all time consumptions, model encoding is the dominant contributor to latency in LLM-based retrievers.
+For example, Full-Llama8b and Full-Qwen7b require over 100 seconds solely for query encoding (109.5s and 100.7s, respectively), making them a bottleneck for large-scale, real-time retrieval.
+In contrast, LightRetriever reduces encoding time to below 50 ms (0.0412–0.0420s), reaching a maximal 2500× speedup of full-sized Llama-8b in the encoding phase.
+This drastic reduction is made possible by replacing full forward passes of large Transformer models with a simple embedding lookup.
+And the overall throughput achieves over 10x QPS speedup.
+Notably, this lightweight encoding does not significantly compromise retrieval performance.
+LightRetriever maintains competitive nDCG@10 scores on both English (BeIR) and Chinese (CMTEB-Retrieval) benchmarks.
+A potent baseline to improve LLM query encoder efficiency is layer reduction.
+We evaluated by training and serving only the first Transformer layer of the Llama3-8b model on the query side under the same settings, meanwhile retaining full-sized model on the document side.
+While this reduces model size during training, it underperforms the LightRetriever (–4.3 on BeIR, –8.6 on CMTEB-R) and introduces inference-time overhead of 2.34s.
+This proves the need for full-sized query modeling during the training phase.
+The above results prove the effectiveness of our extreme decoupled modeling paradigm, which drastically improves online query encoding efficiency while preserving retrieval quality, making it suitable for latency-critical applications.
+[…]
+Benchmark / nDCG@10 Time consumption / s & Throughput / QPS
+Model BeIR CMTEB-R Tokenize Encode Search Total QPS
+Full-Llama8b 56.8 67.6 1.3746 109.4853 8.5133 119.3730 549
+Full-Llama3b 55.6 66.1 1.3250 52.5890 8.5040 62.4180 1050
+Full-Llama1b 53.1 63.8 1.2564 19.9030 8.4897 29.6490 2210
+[…]
+LightRetriever-Llama8b 54.4 63.0 0.8209 0.0412* 8.5010 9.3630 6999*
 Table 2: Performance comparisons on BeIR and CMTEB Retrieval (CMTEB-R) benchmarks.
@@
 Static Embed + BM25 44.7 51.9 59.8 65.1 52.1 65.8 71.9 76.1
+USTAD 44.2 - - 65.3 - - - -
 […]
-[…]
-[…]
+SPLADE-v3 50.2 56.6 63.3 68.6 - - - -
 […]
@@
 Qwen2.5-7b 53.8-$_{\text{-2.8}}$ 60.3 67.5 72.5 66.5-$_{\text{-3.6}}$ 81.1 86.7 89.7
-Performance comparisons
-[…]
-Notably, Static Embedding performs retrieval using a single shared embedding layer for both queries and documents, offering low retrieval performance but highly efficient baseline.
-BM25, as a classical term-frequency-based method, remains a widely adopted sparse retriever.
-The combination of them thus provides a reference for efficient retrieval on both sides.
-In contrast, BGE-m3 is a Transformer-based retriever trained with multi-stage pretraining and fine-tuning, and represents state-of-the-art performance in hybrid retrieval.
-Except for Static Embedding and BM25, all compared retrievers rely on deep Transformer encoders on both query and document sides.
-LightRetriever is, to our knowledge, the first method to address the online efficiency bottleneck on the query side of LLM-based retrievers, striking a balance between retrieval quality and inference efficiency.
-3.2 Main results
-Encoding speeds
-Table 1 presents a detailed breakdown of retrieval time for processing 65,536 Bing queries over 1 million MS MARCO passages, highlighting the efficiency advantages of LightRetriever over full symmetric retrievers.
-Among all time consumptions, model encoding is the dominant contributor to latency in LLM-based retrievers.
-For example, Full-Llama8b and Full-Qwen7b require over 100 seconds solely for query encoding (109.5s and 100.7s, respectively), making them a bottleneck for large-scale, real-time retrieval.
-In contrast, LightRetriever reduces encoding time to below 50 ms (0.0412–0.0420s), achieving over a 2500× speedup in the encoding phase.
-This drastic reduction is made possible by replacing full forward passes of large Transformer models with a simple embedding lookup.
-And the overall throughput achieves over 10x QPS speedup.
-Notably, this lightweight encoding does not significantly compromise retrieval performance.
-LightRetriever maintains competitive nDCG@10 scores on both English (BeIR) and Chinese (CMTEB-Retrieval) benchmarks.
-These results prove the effectiveness of LightRetriever’s decoupled query modeling paradigm, which drastically improves online query encoding efficiency while preserving good retrieval quality, making it suitable for latency-critical applications.
-Performance comparisons
+Overall performance comparisons
 Table 2 compares retrieval effectiveness on English (BeIR) and Chinese (CMTEB-Retrieval) benchmarks.
 The main focus of this comparison is to evaluate the performance degradation introduced by LightRetriever’s lightweight query encoding.
-Overall, by leveraging deep Transformer encoders on both query and document sides, full symmetric retrievers with deep LLM encoders achieve the best scores (e.g., Llama3.1-8b on BeIR with 56.8 nDCG@10, Qwen2.5-7b on CMTEB-R with 70.1 nDCG@10), but at significant computational cost.
+Overall, by leveraging deep Transformer encoders on both query and document sides, full symmetric retrievers with deep LLM encoders achieve the best scores (e.g., Llama3.1-8b on BeIR with 56.8 nDCG@10, Qwen2.5-7b on CMTEB-R with 70.1 nDCG@10), but at a significant computational cost.
 LightRetriever achieves promising effectiveness while drastically reducing query-side computational burden.
@@
 For example, LightRetriever-Qwen2.5-7b achieves 53.8 nDCG@10 on BeIR and 66.5 nDCG@10 on CMTEB-R, only 2.8 and 3.6 lower than its full counterpart.
-Detailed task-level results are shown in Appendix Tables 911.
-Notably, LightRetriever retains strong performance on large-scale open-domain QA, fact-checking, and web search benchmarks such as FEVER [67], SciFact [70], TRECCOVID [16], and MSMARCO [53].
-On smaller, domain-specific datasets (e.g., FiQA [49] and CmedQA [82]), drops can be moderately larger (around 5 points) but remain tolerable, suggesting good generalization with room for future task-specific adaptation.
 […]
+Category-wise performance comparison
+To provide a more fine-grained analysis of our query-lightweight proposal across different retrieval scenarios, we further categorize BeIR datasets into 8 task groups by following the original paper (Thakur et al., 2021) and summarize the results in Table 3.
+Overall, LightRetriever shows strong robustness on general relevance-based retrieval tasks, including Bio-Medical IR, Argument Retrieval, Fact Checking, Duplicate Question Retrieval, Passage Retrieval, and General QA, where it consistently preserves over 93% of the full symmetric retriever’s effectiveness, and even surpasses it in some cases (e.g., TREC-COVID, ArguAna, and Touche2020).
+This indicates that lightweight query encoding is sufficient to capture the semantic signals in typical relevance-based ad-hoc and open-domain retrieval.
+In contrast, moderately larger performance degradations are observed on more challenging tasks, such as Domain-specific QA (HotpotQA, FiQA), Entity Retrieval (DBPedia), and Citation Prediction (SCIDOCS), where the retention drops to around 87-89%.
+Additional benchmarks on CoIR (Li et al., 2025) and BRIGHT (Su et al., 2025) in §A.10 also show inferior performance of query-lightweight design on code and reasoning-intensive retrieval.
+These tasks typically require higher reasoning and query-understanding ability, which is not suitable for an extreme lightweight query encoder.
+Nevertheless, LightRetriever still maintains competitive absolute performance, suggesting promising generalization with room for further task-specific adaptation.
+Table 3: Category-wise performance comparison on BEIR.
+We report the average nDCG@10 of the full symmetric retriever (Llama-3.1-8B), LightRetriever (Llama-3.1-8B), and the performance retention of LightRetriever (LightRetriever/Full-Symmetric).
+Task Datasets Full Light Retention
+Bio-Medical IR TREC-COVID 62.0 71.5 115.3%
+Argument Retrieval ArguAna, Touche2020 39.7 42.5 107.2%
+Fact Checking FEVER, ClimateFEVER, SciFact 70.8 67.6 95.5%
+Duplicate Question Retrieval CQADupstack, Quora 69.6 65.3 93.8%
+Passage Retrieval / General QA MSMARCO, NFCorpus, NQ 50.1 47.0 93.8%
+Domain-specific QA HotpotQA, FiQA2018 68.4 61.1 89.3%
+Entity Retrieval DBPedia 48.2 42.0 87.1%
+Citation Prediction SCIDOCS 23.8 20.7 87.0%
 4 Ablations
-Our core design adopts an asymmetric architecture, with a lightweight query MLP Embedding and a full-sized document model for inference.
+Our core design adopts an asymmetric architecture, with a lightweight query Embedding Bag (Emb) and a full-sized document model for inference.
 Specifically, we utilize a full-sized model on both sides at training time, then cache all query token embeddings in one Embedding Lookup for fast query-side inference.
-In ablation studies below (A1-A3), we are interested in other potential symmetric/asymmetric architectures, which are described in the following Table 4.
-Table 3: Symmetry Ablation settings (A1-A3). 1L-Trans denotes a single-layer Transformer.
+In ablation studies below (A1-A2), we are interested in other potential symmetric/asymmetric architectures, which are described in the following Table 5.
+Table 4: Symmetry Ablation settings (A1-A2). Emb denotes Embedding Bag. 1L-Trans denotes a single-layer Transformer.
 Train Inference
 Query Doc Query Doc
-Ours Full Full MLP Full
-A1 Full Full MLP MLP
-A2 MLP Full MLP Full
-A3 1L-Trans Full 1L-Trans Full
-Table 4: Ablations (A1-3) results (nDCG@10) of LightRetriever.
+Ours Full Full Emb Full
+A1 Full Full Emb Emb
+A2 Emb Full Emb Full
+Table 5: Ablations (A1-2) results (nDCG@10) of LightRetriever.
 Ablations BeIR CMTEB-R
@@
 A1: Both-side Light 34.9-13.8 40.1-18.6
-A2: Encq use a MLP 37.5-11.2 41.3-17.4
-Llama-8b 54.4 63.0
-A3: Encq use 1L-Trans 50.1-4.3 54.4-8.6
+A2: Encq use a Emb 37.5-11.2 41.3-17.4
 A1. Is symmetric lightweight inference effective?
-To test the necessity of the above asymmetry, we evaluate a symmetric lightweight setup where both sides use simplified dense encoders.
+To test the necessity of above asymmetry, we evaluate a symmetric lightweight setup where both sides use simplified dense encoders.
 As shown in Table 5, this leads to severe performance drops on both BeIR (–13.8) and CMTEB-R (–18.6), confirming that full document representations are crucial for maintaining retrieval effectiveness.
 This symmetric setting is not functional for sparse retrieval due to the absence of learnable parameters.
-The best symmetric lightweight retriever is Static Embedding (one MLP) + BM25 (term-based) in Table 2, with nDCG@10 of 44.7 on BeIR and 52.1 on CMTEB-R.
+The best symmetric lightweight retriever is Static Embedding (one Embedding Bag) + BM25 (term-based) in Table 2, with nDCG@10 of 44.7 on BeIR and 52.1 on CMTEB-R.
 However, due to the lack of deep modeling on documents, such a combination is still outperformed.
-We also tested a reversed asymmetry with deep encoding for queries and lightweight documents. However, we observed training instability and degraded performance, likely caused by embedding collapse.
-A2. Can we use only a query MLP for training?
-To confirm that a deep query encoder is still needed during training, we replaced the Transformer-based query encoder with a simple MLP that reuses only the LLM’s embedding layer.
-This setup significantly degrades performance (–11.2 on BeIR, –17.4 on CMTEB-R), indicating that the full-sized modeling for query side training remains essential during training to learn effective token-level representations.
-A3. Is adding one Transformers layer better on the query side?
-We also evaluated using only the first Transformer layer on the query side.
-While this reduces model size during training, it still underperforms the LightRetriever (–4.3 on BeIR, –8.6 on CMTEB-R) and introduces inference-time overhead, making it incompatible with the design goals of LightRetriever.
-This again proves the need for full-sized query modeling during the training phase.
-Figure 5: Ablations of retrieval performances (nDCG@10) of LightRetriever on BeIR and CMTEB-R benchmarks.
-A4. Breakdown Ablation of Retrieval Performance
+We also tested a reversed asymmetry with deep query encoding and lightweight documents.
+However, we observed training instability and degraded performance, likely caused by embedding collapse.
+A2. Can we use only a query Embedding Bag for training?
+To confirm that a deep query encoder is still needed during training, we replaced the Transformer-based query encoder with a simple Embedding Bag that reuses only the LLM’s embedding layer.
+This setup significantly degrades performance (–11.2 on BeIR, –17.4 on CMTEB-R), indicating that although deep token interaction could be limited with a durable cost during query inferencing, the full-sized modeling for query side training remains essential to learn effective representations.
+Figure 5: Ablations of retrieval performances (nDCG@10) of LightRetriever on BeIR and CMTEB-R benchmarks. For detailed results, please refer to Table 14, 15 and 16 in Appendix.
+A3. Breakdown Ablation of Retrieval Performance
 Figure 5 reports LightRetriever’s performance under dense, sparse, and hybrid configurations.
@@
 While documents could be pre-encoded and indexed offline, queries arrive in real-time and need online encoding.
-The deployment of large LLMs results in inefficient online services.
+The deployment of LLMs results in inefficient online services.
 In this paper, we propose LightRetriever, a novel LLM-based hybrid retrieval architecture capable of extremely lightweight query encoding.
-Our approach achieves over a 1000x speedup for query inference on an H800 GPU, even a 20x speedup without GPU acceleration.
+Our approach achieves over a 1000x speedup for query inference and 10x increase in end-to-end retrieval throughput on an A800 GPU.
 Experiments show great robustness and generalization abilities across different foundation LLMs and retrieval tasks.
-Appendix A Technical Appendices and Supplementary Material
+Acknowledgments
+This work is supported by the National Natural Science Foundation of China (No. U24A20335).
+Appendix A Appendix
 A.1 Limitations
-Our work focuses on online efficiency optimizations of recent LLM-based architecture, whose giant parameters pose a greater need for such optimizations.
+Our work focuses on online efficiency optimizations of recent LLM-based architectures, whose giant parameters pose a greater need for such optimizations.
 Other readers may be interested in the application of our methods on non-LLM architectures, e.g., BERT.
 While such a combination is possible, due to limited time constraints and lesser needs for non-LLM, we leave this to future work.
+Moreover, as stated in Section §3.2, our query-lightweight design is applicable to normal relevance-based retrieval tasks.
+Tasks that require complex query understanding, such as reasoning-intensive retrieval, are not suitable for our work.
 A.2 Related works
 A.2.1 Text retrieval
-Text retrieval builds on top of the vector space model [63], which encodes query and document representation vectors, and performs relevance search based on vector similarities.
+Text retrieval builds on top of the vector space model (Salton et al., 1975), which encodes query and document representation vectors, and performs relevance search based on vector similarities.
 Dense retrieval
-Dense retrieval [30] trains PLM-based encoders to condense queries and passages into low-dimensional dense vectors, then performs relevance search based on the Maximum Inner Product Search (MIPS) [52] algorithm.
-In recent years, dense retrievers, such as BGE [11] and E5 [71], have gained popularity for strong retrieval abilities across different tasks [51, 66], languages [84], and context granulaties [75, 53], owing to diverse training data [60], retrieval-customized pre-training [22, 76, 47], improved negative-mining [79], and so on.
-Recent SOTA retrievers [72, 5, 39] start to utilize LLMs [27, 68] as backbone encoders.
+Dense retrieval (Karpukhin et al., 2020) trains PLM-based encoders to condense queries and passages into low-dimensional dense vectors, then performs relevance search based on the Maximum Inner Product Search (MIPS) (Mussmann and Ermon, 2016) algorithm.
+In recent years, dense retrievers, such as BGE (Chen et al., 2024) and E5 (Wang et al., 2024b), have gained popularity for strong retrieval abilities across different tasks (Muennighoff et al., 2023; Thakur et al., 2021), languages (Zhang et al., 2023), and context granulaties (Williams et al., 2018; Nguyen et al., 2016), owing to diverse training data, retrieval-customized pre-training (Gao and Callan, 2021; Xiao et al., 2022; Ma et al., 2024), improved negative-mining (Xiong et al., 2021), and so on.
+Recent SOTA retrievers (Wang et al., 2024a; BehnamGhader et al., 2024; Lee et al., 2024) start to utilize LLMs (Jiang et al., 2023; Touvron et al., 2023) as backbone encoders.
 With larger parameters and pre-training data, LLM-based retrievers enable significantly powerful retrieval capacities and domain adaptation abilities.
@@
 Sparse retrieval
-Traditional statistical sparse retrieval algorithms (e.g., TF-IDF [64], and BM25 [61]) do not involve learnable language models.
+Traditional statistical sparse retrieval algorithms (e.g., TF-IDF (Salton and Buckley, 1988), and BM25 (Robertson et al., 1994)) do not involve learnable language models.
 Instead, they directly analyze lexical frequencies (e.g., TF) and importance (e.g., IDF) in a large vocabulary space.
-These lexicals are sparsely distributed and require building inverted indexes [46] for feasible similarity metric calculation, such as Apache Lucene [6] and Tantivy [59].
+These lexicals are sparsely distributed and require building inverted indexes (Luhn, 1957) for feasible similarity metric calculation, such as Apache Lucene (Bialecki et al., 2012) and Tantivy.
 In recent years, sparse retrieval has started to work with pre-trained LMs (PLMs).
-BGE [11], SPLADE [21], SparTerm [3], and SPARTA [85] use customized XLM-RoBERTa [14] or BERT [19] as backbone encoders to learn term frequencies in an end-to-end manner and directly perform impact search based on the dot product of sparse vectors.
+BGE (Chen et al., 2024), SPLADE (Formal et al., 2021), SparTerm (Bai et al., 2020), and SPARTA (Zhao et al., 2021) use customized XLM-RoBERTa (Conneau et al., 2020) or BERT (Devlin et al., 2019) as backbone encoders to learn term frequencies in an end-to-end manner and directly perform impact search based on the dot product of sparse vectors.
 Because of sparsity, these works also search with inverted indexes.
-Recent works [86] also try to incorporate LLMs with sparse retrieval.
+Recent works (Zhuang et al., 2024) also try to incorporate LLMs with sparse retrieval.
 However, the online efficiency issue still exists.
+Learning dynamic analysis (Mackenzie et al., 2023) on learned sparse retreival shows that semantic information can be fully carried on the document side regardless of token semantics.
+Motivated by this finding, in this paper we hypothesize that if the semantic capacity is concentrated on the document side, then the query side can be made extremely lightweight with minimal performance degradation.
 Hybrid retreval
-Hybrid retrieval is a technique that combines retrieval scores [73] or ranks [15] from multiple retrieval systems, such as dense [30] / sparse [61, 21] retrieval, and single [30] / multiple [32, 11] vector retrieval.
+Hybrid retrieval is a technique that combines retrieval scores (Wang et al., 2021) or ranks (Cormack et al., 2009) from multiple retrieval systems, such as dense (Karpukhin et al., 2020) / sparse (Robertson et al., 1994; Formal et al., 2021) retrieval, and single (Karpukhin et al., 2020) / multiple (Khattab and Zaharia, 2020; Chen et al., 2024) vector retrieval.
 It enables better retrieval abilities by interpolating results from multiple sources.
-Linear interpolation [73] of dense and sparse scores is used in our work.
+Linear interpolation (Wang et al., 2021) of dense and sparse scores is used in our work.
 A.2.2 Improving inference efficiency
@@
 Previous works have tried to distill small query encoders from large encoders for better query inference efficiency.
-KALE [9] prunes the BERT-based query encoder and distills it during post-training.
-However, when the query encoder is pruned to one Transformers layer, its recall@20 on NQ [37] decreases by 25-30%.
-[74] proposes asymmetric dual-encoders by pruning BERT layers or using smaller BERT models.
+KALE (Campos et al., 2023) prunes the BERT-based query encoder and distills it during post-training.
+However, when the query encoder is pruned to one Transformers layer, its recall@20 on NQ (Kwiatkowski et al., 2019) decreases by 25-30%.
+(Wang and Lyu, 2023) proposes asymmetric dual-encoders by pruning BERT layers or using smaller BERT models.
 Its experiments show that a 2-layer BERT query encoder retains 92.5% of the full-sized performances via distillation and proper student initialization.
-USTAD [33] also proposes similar asymmetry and distillation strategies, which choose the 12-layer BERT as teachers and the 6-layer DistilBERT or 2-layer BERT-mini as students.
+USTAD (Kim et al., 2024) also proposes similar asymmetry and distillation strategies, which choose the 12-layer BERT as teachers and the 6-layer DistilBERT or 2-layer BERT-mini as students.
 Its experiments show that the distilled BERT-mini query encoder achieves 95-97% of the teacher’s performance because of better distillation.
-However, these studies have primarily focused on small language models (SLMs), such as BERT [19], and overlooked feasible approaches for improving online inference efficiency of LLM-based retrievers, where such optimization is even more critical.
+However, these studies have primarily focused on small language models (SLMs), such as BERT (Devlin et al., 2019), and overlooked feasible approaches for improving online inference efficiency of LLM-based retrievers, where such optimization is even more critical.
 Furthermore, these existing methods typically retain Transformer layers in the query encoders, which limits their efficiency gains.
-In addition, most of these works rely on SLMs and are trained on relatively narrow datasets, such as MS-MARCO [53], thus unable to compete with LLM-based retrievers.
+In addition, most of these works rely on SLMs and are trained on relatively narrow datasets, such as MS-MARCO (Nguyen et al., 2016), thus unable to compete with LLM-based retrievers.
 In contrast, with the rapid advancement of LLM-based text retrievers, driven by larger and more diverse foundation models and datasets, retrieval performance has significantly improved.
@@
 Previous works have also been dedicated to reducing index sizes for smaller disk space usage.
-Matryoshka representation learning (MRL) [36] supports adjusting the dense vector dimensions by directly taking the top-k dimension of the original vector as shrink vectors.
+Matryoshka representation learning (MRL) (Kusupati et al., 2022) supports adjusting the dense vector dimensions by directly taking the top-k dimension of the original vector as shrink vectors.
 It aligns and trains the corresponding multi-dimensional vectors in the end-to-end contrastive learning.
 For example, if $k\in\{64,128,256\}$, query-passage dense vectors with $\{64,128,256\}$-dims could be obtained together after one encoder forward operation.
-For LM-based sparse retrieval, the similarity scores are directly affected by the impacts of corresponding lexicals [21].
+For LM-based sparse retrieval, the similarity scores are directly affected by the impacts of corresponding lexicals (Formal et al., 2021).
 Thus, it naturally supports reducing the index sizes by retaining Top-k terms/lexicals.
 To explore the potential of controlling index sizes, our work also explores shrinking index sizes with dense MRL and sparse Top-k in the Section A.4, which also shows strong generalization capacities.
-Figure 6: Performances (Llama-3.1-8b) with different Top-k dimensions or sparsities.
-A.3 Additional results about controls of vector dimension and sparsity
+A.3 Case Study
+Here, we report some hand-crafted ranking cases to show LightRetriever’s ability to understand complex queries rather than simple lexical matches, even if the query encoder is simplified and limited for inference efficiency.
+Table 6: Case Study 1: Distinguish compositional semantics.
+Texts colored in red are important for semantic matching, while the blue ones represent unrelated words or distractions.
+$\checkmark$ means ranking correct, while $\times$ means rank fail.
+Query Which camera captures fast-moving subjects with minimal motion blur?
+Doc A (relevant) The X200 features a 1/8000s shutter and phase-detection AF, enabling crisp photos of sports and fast action with minimal motion blur.
+Doc B The Y7 is praised for its wide-angle lens and superior low-light performance, often used for landscapes and night photography.
+Results Full-sized Llama8b $\checkmark$ Doc A >> Doc B LightRetriever-Llama8bDense $\checkmark$ Doc A >> Doc B LightRetriever-Llama8bSparse $\checkmark$ Doc A >> Doc B LightRetriever-Llama8bHybrid $\checkmark$ Doc A >> Doc B
+Case Study 1: Distinguish compositional semantics.
+This case tests the ability to search multiple attributes together, capturing fast-moving subjects with minimal motion blur.
+While both candidate documents discuss camera capabilities, only Doc A contains the correct composition of features.
+Doc B, although relevant to photography, emphasizes low-light and landscape use cases, which do not align with the compositional semantics.
+LightRetriever, despite using a simplified query encoder, successfully prioritizes Doc A across dense, sparse, and hybrid variants.
+Table 7: Case Study 2: Negation.
+Query Articles arguing against remote work for improving team collaboration.
+Doc A (relevant) A recent study argues that remote work hinders spontaneous collaboration, noting dropoffs in serendipitous conversations and mentoring opportunities among junior staff.
+Doc B Many experts advocate remote work for flexibility and increased individual productivity, highlighting asynchronous tools that maintain collaboration.
+Results Full-sized Llama8b $\checkmark$ Doc A >> Doc B LightRetriever-Llama8bDense $\checkmark$ Doc A >> Doc B LightRetriever-Llama8bSparse $\checkmark$ Doc A >> Doc B LightRetriever-Llama8bHybrid $\checkmark$ Doc A >> Doc B
+Case Study 2: Negation.
+This case tests whether models can correctly interpret affirmation/negation.
+The query explicitly seeks articles against remote work for improving collaboration.
+Although both candidate documents discuss remote work and collaboration, only Doc A aligns with the negative phase by describing how remote work hinders interactions.
+Doc B, on the contrary, advocates for remote work, which is semantically opposite to the query intent.
+Correctly distinguishing the affirmation/negation requires more than lexical overlap with remote work and collaboration.
+As shown in the results, both the full-sized Llama8b and all LightRetriever variants successfully rank Doc A higher, demonstrating that LightRetriever can preserve such negation discrimination ability.
+Table 8: Case Study 3: Complex understanding.
+Query How to revert the last Git commit without losing the changes?
+Doc A (relevant) Use git reset –soft HEAD 1 to undo the last commit while keeping changes in the working tree and index; this retains your edited files for further edits.
+Doc B Use git revert HEAD to create a new commit that undoes the changes introduced by the previous commit, preserving history and not altering the working tree.
+Results Full-sized Llama8b $\checkmark$ Doc A >> Doc B LightRetriever-Llama8bDense $\times$ Doc A << Doc B LightRetriever-Llama8bSparse $\checkmark$ Doc A >> Doc B LightRetriever-Llama8bHybrid $\times$ Doc A << Doc B
+Case Study 3: Complex understanding.
+This case tests the ability to understand complex technical instructions.
+The query asks how to revert the last Git commit without losing the changes, a subtle requirement that distinguishes between keeping edits in the working tree (Doc A) versus creating a new commit that reverts prior changes (Doc B).
+Despite both documents mentioning ways to undo commits, only Doc A satisfies the instructions.
+The full-sized Llama8b is able to complete such complex technical instructions.
+However, LightRetriever-Dense and Hybrid fail, likely because they focus on the strong lexical signal of revert but neglect the constraint without losing the changes.
+Interestingly, LightRetriever-Sparse ranks Doc A correctly, suggesting that token-level sparse matching may emphasize critical keywords such as keeping changes and working tree.
+This highlights a limitation of lightweight dense encoding: complex compositional understanding may fail, and success may depend on complementary sparse signals.
+A.4 Additional results about controls of vector dimension, sparsity, and embedding serving size
 Retrieval systems usually involve indexing millions of documents into search databases, which consumes large amounts of disk space, especially for dense vectors.
-For example, indexing 8.8M MS-MARCO passages [53] with Llama-8b model consumes 134.92 GB for Faiss dense retrieval and 11.24 GB for Lucene sparse retrieval.
-Thus, it’s essential to explore the possibility of controlling index sizes by shrinking vectors.
-1) For dense retrieval, we incorporate Matryoshka Representation Learning (MRL) [36] with LightRetriever to support flexible controls of dense vector dimensions.
+For example, indexing 8.8M MS-MARCO passages (Nguyen et al., 2016) with the Llama-8b model consumes 134.92 GB for Faiss dense retrieval and 11.24 GB for Lucene sparse retrieval.
+Besides index sizes, LightRetriever’s dense part also serves as an Embedding matrix in RAM with shape [V, H] in Float16 precision.
+For the Llama-8b model, if the serving size is not properly controlled, such online embedding will consume 1.05 GB of RAM.
+Thus, it’s essential to explore the possibility of controlling index sizes and serving sizes by shrinking the vector dimensions.
+1) For dense retrieval, we incorporate Matryoshka Representation Learning (MRL) (Kusupati et al., 2022) with LightRetriever to support flexible controls of dense embedding dimensions.
 Given a dense vector $v^{den}$ with dimension $H$, MRL derives dimension-shrunk vectors by simply cutting out the Top-k dimensions $v^{den}[:k]$, then applies the same training loss as the original vectors.
@@
 Thus, LightRetriever supports controlling sparsity by directly taking the Top-k terms, where the larger impact terms matter more.
+Figure 6: Performances (Llama-3.1-8b) with different Top-k dimensions or sparsities.
+Table 9: Performances, dense serving size, and index size of LightRetriever-Llama8b across different top-k dimensions (Dim). Note that dim=4096 is the untruncated setting.
+Dimension Performance / nDCG@10 Serving Size (MB) Index Size (GB)
+Top-k BEIR CMTEB-R Embedding Dense Sparse
+128 49.3 55.2 32.83 4.22 3.03
+256 51.6 57.5 65.67 8.43 5.42
+512 53.0 58.9 131.33 16.87 8.89
+1024 53.1 59.8 262.67 33.73 10.85
+2048 53.3 60.4 525.34 67.46 11.22
+4096 54.4 63.0 1050.67 134.92 11.24
+Index size controls via tuning vector dimension and sparsity
 We conduct experiments of controlling dense vector dimension and sparse vector sparsity with Top-k of $\{128,256,512,1024,2048,4096\}$ on Llama-3.1-8b.
-As presented in Figure 6, the performances of LightRetriever decrease marginally when shrinking Top-k to 4x ($4090\rightarrow 1024$).
-The lowest Top-k setting ($128$) has 7.9% and 8.8% performance gaps on BeIR and CMTEB-R benchmarks compared to the highest Top-k ($4096$), but its disk usages with MS-MARCO are only 4.22 GB for dense trieval and 3.03 GB for sparse retrieval.
+As presented in Figure 6 and Table 9, LightRetriever has affordable performance cost when truncating the Top-k.
+The lowest Top-k setting ($128$) has 5.1 and 7.8 nDCG@10 performance gaps on BeIR and CMTEB-R benchmarks compared to the highest Top-k ($4096$), but its disk usages with MS-MARCO are only 4.22 GB for dense trieval and 3.03 GB for sparse retrieval.
 This shows that our LightRetriever is capable of flexible index controls via simple vector shrinking techniques, with limited effect on retrieval performances.
-A.4 Additional ablation about the effect of auxiliary KL loss
+Dense embedding serving size controls via dimension truncation
+[…]
+A.5 Additional ablation about the effect of auxiliary KL loss
 Table 10: Ablations (nDCG@10, w/Llama-3.1-8b) on BeIR and CMTEB-R about auxiliary KL loss.
@@
 Hybrid 54.4 63.0 53.9-0.5 62.1-0.9
-Previous works for BERT-based asymmetric retrievers [74, 33] emphasize the importance of alignment between asymmetric students and full symmetric teachers.
+Previous works for BERT-based asymmetric retrievers (Wang and Lyu, 2023; Kim et al., 2024) emphasize the importance of alignment between asymmetric students and full symmetric teachers.
 To explore the effect of such alignment, besides contrastive loss as the main loss function, we also explore to adopt the KL loss as an auxiliary to align the asymmetric similarity scores $S^{student}=v_{q}^{den}\cdot v_{d}^{den}$ to full-sized symmetric similarity scores $S^{teacher}=Enc_{q}(q)\cdot Enc_{d}(d)$,
@@
 Results in Table 10 show that removing such loss incurs limited drops of -0.5 and -0.9 on BeIR and CMTEB-R. Thus, such auxiliary KL loss is not mandatory needed in our experiments.
-A.5 Training datasets
-Dataset infos
-Following the settings of previous LLM-based retrievers [72, 5, 48], our work uses 23 training sets pre-built by tDRO [48] or Sentence Transformers Training Data [60], including 20 English and 3 Chinese datasets.
-Detailed dataset information is listed in Table 6 as follows.
 Table 11: Training Dataset informations.
 Dataset Language Category Deduped Size Epoch Ratio
-Amazon Review (2018) [54] English Amazon $999$ 0.01 0.58%
-cMedQA2 [82] Chinese Chinese Medical 99936 1 5.84%
-DuReader [58] Chinese Chinese Web Collections 86366 1 5.05%
-Eli5 [20] English Reddit 325390 0.05 0.95%
-Fever [67] English Wikipedia QA 109808 0.8 5.13%
-FiQA [49] English Financial 5498 1 0.32%
-GooQA Pairs [31] English Web Collections 3012347 0.025 4.40%
-HotpotQA [81] English Wikipedia QA 85000 0.5 2.48%
-MSMARCO [53] English Web Collections 502854 1 29.39%
-NFCorpus [7] English Medical 2585 1 0.15%
-NPR [45] English News 594376 0.05 1.74%
-NQ [37] English Wikipedia QA 58800 1 3.44%
-PQA Pairs [41] English Wikipedia QA 99999 0.5 2.92%
-Quora Duplicates Triples [26] English Forum Duplicates 97011 0.5 2.83%
-S2ORC Title-Abstract [43] English Semantic Scholar 99998 0.5 2.92%
-SciFact [70] English S2ORC 806 1 0.05%
-SPECTER [13] English Semantic Scholar 136642 0.25 2.00%
-StackExchange Duplicates (Title-Body) [17] English Forum Duplicates 250516 0.25 3.66%
-E5 Synthetic [4] English GPT Synthetic Data 224791 1 13.14%
-T2Ranking [78] Chinese Chinese Web Collections 200362 0.5 5.86%
-Trivia [29] English Wikipedia QA 60370 0.5 1.76%
-Wikihow [34] English WikiHow 128543 0.25 1.88%
-Yahoo Answers (Title-Answer) [83] English Yahoo 1198018 0.05 3.50%
-Many downstream retrieval benchmarks [66, 77] are evaluated out-of-domain, where no training samples from the downstream sources are available.
-To avoid the overfitting of training sets, as is shown in the above Table 6, we slightly tune the epoch used in the dataset sampling, by following previous works in [55, 72, 48].
-Training set decontamination
-Training set decontamination is essential for developing fair, unbiased, and robust retrieval systems.
-Existing training datasets are derived from multiple sources, whose training data have possible duplicates with the test sets.
-Following tDRO [48], we perform strict training set decontamination with SimHash [50] to ensure no training queries appear in the BeIR [66] and CMTEB Retrieval [77].
-The original sizes, deduplicated sizes, and duplicate numbers as listed in the Table 7.
+Amazon Review (2018) (Ni et al., 2019) English Amazon $999$ 0.01 0.58%
+cMedQA2 (Zhang et al., 2018) Chinese Chinese Medical 99936 1 5.84%
+DuReader (Qiu et al., 2022) Chinese Chinese Web Collections 86366 1 5.05%
+Eli5 (Fan et al., 2019) English Reddit 325390 0.05 0.95%
+Fever (Thorne et al., 2018) English Wikipedia QA 109808 0.8 5.13%
+FiQA (Maia et al., 2018) English Financial 5498 1 0.32%
+GooAQ Pairs (Khashabi et al., 2021) English Web Collections 3012347 0.025 4.40%
+HotpotQA (Yang et al., 2018b) English Wikipedia QA 85000 0.5 2.48%
+MSMARCO (Nguyen et al., 2016) English Web Collections 502854 1 29.39%
+NFCorpus (Boteva et al., 2016) English Medical 2585 1 0.15%
+NPR (Lucas et al., 2023) English News 594376 0.05 1.74%
+NQ (Kwiatkowski et al., 2019) English Wikipedia QA 58800 1 3.44%
+PQA Pairs (Lewis et al., 2021) English Wikipedia QA 99999 0.5 2.92%
+Quora Duplicates Triples (Iyer et al., 2017) English Forum Duplicates 97011 0.5 2.83%
+S2ORC Title-Abstract (Lo et al., 2020) English Semantic Scholar 99998 0.5 2.92%
+SciFact (Wadden et al., 2020) English S2ORC 806 1 0.05%
+SPECTER (Cohan et al., 2020) English Semantic Scholar 136642 0.25 2.00%
+StackExchange Dups (Title-Body) (da Silva et al., 2018) English Forum Duplicates 250516 0.25 3.66%
+E5 Synthetic (BeastyZ, 2024) English GPT Synthetic Data 224791 1 13.14%
+T2Ranking (Xie et al., 2023) Chinese Chinese Web Collections 200362 0.5 5.86%
+Trivia (Joshi et al., 2017) English Wikipedia QA 60370 0.5 1.76%
+Wikihow (Koupaee and Wang, 2018) English WikiHow 128543 0.25 1.88%
+Yahoo Answers (Title-Answer) (Zhang et al., 2015) English Yahoo 1198018 0.05 3.50%
+A.6 Training datasets
+Dataset infos
+Following the settings of previous LLM-based retrievers (Wang et al., 2024a; BehnamGhader et al., 2024; Ma et al., 2025), our work uses 23 training sets pre-built by tDRO (Ma et al., 2025) or Sentence Transformers Training Data, including 20 English and 3 Chinese datasets.
+Detailed dataset information is listed in Table 11 as follows.
+Many downstream retrieval benchmarks (Thakur et al., 2021; Xiao et al., 2024) are evaluated out-of-domain, where no training samples from the downstream sources are available.
+To avoid the overfitting of training sets, as is shown in the above Table 11, we slightly tune the epoch used in the dataset sampling, by following previous works in (Ni et al., 2022; Wang et al., 2024a; Ma et al., 2025).
 Table 12: Training set deduplications. We report each dataset with the original size, deduplicated size, and number of duplicates detected in both the train and test sets.
@@
 FiQA 5498 5498 0
-GooQA Pairs 3012496 3012347 149
+GooAQ Pairs 3012496 3012347 149
 HotpotQA 85000 85000 0
@@
 Yahoo Answers (Title-Answer) 1198260 1198018 242
-Additionally, we also listed some samples of the duplicated training queries, as well as the corresponding test sets in Table 8.
 Table 13: Samples of duplicate queries found in both the train and test sets.
@@
 Eli5 train-3088 How does unemployment insurance work? FiQA 2648 How does unemployment insurance work?
-A.6 Task instructions
-As previously mentioned in Section 2.1, modern LLM-based retrievers [72, 5] introduce task-specific instructions to format input queries, which enables higher retrieval capacities.
-These LLM-based instruction-tuned retrieval is similar to instruction-based supervised fine-tuning (SFT) in LLM generation [56].
-Following Mistral-E5 [72], we reuse most of its task instructions, and format the instruction and query in the following format:
+Training set decontamination
+Training set decontamination is essential for developing fair, unbiased, and robust retrieval systems.
+Existing training datasets are derived from multiple sources, whose training data have possible duplicates with the test sets.
+Following tDRO (Ma et al., 2025), we perform strict training set decontamination with SimHash (Manku et al., 2007) to ensure no training queries appear in the BeIR (Thakur et al., 2021) and CMTEB Retrieval (Xiao et al., 2024).
+The original sizes, deduplicated sizes, and duplicate numbers as listed in the Table 12.
+Additionally, we also listed some samples of the duplicated training queries, as well as the corresponding test sets in Table 13.
+A.7 Task instructions
+As previously mentioned in Section 2.1, modern LLM-based retrievers (Wang et al., 2024a; BehnamGhader et al., 2024) introduce task-specific instructions to format input queries, which enables higher retrieval capacities.
+These LLM-based instruction-tuned retrieval is similar to instruction-based supervised fine-tuning (SFT) in LLM generation (Ouyang et al., 2022).
+Following Mistral-E5 (Wang et al., 2024a), we reuse most of its task instructions, and format the instruction and query in the following format:
 […]
-The detailed instructions are listed below, most of which are directly copied from the previous works E5-Mistral [72] and tDRO [48] for reproducibility.
+The detailed instructions are listed below, most of which are directly copied from the previous works E5-Mistral (Wang et al., 2024a) and tDRO (Ma et al., 2025) for reproducibility.
 Note that one training set may correspond to multiple instructions to avoid overfitting.
@@
   6. 6.
-    GooQA Pairs: Given a web search query, retrieve the corresponding answers from Google
+    GooAQ Pairs: Given a web search query, retrieve the corresponding answers from Google
   7. 7.
@@
     VideoRetrieval: Given a video search query, retrieve the titles of relevant videos
-A.7 Efficiency optimizations techniques
+A.8 Efficiency optimizations techniques
 Customized causal mask
@@
 The customized causal mask ensures that the token blocks can attend to the common prompt area, but not attend to each other.
-Let the sequence length be $L$, prompt length be $P$, micro block (each token + <eos>) width be $w$, and let $q\in[0,L-1]$ denote the query index, $k\in[0,L-1]$ the key index.
+Let the sequence length be $L$, prompt length be $P$, micro block (each token + eos) width be $w$, and let $q\in[0,L-1]$ denote the query index, $k\in[0,L-1]$ the key index.
 The attention mask (0 means attend, -$\infty$ means masked) is defined as,
@@
 , where the vocabulary dimension $V$ is much larger than the hidden dimension $H$.
-For example, for Llama3.1-8b [68] model, $V=128256$ and $H=4096$, where $\lfloor V/H\rfloor=31$.
+For example, for Llama3.1-8b (Touvron et al., 2023) model, $V=128256$ and $H=4096$, where $\lfloor V/H\rfloor=31$.
 This means that, given batch size B, sequence length T, hidden size H, vocabulary size V, the intermediate tensors with shape $[B,T,V]$ in the projection, ReLU, and Log are 31x larger than a dense tensor $[B,T,H]$, which consumes huge amounts of GPU memory during training and inference.
@@
 Proper optimizations are important.
-For example, we found it hard to reproduce quickly with some of the open-sourced baselines [5], which is roughly 6x slower than our implementation when testing with the same size of LLM.
+For example, we found it hard to reproduce quickly with some of the open-sourced baselines (BehnamGhader et al., 2024), which is roughly 6x slower than our implementation when testing with the same size of LLM.
 Our experiments involve training with 23 train sets and inference with another 23 test sets, which requires proper optimizations urgently.
-We import other optimizations based on the implementations  [^3] Please refer to [48]’s implementation on https://github.com/ma787639046/tdro.
-released by tDRO [48].
-These optimizations include 1) A distributed multi-node inference framework based on PyTorch RPC [42]; 2) Sequence packing [35] with Flash Attention 2 [18], which gives roughly 50% training speedup due to the removal of the pad tokens.; 3) Fused Kernels from Liger-Kernel [24], which gives roughly 10% speedup to both training and inferencing; 4) Multi-vector gradient accumulation support with GradCache [23] for optionally enlarging batch sizes if necassary.
-A.8 Detailed main results
+We import other optimizations based on the implementations released by tDRO (Ma et al., 2025).
+These optimizations include 1) A distributed multi-node inference framework based on PyTorch RPC (Li et al., 2023); 2) Sequence packing (Kundu et al., 2024) with Flash Attention 2 (Dao, 2024), which gives roughly 50% training speedup due to the removal of the pad tokens.; 3) Fused Kernels from Liger-Kernel (Hsu et al., 2024), which gives roughly 10% speedup to both training and inferencing; 4) Multi-vector gradient accumulation support with GradCache (Gao et al., 2021) for optionally enlarging batch sizes if necassary.
+A.9 Detailed main results
 The detailed main results (nDCG@10) on each dataset are shown in the tables below, which are full versions of Table 2.
@@
 BM25 31.5 29.9 21.3 31.3 75.3 23.6 60.3 22.8
+USTAD 34.9 30.6 22.5 35.9 76.9 29.5 56.0 46.6
 BGE-m3-dense 53.9 38.4 26.5 38.5 77.2 40.8 68.6 37.6
 BGE-m3-sparse 36.2 30.5 26.7 24.4 85.9 26.9 70.3 17.1
+SPLADE-v3 50.9 34.7 23.3 45.0 79.6 37.4 69.2 46.8
 LLM2Vec-llama3-8b 62.8 48.3 34.3 48.3 90.2 55.3 71.8 43.2
@@
 BM25 32.5 32.9 78.9 15.8 66.5 65.6 36.7 41.7
+USTAD 30.7 50.8 81.4 14.4 55.5 72.3 24.7 44.2
 BGE-m3-dense 31.5 60.1 88.4 15.2 63.6 47.0 20.6 47.2
 BGE-m3-sparse 28.3 20.3 73.4 12.2 63.3 52.7 27.8 39.7
+SPLADE-v3 35.7 58.6 81.4 15.8 71.0 74.8 29.3 50.2
 LLM2Vec-llama3-8b 41.8 64.2 87.2 23.0 78.2 80.3 20.5 56.6
@@
 LightRetriever Hybrid 34.0 76.7 88.6 58.0 76.2 53.9 83.0 61.4 66.5
+A.10 Additional results on CoIR and BRIGHT
+Benchmarks of our query-lightweight design on CoIR (Code Retrieval) and BRIGHT (Reasoning-intensive Retrieval) have been conducted using three backbones (Llama-3.2-1B, Llama-3.1-8B, Qwen-2.5-7B).
+Results are shown below in Table 17 and 18.
+As code retrieval and reasoning-intensive retrieval are more challenging out-of-domain (OOD) tasks requiring deep query or structure understanding, our proposed query-lightweight encoders are less performant compared to strong LLM embedding models like E5-Mistral (-6.2pp on CoIR, -6.3pp on BRIGHT), which is expected in our work.
+Our method is more suitable for normal relevance-based retrieval.
+Queries requiring complicated instructions or code understanding are not suitable to be applied to the query-lightweight encoder.
+Table 17: Retrieval performance (nDCG@10) on the CoIR benchmark.
+Model Apps CosQA SyntheticText2SQL CodeSearchNet CodeSNCC CodeTransContest CodeTransDL StackOverFlowQA CodeFeedBackMT CodeFeedBackST Avg
+BM25 1.0 14.0 16.9 26.8 34.7 50.1 8.7 56.8 34.7 54.3 29.8
+BGE-M3 7.4 22.7 48.8 43.2 47.6 47.9 31.2 61.0 33.5 49.9 39.3
+E5-Mistral 21.3 31.3 66.0 54.3 65.3 82.6 33.2 91.5 33.7 72.7 55.2
+LightRetriever
+Llama3.2-1B 11.2 30.1 52.0 60.1 58.6 57.7 27.9 77.8 29.6 64.5 47.0
+Llama3.1-8B 18.1 29.5 54.5 57.8 59.0 59.7 25.7 82.0 33.8 68.1 48.8
+Qwen2.5-7B 18.4 30.0 54.0 57.7 61.8 61.5 22.3 84.4 33.1 66.3 49.0
+Table 18: Retrieval performance (nDCG@10) on the BRIGHT benchmark.
+Model Bio. Earth. Econ. Psy. Rob. Stack. Sus. Leet. Pony AoPS TheT. TheQ. Avg
+BM25 18.9 27.2 14.9 12.5 13.6 18.4 15.0 24.4 7.9 6.2 10.4 4.9 14.5
+BGE-large 11.7 24.6 16.6 17.5 11.7 10.8 13.3 26.7 5.7 6.0 13.0 6.9 13.7
+E5-Mistral 18.6 26.0 15.5 15.8 16.3 11.2 18.1 28.7 4.9 7.1 26.1 26.8 17.9
+LightRetriever
+Llama3.2-1B 13.7 21.4 11.2 11.1 18.5 13.0 12.7 15.2 4.0 2.9 6.5 6.0 11.4
+Llama3.1-8B 14.2 22.3 12.6 11.1 17.1 13.9 15.0 16.6 8.1 0.7 4.0 5.7 11.8
+Qwen2.5-7B 19.3 24.0 13.8 13.1 14.1 11.3 14.3 11.8 5.2 1.1 5.8 5.7 11.6